### v1

11email: tshenar@tau.ac.il 22institutetext: ESO - European Southern Observatory, Karl-Schwarzschild-Strasse 2, 85748 Garching bei München, Germany 33institutetext: Institute of Astronomy, KU Leuven, Celestijnenlaan 200D, 3001 Leuven, Belgium 44institutetext: Department of Physics & Astronomy, Hounsfield Road, University of Sheffield, Sheffield, S3 7RH, United Kingdom 55institutetext: Instituto de Astrofísica de Canarias, C. Vía Láctea, s/n, 38205 La Laguna, Santa Cruz de Tenerife, Spain 66institutetext: Universidad de La Laguna, Dpto. Astrofísica, Av. Astrofśico Francisco Sánchez, 38206 La Laguna, Santa Cruz de Tenerife, Spain 77institutetext: Escola de Ciências e Tecnologia, Universidade Federal do Rio Grande do Norte, Natal, RN 59072-970, Brazil 88institutetext: Zentrum für Astronomie der Universität Heidelberg, Astronomisches Rechen-Institut, Mönchhofstr. 12-14, 69120 Heidelberg, Germany 99institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Newcastle upon Tyne, NE1 7RU, UK 1010institutetext: Heidelberger Institut für Theoretische Studien, Schloss-Wolfsbrunnenweg 35, 69118 Heidelberg, Germany 1111institutetext: Universität Heidelberg, Department of Physics and Astronomy, Im Neuenheimer Feld 226, 69120 Heidelberg, Germany 1212institutetext: Royal Observatory of Belgium, Avenue Circulaire/Ringlaan 3, B-1180 Brussels, Belgium 1313institutetext: Anton Pannekoek Institute for Astronomy, University of Amsterdam, Science Park 904, 1098 XH Amsterdam, the Netherlands 1414institutetext: Max-Planck-Institute for Astrophysics, Karl-Schwarzschild-Strasse 1, 85748 Garching, Germany 1515institutetext: European Space Agency (ESA), ESA Office, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 1616institutetext: ICREA, Pg. Lluís Companys 23, E08010 Barcelona, Spain 1717institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí Franquès 1, E08028 Barcelona, Spain 1818institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, United Kingdom 1919institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 2020institutetext: Institute of Science and Technology Austria (ISTA), Am Campus 1, 3400 Klosterneuburg, Austria 2121institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2222institutetext: Department of Astronomy and Physics, Saint Mary’s University, 923 Robie Street, Halifax, B3H 3C3, Canada 2323institutetext: Gemini Observatory/NSF’s NOIRLab, Casilla 603, La Serena, Chile 2424institutetext: Center for Computational Astrophysics, Division of Science, National Astronomical Observatory of Japan, 2-21-1, Osawa, Mitaka, Tokyo 181-8588, Japan 2525institutetext: Astronomical Institute, Academy of Sciences of the Czech Republic, Fričova 298, CZ-251 65 Ondřejov, Czech Republic 2626institutetext: Department of Astronomy and Astrophysics, University of Toronto, 50 St. George Street, Toronto, Ontario, M5S 3H4, Canada 2727institutetext: Dublin Institute for Advanced Studies, DIAS Dunsink Observatory, Dunsink Lane, Dublin 15, Ireland 2828institutetext: Centro de Astrobiología (CSIC-INTA), campus ESAC, camino bajo del castillo s/n, 28 692 Villanueva de la Cañada, Spain 2929institutetext: School of Physics and Astronomy, Monash University, Clayton VIC 3800, Australia 3030institutetext: ARC Centre of Excellence for Gravitational-wave Discovery (OzGrav), Melbourne, Australia 3131institutetext: IAASARS, National Observatory of Athens, GR-15236, Penteli, Greece 3232institutetext: Institute of Astrophysics, FORTH, GR-71110, Heraklion, Greece 3333institutetext: Centro de Astrobiología (CSIC-INTA), Ctra. Torrejón a Ajalvir km 4, 28850 Torrejón de Ardoz, Spain 3434institutetext: Institut für Physik und Astronomie, Universität Potsdam, Karl-Liebknecht-Str. 24/25, 14476 Potsdam, Germany 3535institutetext: Lund Observatory, Division of Astrophysics, Department of Physics, Lund University, Box 43, SE-221 00, Lund, Sweden 3636institutetext: Department of Astronomy & Steward Observatory, 933 N. Cherry Ave., Tucson, AZ 85721, USA 3737institutetext: Observatório Nacional, R. Gen. José Cristino, 77 - Vasco da Gama, Rio de Janeiro - RJ, 20921-400, Brazil 3838institutetext: Institute of Astronomy, Faculty of Physics, Astronomy and Informatics, Nicolaus Copernicus University, Grudziadzka 5, 87-100 Torun, Poland 3939institutetext: Department of Particle Physics and Astrophysics, Weizmann Institute of Science, Rehovot 7610001, Israel 4040institutetext: Lennard-Jones Laboratories, Keele University, ST5 5BG, UK 4141institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA 4242institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 4343institutetext: Armagh Observatory, College Hill, Armagh, BT61 9DG, Northern Ireland, UK
Binarity at LOw Metallicity (BLOeM)††thanks: Based on observations collected at the European Southern Observatory under ESO program ID 112.25W2.,††thanks: Tables … are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
Surveys in the Milky Way and Large Magellanic Cloud revealed that the majority of massive stars will interact with companions during their lives. However, knowledge of the binary properties of massive stars at low metallicity, which approaches the conditions of the Early Universe, remains sparse. We present the Binarity at LOw Metallicity (BLOeM) campaign – an ESO large programme designed to obtain 25 epochs of spectroscopy for 929 massive stars in the Small Magellanic Cloud (SMC) – the lowest metallicity conditions in which multiplicity is probed to date (). BLOeM will provide (i) the binary fraction, (ii) the orbital configurations of systems with periods yr, (iii) dormant black-hole binary candidates (OB+BH), and (iv) a legacy database of physical parameters of massive stars at low metallicity.
The stars are observed with the LR02 setup of the giraffe instrument of the Very Large Telescope (3960 – 4570Å resolving power ; typical signal-to-noise ratio S/N) . This paper utilises the first 9 epochs obtained over a three-month time. We describe the survey and data reduction, perform a spectral classification of the stacked spectra, and construct a Hertzsprung-Russell diagram of the sample via spectral-type and photometric calibrations.
The sample covers spectral types from O4 to F5, spanning the effective temperature and luminosity ranges and and initial masses . It comprises 159 O-type stars, 324 early B-type (B0–3) dwarfs and giants (luminosity classes V–III), 309 early B-type supergiants (II–I), and 137 late-type supergiants. At least 75 stars are Oe/Be stars: 20 O-type and 55 B-type (13% and 10% of the respective samples). In addition, it includes four high-mass X-ray binaries, three stars resembling luminous blue variables, two bloated stripped-star candidates, two candidate magnetic stars, and 74 eclipsing binaries.
Key Words.:
stars: massive – binaries: close – binaries: spectroscopic – Magellanic Clouds1 Introduction
Massive stars (), typically classified as OB-type stars on the main sequence, play an increasingly important role in modern astrophysics, with direct links to stellar dynamics, stellar transients, and gravitational-wave (GW) astrophysics (Portegies Zwart & Verbunt, 1996; Langer, 2012; Woosley, 2017; Wang et al., 2020; Marchant & Bodensteiner, 2023). Spectroscopic and interferometric surveys in the Milky Way (MW) and the Large Magellanic Cloud (LMC) have shown that stellar multiplicity is common among massive stars (e.g., Abt, 1983; Kobulnicky & Fryer, 2007; Mason et al., 2009; Sana et al., 2012, 2014; Bordier et al., 2023); over 50% of massive stars are expected to interact with a companion during their lifetime (Sana et al., 2012, 2013; Stegmann et al., 2022; Offner et al., 2023; Kummer et al., 2023). Such interactions dramatically alter the evolutionary paths and final fates of massive stars, resulting in a plethora of astronomical phenomena such as stripped helium stars and Wolf-Rayet (WR) stars (Paczyński, 1967; Shenar et al., 2016; Götberg et al., 2018; Pauli et al., 2022a; Drout et al., 2023), rapidly rotating stars with decretion disks (Oe/Be stars; Pols et al., 1991; Rivinius et al., 2013; de Mink et al., 2013; Wang et al., 2017; Bodensteiner et al., 2020; Britavskiy et al., 2023; Renzo & Götberg, 2021), stellar mergers and magnetic stars (Ferrario et al., 2009; de Mink et al., 2014; Schneider et al., 2019; Shenar et al., 2023; Frost et al., 2024), single-degenerate binaries and high-mass X-ray binaries (Corral-Santana et al., 2016; Shenar et al., 2022a; Mahy et al., 2022), and GW sources (de Mink & Mandel, 2016; Marchant et al., 2016; Tauris et al., 2017; Mandel & Broekgaarden, 2022; Mandel & Farmer, 2022).
Of special interest is massive-star research in low metallicity (low ) environments, which reflect the conditions prevalent in the distant Universe. An increasing number of transients, such as long-duration -ray bursts (LGRBs; Yoon & Langer, 2005; Woosley & Heger, 2006), superluminous supernovae (SNe; Quimby et al., 2011; Gal-Yam, 2012), broad-lined type Ic supernovae (Modjaz et al., 2008), and pair-instability SNe (Barkat et al., 1967; Fryer et al., 2001; Langer et al., 2007; Woosley, 2017; Farmer et al., 2019) are thought to be associated mainly or exclusively with low- conditions. Similarly, the bulk of black-hole (BH) mergers observed with the LIGO-Virgo-KAGRA collaboration are thought to originate in low metallicity conditions (Abbott et al., 2016; Giacobbo et al., 2018; Klencki et al., 2018).
Modern investigations expose deficiencies in our understanding of massive stars at low . For example, the rates and mass distribution of BH mergers defy original expectations (Broekgaarden et al., 2021; Mandel & Broekgaarden, 2022; van Son et al., 2022), and observables such as the rate of LGRBs (Graham & Fruchter, 2017; Chen et al., 2017) and the fraction of Oe/Be stars and Be X-ray binaries (Haberl & Sturm, 2016; Schootemeijer et al., 2022) as a function of are not reproduced by contemporary models (Graham & Fruchter, 2017; Chen et al., 2017; Hastings et al., 2021). Such discrepancies are likely related to insufficient knowledge of massive-star evolution at low , or to a false extrapolation of the initial conditions of massive stars (e.g., binary fraction, orbital configurations) to low . The recent detection of a BH with a low-metallicity companion through Gaia astrometry provides additional evidence suggesting that low- environments are crucial for the formation of massive BHs (Gaia Collaboration et al., 2024).
While it has been shown that for solar-type stars binary properties can depend on the natal metallicity, this remains only a prediction for massive stars (Kroupa, 2001; Saigo et al., 2004; Machida, 2008; Marks et al., 2012; Moe et al., 2019; Price-Whelan et al., 2020). To mitigate this, spectroscopic monitoring surveys of massive-star populations at different environments are needed, which are sensitive to the regime of binary interactions (i.e. orbital periods ). A notable spectroscopic campaign in this context was the VLT-FLAMES Tarantula Survey (VFTS; PI: Evans), which monitored massive stars in the Tarantula nebula of the LMC (Evans et al., 2011), with a metallicity of . The VFTS survey yielded a comparable intrinsic binary fraction for OB-type stars in the LMC (50-60%; Sana et al. 2013; Dunstall et al. 2015) to that observed in different Galactic environments (50-70%, e.g., Sana et al., 2012; Kiminki & Kobulnicky, 2012; Banyard et al., 2022). The follow-up Tarantula Massive Binary Monitoring (TMBM; PI: Sana) and B-type Binary Characterisation (BBC; PI: Taylor) programmes also revealed an overall similar distribution of orbital parameters and mass ratios to Galactic samples (Almeida et al., 2017; Villaseñor et al., 2021; Shenar et al., 2022b; Mahy et al., 2020b). However, the LMC metallicity only differs by a factor of from that of the MW.
The Small Magellanic Cloud (SMC) is a neighbouring dwarf galaxy with (Hunter et al., 2007) located about 62 kpc from Earth (Graczyk et al., 2020), hosting thousands of massive stars (Humphreys & McElroy, 1984). It had a star formation peak Myr ago (Antoniou et al., 2010; Rubele et al., 2015; Schootemeijer et al., 2021), potentially triggered by a collision with the LMC about 100-150 Myr ago (e.g. Zivick et al., 2018). While galaxies of lower metal content exist in the Local Group (e.g., Sextans A, Skillman et al. 1989; Lorenzo et al. 2022; Leo P, McQuinn et al. 2015; Evans et al. 2019; Telford et al. 2023), the SMC is the only galaxy in which a large sample of massive stars at low can currently be resolved and spectroscopically monitored with sufficient spectral resolution and signal-to-noise (S/N). Previous or ongoing surveys addressed aspects related to the SMC massive-star contents (Humphreys & McElroy, 1984; Evans et al., 2006; Martayan et al., 2007; Evans et al., 2004; Schootemeijer et al., 2021), stellar winds and mass-loss (Ramachandran et al., 2019; Vink et al., 2023) and runaway status (RIOTS4, Lamb et al., 2016), but multiplicity has been largely neglected beyond analyses of selected eclipsing binaries (Hilditch et al., 2005), clusters (Dufton et al., 2019; Bodensteiner et al., 2021), and individual objects of interest (e.g., Pauli et al., 2022b). Moe & Di Stefano (2013) investigated the frequency of eclipsing massive binaries among B-type stars in the SMC, LMC, and MW, and found no significant differences between the populations. However, their study was limited to the period range d, and is generally bias-dominated given the small fraction of eclipsing binaries. Further information regarding multiplicity in the SMC is not known.
The need to establish the multiplicity of massive stars at low is not the only reason to monitor massive stars in the SMC. Recent spectroscopic monitoring of massive stars in the MW (Mahy et al., 2022) and the LMC (Shenar et al., 2022b, a) uncovered the tip of the iceberg of a new population of massive single-degenerate binaries: X-ray dormant OB+BH binaries (e.g., Giesers et al., 2018). Such binaries yield precious constraints on core-collapse mechanisms and the presence of SN explosions and possible natal kicks during the collapse into BHs (Mirabel & Rodrigues, 2003; Renzo et al., 2019; Atri et al., 2019; Langer et al., 2020; Banagiri et al., 2023). For example, Vigna-Gómez et al. (2024) recently used the dormant OB+BH binary VFTS 243 to derive a natal kick of 4 km and an ejection of 0.3 neutrino mass during the collapse of the progenitor. The Gaia mission will likely uncover dozens more OB+BH binaries in the MW via high-precision astrometry (Mashian & Loeb, 2017; Breivik et al., 2017; Janssens et al., 2022, 2023), though so far only low-mass stars with BH companions were discovered (El-Badry et al., 2023b, a; Shahaf et al., 2023; Gaia Collaboration et al., 2024). The formation scenarios of these objects are still debated, and it is possible that they originated via dynamical captures, making them less constraining for SN physics (Rastello et al., 2023; El-Badry, 2024; Marín Pina et al., 2024). In any case, Gaia will not uncover extragalactic OB+BH binaries. Finding the first dormant OB+BH binaries in the SMC via spectroscopic monitoring of massive stars has the potential to yield unprecedented constraints on BH formation at low .
Finally, binary monitoring provides a crucial testbed for single-star models. Eclipsing binaries enable the determination of accurate stellar masses and radii (Hilditch et al., 2005; Torres et al., 2010; Mahy et al., 2020a). Moreover, well-separated binaries are less likely to have been affected by binary interaction, and hence provide a more solid basis for benchmarking endeavours of single-star models (de Mink et al., 2014).
Motivated by these reasons, we initiated a novel spectroscopic monitoring survey of a large population of massive stars in the SMC. The Binarity at LOw Metallicity (BLOeM) campaign is a European Southern Observatory (ESO) Large Programme (PI: Shenar, dPI: Bodensteiner; ID: 112.25R7) scheduled for 2023 – 2025. Relying on 116 hr of observing time with the Fibre Large Array Multi Element Spectrograph (FLAMES; Pasquini et al. 2002) of the Very Large Telescope (VLT), the survey is underway and assembling 25 epochs of spectroscopy for 929 massive stars for a total baseline of two years (four semesters). The survey will enable a full characterisation of the binary fraction and orbital parameters of the stars up to orbital periods of a few years and mass ratios as low as , discovery of dormant OB+BH binaries, and a complete analysis of the binary and single-star content of the sample.
This first paper in the series provides an overview of the sample and sample selection (Sect. 2), a description of the data reduction and quality (Sect. 3), a detailed spectral classification (Sect. 4), and a first characterisation of the physical mass range and evolutionary status of the sample stars (Sect. 5), followed by our main conclusions (Sect. 6).
2 Sample selection
The BLOeM sample (see Fig. 1) was selected from the third Gaia data release catalogue (Gaia DR3, Gaia Collaboration et al., 2023). The catalogue was retrieved from the Gaia database using a search radius of around the SMC center ( = 00:52:38, -72:48:01; epoch J2000). To achieve S/N per pixel (0.2 Å spectral bin), only stars with mag were retrieved. Foreground objects were filtered via two constraints. First, the parallax was required to be consistent with zero111Gaia is not sensitive enough to measure non-vanishing parallaxes within errors at the SMC distance. within 5, that is . Second, the proper motions of the stars were required to be consistent within 15 with the SMC proper motion ( and ) following Yang et al. (2019) and Schootemeijer et al. (2021).
We omitted the 12 known SMC Wolf-Rayet (WR) stars (Neugent et al., 2018) from the sample, since they have been previously monitored (Foellmi et al., 2003; Hainich et al., 2015; Shenar et al., 2016; Schootemeijer et al., 2024). Finally, we omitted potential red supergiants (RSGs) from the sample by imposing mag: due to the large radii of RSGs, RSG binaries have periods that exceed a few years (e.g. Patrick et al., 2019; Neugent et al., 2020) and hence exceeding the 2-year baseline of our programme.
To avoid crowding, we removed objects that have a Gaia source brighter than mag closer than 1.2”, which corresponds to the FLAMES fibre size. We also explicitly excluded stars within 30” around the centres of the dense SMC clusters NGC 330 and NGC 346.
In an attempt to select only massive stars, we made use of an evolutionary track of a star computed by Schootemeijer et al. (2019) with the Modules for Experiments in Stellar Astrophysics (MESA) stellar evolution code (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). We converted the physical parameters along the track to a -band magnitude and colours using bolometric corrections taken from the MIST webpage222http://waps.cfa.harvard.edu/MIST/model_grids.html; the bolometric corrections were retrieved by fitting a 5th order polynomial to MIST values for and . (Dotter, 2016; Choi et al., 2016). We adjusted the track on the color-magnitude diagram (CMD) by adopting a distance of 62 kpc (Graczyk et al., 2020), and assuming an average value for the reddening of mag and extinction of mag (Schootemeijer et al., 2021). We then only selected targets whose CMD positions lie above this track before it becomes a RSG (effective temperature kK; see Fig. 2). This resulted in a massive-star catalogue of 5576 stars subject to the criteria above. We also made use of a MESA track computed for by Schootemeijer et al. (2019) to divide the sample into stars with (born as O-type stars) and (born as B-type stars)333The threshold mass for O-type stars is typically taken as for the Galaxy (e.g. Martins et al., 2005). However, stars in the SMC are more compact and hot at a fixed mass (e.g. Georgy et al., 2013), such that this threshold is likely lower at low ..
Our programme includes a total of eight FLAMES plate configurations, each with a field-of-view (FoV) of diameter 25’, although the instrument setup ensures visibility of targets only within a 20’ diameter (Fig. 1). For each FLAMES plate configuration, 130 giraffe fibres are available. We allocated 14 fibres for sky, leaving each field with 116 science targets. The only exception is field 8, for which 13 sky fibres and 117 science targets are available. This makes a total of science targets.
To obtain a balanced sample of 929 science targets out of the 5576 available targets, we aimed to achieve a band magnitude distribution that is as homogeneous as possible, while prioritising the brightest and hence most massive stars, which are rarer. The target allocation was then followed via the two steps described below.
First, the centre of a FLAMES pointing was chosen. The choice of field centre followed automatically by identifying the coordinate which encloses as many massive and bright stars as possible within a 25’ diameter, after removing stars which were already allocated in previous field allocations. Specifically, the pointings were selected by identifying the centre coordinates that result in the largest number of stars with (see above) and mag. This resulted in fields 1 – 8 shown in Fig. 1, which provide a good coverage of the massive-star content of the SMC. We note that while some of the fields overlap (e.g., 1 & 4), there is no overlap in the allocated stars within each field.
As a second step, for each plate configuration, targets within a 20’ diameter were allocated to the available fibres, starting from the brightest ones, while aiming at a homogeneous sampling across the -band. Specifically, we divided the magnitude range 10 – 16.5 mag into 15 bins, and aimed at a homogeneous coverage across these bins, resulting in stars per magnitude bin, per field. This could not always be achieved, given the rarity of bright stars, and the smaller parameter range of massive stars at lower magnitudes. Since the final fibre allocation depends on limitations related to the FLAMES fibre positioner, the remainder of the massive stars in each field in a diameter of 25’ were taken as backup targets, with their priority sorted by brightness.
The final fibre allocation was then performed using ESO’s Fibre Positioner Observation Support Software (FPOSS). The majority of our input targets made it to the final allocation, but the final sample includes a few dozens of backup targets. Sky fibres were allocated to 14 fibres in each field (13 for field 8) selected from concentric rings around the field centre where no known Gaia source is located. Finally, guide stars were selected following requirements in the ESO FLAMES manual for cycle P112.
The final sample of 929 targets is shown on a CMD in Fig. 2, along with a magnitude histogram. We also show the completeness fraction with respect to the underlying SMC Gaia catalogue as a function of -band magnitude. Out of the 929 stars, and based on the single-star tracks in Fig. 2, 323 have initial masses above track (”born as O-type stars”) and 606 are below it (”born as B-type stars”). Evidently, the sample reaches a completeness fraction of for the sub-sample, and for the sample.
The naming convention for the sample stars follows the format F-NNN, where F is the field number (1 – 8), and NNN is the target number (001 – 117), sorted by ascending right ascension per field.
3 Observations and data reduction
At the time of writing, nine out of 25 epochs were obtained during the first semester and processed in the framework of the BLOeM survey. The field centres, along with the MJD values of the epochs so far acquired, are provided in Table 2 of Appendix A.
The data reduction was performed with the giraffe pipeline v. 2.16.11 under the ESO CPL environment (v. 3.13.5). Each exposure was split into two sub-exposures for a robust removal of cosmics (Sect. 3.5). The data reduction itself consisted of four steps: bias and dark subtraction, flatfield correction, and wavelength calibration. All spectra were resampled by the ESO CPL pipeline to a constant wavelength step of 0.2Å (see also Sect. 3.2) and science spectra were then sky-subtracted and corrected for the barycentric motion. As a final step, we resampled the individual spectra to a common wavelength grid and co-added the spectra of individual targets to boost S/N in order to improve spectral typing. The spectra have a spectral resolving power of and cover the spectral range Å, with a median S/N of per pixel and epoch; details are provided below. We provide further details below on specific aspects of the data reduction process444The reduced data and co-added spectra will be made available via ESO phase 3 upon termination of propriety time; they are currently available on http://www.astro.tau.ac.il/~tshenar/DR3/; please contact T. Shenar or J. Bodensteiner for credentials. .
3.1 Temporal sampling
The temporal sampling of the epochs is not strictly defined a-priori, to allow for sufficient scheduling flexibility. A minimum separation between each epoch of 1 d is requested, and no larger than 20 d to ensure that all epochs would be acquired within a semester. The typical separations between the epochs used here, acquired during Sep. 2023 – Dec. 2023, are days to weeks (see Table 2), for a total time baseline of 50 – 70 d, depending on the field. The fact that the acquisition will take place across four semesters ensures that both short-scale and long-scale variability will be covered by the survey. Since multiplicity analysis is beyond the scope of this current paper, we refrain from a complete characterisation and Fourier mapping of the temporal sampling here.
3.2 Wavelength calibration
We paid particular attention to the quality of the wavelength solution. For FLAMES giraffe, a reference ThAr calibration frame in the LR02 setup is obtained at the end of an observing night by illuminating each fibre on a given plate with the light of a ThAr lamp. As a result, each fibre of each epoch has its own calibration ThAr spectrum which is used by the CPL pipeline to produce a 2D polynomial dispersion solution. The wavelength calibration solution was performed in two steps. First we used the instrument model provided by the pipeline static calibration v2.16.11 to compute a first-guess solution and run a first iteration of the giwavecalibration CPL recipe. We modified the standard options to allow for a large detection window of 20 pixels at first, for five iterations, then we progressively reduced it to 15 and ultimately 10 in the remaining 5 iterations. We decreased the rejection threshold from 1.2 to 3, allowing us to retain more lines and provide a first wavelength solution with a root-mean-square (rms) residual of 0.45-0.56 pixels. By comparing ThAr spectra of different nights we noticed a slight drift in the wavelength solution in various parts of the ThAr spectrum, with a higher stability in the centre of the wavelength range and a larger epoch-to-epoch variation at the blue and red part of the wavelength solution. We estimated the internal consistency to be no better than a few km s-1. To improve on this, we performed a second iteration by using the first solution as a new input guess-solution and re-iterating the giwavecalibration recipe. This yields a final dispersion solution, characterised with a rms residual in the range of 0.22 to 0.25 pixels.
Three epochs of field 1 have been observed with the SIMCAL lamp on at the beginning of the survey. The SIMCAL lamp yields a set of five ThAr spectra spread across the detector and acquired simultaneously with the science observations. This setup was discontinued because the glow of the strongest ThAr lines impacted the adjacent fibres signal, leaving a noticeable imprint of ThAr on nearby sky and weak objects. Yet, we noticed no difference in the quality of the wavelength solution with or without the SIMCAL lamp.
We experimented with the pipeline rebin pixel size (--rbin-lstep) but did not find this to yield any significant improvement and we decided to resort to the default resampling of 0.2 Å. The average spectral resolving power measured on the ThAr lines across all epochs is .
Finally, we investigated the stability and consistency of the individual dispersion solutions across the BLOeM dataset. We specifically investigated the inter-epoch stability for given fibres as well as the intra-epoch consistency across all fibres of a given epoch. For the first one, which informs us about the temporal stability of the data, we cross-correlated the wavelength-calibrated ThAr spectra of a given fibre and field with that of the same fibre and field across all the epochs obtained so far. We found maximum peak-to-peak differences to be of 130 m s-1, with a standard deviation below 50 m s-1. For the latter, which informs us about the consistency of the dispersion solution across the BLOeM data sets, we cross-correlated the ThAr spectra of all fibres for a given epoch and field with one arbitrary ThAr spectrum (fibre #10) for the epoch under consideration. While the peak-to-peak and rms variations are slightly larger (750 and 150 m s-1 in the worse case), they remain well within the specifications of the instrument. For a few nights, no ThAr calibration frame could be obtained in the morning following the observations. In such cases, we use the frame closest in time, typically the one from the morning before. In eight cases, however, we had to resort to ThAr calibration frames taken 30 to 40 h apart of the science. Yet, no noticeable difference in the quality of the calibration could be found, again suggesting that temporal drifts are limited.
3.3 Sky-subtraction
Sky spectra were obtained simultaneously with the science integration through a set of fibres allocated to empty patches of sky. These are dubbed SKY fibres and record all background signal, including the moon, nebular and sky emission spectra depending on the wavelength regime. Hence not all signal recorded by SKY fibres are from the ‘sky’ itself, but we nonetheless adopt the generic denomination here.
As described earlier, we typically used 14 SKY fibres in each field and kept the location constant across all epochs of a given field. We visually inspected the SKY spectra from each field and each epoch and flag spectra that seemed to be significantly higher than the median of the epoch and field. These are possibly contaminated by faint objects and therefore not representative of the true background signal. Once a sky location has been flagged as contaminated in any of the epochs of a given field, it is rejected from all epochs so that the median sky is always computed with the same set of input locations. In this process, we rejected 4, 3, 1 and 2 sky fibres for fields 1, 2, 5 and 6 respectively. The sky correction was finally performed by subtracting the median spectra of the ‘valid’ SKY fibres at the corresponding field and epoch. The error sky spectrum was computed as the rms around the median sky at each pixel after masking sky pixels affected by cosmic rays through a iterative filtering. The error sky spectrum was added quadratically to the error science non-sky-subtracted spectrum produced by the pipeline in order to compute the error spectrum of the sky-subtracted science data.
Of importance, any nebular component in the sky spectra is obtained at the position of the SKY fibres. In our adopted observational setup, these are located tens of arcsecs to several arcminutes away from any science spectra hence they do not reflect the local nebular conditions of any science targets. The process of taking the median across 10 to 14 sky positions ensures that any local nebulosity in the SKY fibres is averaged out. As a corollary, the sky-subtracted science spectra are not corrected for nebulosity and therefore retain their nebular component.
3.4 Normalisation
The sky-subtracted spectra of each sub-exposure were individually and automatically normalised using a designated Python script to remove the underlying continuum, which is a combination of the stellar continuum, reddening, and instrumental response. Overall, the non-normalised spectra are well-behaved and have a smooth, typically monotonically decreasing behaviour across the spectral range. The normalisation was performed by fitting a polynomial to automatically-identified continuum points along the spectrum. Following several independent attempts, we found that a polynomial of degree eight provided the best results in terms of robust normalisation; lower polynomial degrees resulted at times in underfit regions, while higher degrees introduced wavy patterns in the spectra which can impact subsequent analysis.
The continuum points were automatically selected in an iterative manner. For this purpose, we computed an average spectrum with a small 3-pixel window and a median spectrum with a large 250-pixel window, which is a first approximation for the continuum. A first batch of continuum points was defined as the set of points whose average flux is 2 above the median flux, with originating in the error spectrum. To ensure that the edges of the spectra are included, we always include the first and last 10 pixels not affected by cosmics in this set. We then fit a polynomial of degree eight through these points to obtain an approximation for the continuum. This process is then repeated three times, with the polynomial fit for the continuum replacing the role of the median spectrum of the original iteration.
3.5 Cosmic correction and combination of sub-exposures
To boost the S/N, the two sub-exposures of each epoch and star were combined into one exposure. Moreover, we utilised the availability of two sub-exposures for a robust cosmic correction using a designated Python script. The process is as follows. First, positive flux outliers are identified in each of the two sub-exposures via the condition , where is the minimum of the error spectra of both sub-exposures. Cosmics are identified as points flagged as such in one spectrum, but not in the other – this ensures that intrinsic emission features (e.g., disk features, wind features, nebular lines) are not removed via this process. In principle, cosmics may be present in the same pixels in both sub-exposures, but such cases occur so rarely that this does not warrant further consideration. The cosmics are then removed by replacing each pixel identified as cosmic in one sub-exposure with the flux value of the same pixel in the second sub-exposure. The flux of the remaining pixels is formed by quadratically adding the two sub-exposures. The S/N per pixel ranges between 20 – 300, with a median value between 70 – 120, depending on the field (Fig. 3).
3.6 Co-added spectra
For a more robust spectral classification, we produced high S/N spectra by stacking the nine available epochs for each target. To achieve this, the cross-correlation occurred in two steps. First, the spectra were shifted to the rest frame by cross-correlating them with a spectral model and then co-added, which enabled us to generate a first co-added spectrum. In this step, the entire spectral range was considered, which is dominated by the Balmer lines H and H. As a second step, we cross-correlated all the epochs against the co-added spectrum formed in the previous step, which is calibrated to the rest frame. This time, however, we used a narrow spectral window around He i (namely 4460–4485 Å), which is present in essentially all objects555For AF supergiants, this range is rich in other spectral lines such as Mg ii , making the cross-correlation in this range equally possible.. This allows for a more accurate radial-velocity (RV) measurement that is tuned to the spectral morphology of the individual star. For the first step, we used a pre-computed model from the TLUSTY O-star model atmosphere grid (Hubeny & Lanz, 1995; Lanz & Hubeny, 2003) for 666This is the closest available metallicity to that of the SMC. In any case, the impact of the metallicity on RV measurements is negligible, an effective temperature of kK, and a surface gravity of []. The exact choice of the model can impact the absolute RV calibration, but has no impact on the process of spectral classification which is the focus of this paper.
We note that for double-lined spectroscopic binaries (SB2), the process of co-adding the data will smear the spectral features of the two components and result in an average spectrum, which may well contaminate the classification. Clear cases of SB2s (40/929 in total) were identified via visual inspection of the available epochs (see also Sect. 4). A montage of the co-added spectra of selected stars, ordered by descending spectral type (Sect. 4), is shown in Figs. 4 – 7.
4 Spectral classification
We have established SMC reference OB stars from comparison with Galactic O star templates from Sota et al. (2011) and Maíz Apellániz et al. (2016) or Galactic B star templates from Negueruela et al. (2024). These stars are drawn from BLOeM datasets, supplemented by archival VLT/FLAMES (Evans et al., 2006; Dufton et al., 2019) or VLT/X-Shooter (Vink et al., 2023) datasets. Reference stars are ideally sharp-lined, permitting rotational broadening to be applied for comparison with fast rotators. We assign e, e+ and pe for stars exhibiting (i) H i, (ii) H i & Fe ii, and (iii) H i & He i emission, respectively. The qualifier ”:” deems the classification uncertain. We note that the BLOeM spectral range includes most diagnostic lines, but misses a few lines which can help refine the classification such as N iii , He ii , H, and H.
Spectral types of O to B0.7 stars are determined following the Galactic O star scheme of Sota et al. (2011) and Maíz Apellániz et al. (2016) which utilises the ratio of He ii 4542 to He i 4471 or 4388 for late subtypes (O8.5+). Luminosity classes are usually assigned courtesy of He ii 4686. Since this line is not available for the BLOeM dataset, luminosity classes for O4–7 stars are estimated from the detection of N iv 4058 and Si iv 4088-4116 emission lines. For late O and early B subtypes we use a combination of H and H line profile morphologies supplemented by the ratio of Si iv 4088, 4116 to He i 4121, 4144 lines, following Walborn & Fitzpatrick (1990). It is well known that incorrect luminosity classes would be assigned for silicon-poor SMC stars based on Galactic templates (Walborn, 1983). For example, Walborn et al. (2014) showed that using Si-He criteria results in different luminosity classes than He-line criteria alone. Hence, SMC reference stars are essential for this approach. Nomenclature linked to the region of He ii 4686 such as f and (f), which reflect the presence of emission in the line, is not possible, but we are able to flag potential rapidly rotating O stars via (n), n, nn etc. on the basis of the FWHM of He ii 4542 (early and mid O-types) and He i 4471 (late O-types). For a few O-type stars, spectral types (mostly luminosity classes) were adjusted based on previous literature, given the lack of diagnostics such as He ii and H in the BLOeM dataset. These cases are documented in the comments in Table LABEL:tab:SpT.
We follow the SMC B-type scheme of Lennon (1997) in assigning spectral types of B1–9 stars from Si iv 4088–4116, Si iii 4553, Si ii 4128–32, plus the ratio of He i 4471 to Mg ii 4481 at late subtypes. Luminosity classes are obtained from line morphologies of H and H with respect to reference stars. Unusual B-type systems with Balmer emission lines and forbidden [Fe ii] lines are classified as supergiant B[e] (sgB[e]) following Lamers et al. (1998); Kraus (2019).
For A and F supergiants we are unable to follow the Ca ii H+K criteria of Evans & Howarth (2003) since our dataset excludes the Fraunhofer K line, so again we have selected SMC reference A0–5 supergiants from archival VLT/FLAMES spectroscopy, with luminosity classes assigned from the equivalent width of H from comparison with Millward & Walker (1985). Spectral types assigned to late A and F stars are relatively coarse, and are primarily based on the strength of the CH G-band in BLOeM datasets spectrally degraded to the resolution of SMC AF supergiants presented by Evans & Howarth (2003).
The derived spectral types are compiled in Table LABEL:tab:SpT for each of our targets. While a multiplicity analysis is beyond the scope of this paper, a few dozen () already portray clear evidence for the presence of at least two stellar components in their spectra. These are visually identified and, when possible, a preliminary spectral type for the companion(s) is provided.
5 Results
5.1 Census
A histogram of the spectral type distribution across the range O4 to F5, subdivided into dwarfs/giants and supergiants, is shown in Fig. 8. The BLOeM sample includes the complete spectral range from the earliest stars in the SMC (excluding Wolf-Rayet stars) to yellow supergiants. Overall, the sample includes 159 O-type stars, 324 early B-type dwarfs and giants (B0 – B3 V–III), 309 early B-type supergiants (B0 – B3 II–I), and 137 late-type supergiants (B4 and later). Of these, 75 objects are classified as Oe/Be stars, portraying characteristic emission in their Balmer lines: 20 Oe stars (including four uncertain cases, marked as ”e?”), and 55 Be stars (Sect. 5.2).
The number of dwarfs/giants (generally interpreted as main-sequence objects) steeply decreases around a spectral type of B2, which roughly corresponds to (Harmanec, 1988). Among the B1-B2 dwarfs/giants, the vast majority are giants, bright enough to have been included in our survey. These are likely evolved main sequence stars. In the O-star regime, we are sensitive to the full extent of the main sequence down to the youngest stars, assuming those are not dust-enshrouded.
5.2 Oe/Be fraction: complementary Gaia low-resolution spectra
As noted in Sect. 5.1, 75 targets have been identified as Oe/Be stars in our sample. However, the BLOeM dataset doesn’t cover the H line, which is the most sensitive line for circumstellar material in the visual range. Due to this, it is likely that we have missed Oe/Be stars in the sample which exhibit H emission but not H, which is covered by BLOeM. To mark additional Oe/Be candidates, we make use of Gaia XP low-resolution flux-calibrated spectra (De Angeli et al., 2023), which exist for all sources except BLOeM 4-041, 6-044, 7-012 and 8-033. While these data are of very low resolution (), and some are affected by problematic wiggles in the data, they may be used to detect the presence of strong emission, or candidate OBe stars in our case. To be pragmatic, we adopt a straightforward approach of measuring the H equivalent width (EW) in the XP ‘sampled’ data, using broad windows on both sides of the line to define a local continuum. While more complex approaches are possible (e.g., Weiler et al., 2023), our approach should be adequate for detecting strong H emission. Candidate OBe stars are then flagged as 1- outliers in the distribution of equivalent width as a function of magnitude. As a check on this approach, a visual inspection of all spectra was also performed to define a second category of ‘by eye’ OBe candidates. Both samples are illustrated in Fig. 9. Both methods agree for most targets, although given the contamination induced by the spectral wiggles on the EWs, we favour the by-eye classification of H emitters.
Overall, in addition to the 75 Oe/Be stars classified using the BLOeM dataset, we identify 28 additional H emitters. However, the majority of those are supergiants, which are not Oe/Be stars per definition, or stars with identified nebular contamination. Only seven targets are potential new Be stars. Adding those, the total Oe/Be fraction among the non-supergiant OB stars is 11%. This fraction is much lower than the fraction of 30% that has been found in previous work (Bonanos et al., 2010; Schootemeijer et al., 2022). The difference can be explained by the diagonal CMD cut that we use to select our sample, which favours blue stars and hence disfavours OBe stars, which are about 0.3 magnitudes redder than OB stars (Schootemeijer et al. 2022, their figure A.5).
The bias mentioned above likely does not impact the Oe fraction derived here, which is 20/159 = 13%, comparable to the 10% fraction reported by Bonanos et al. (2010). However, the Be fraction among early B-type stars in the sample (removing supergiants) is 55/560 = 10%, which is likely strongly underestimated due to the exclusion of reddened stars.
5.3 Notable targets
We highlight a few unique objects in our sample. BLOeM 2-116, BLOeM 3-012, and BLOeM 4-055 are classified as sgB[e] stars. BLOeM 2-116 (alias LHA 115-S 18) was subject of several studies (e.g., Shore et al., 1987; Clark et al., 2013; Maravelias et al., 2014), as was BLOeM 3-012 (alias RMC 4; Zickgraf et al. 1996; Graus et al. 2012; Pasquali et al. 2000; Wu et al. 2020). Such objects are thought to represent a brief evolutionary phase of massive stars. Their spectra resemble those of luminous blue variables (LBVs), and their origin is debated in the literature (e.g. Podsiadlowski et al., 2006; Clark et al., 2013). The spectra of all these objects indicate variability. Whether this variability stems from binary motion is not yet clear, though BLOeM 3-012 has previously been reported to be a binary (Zickgraf et al., 1996) and seems well explained as the product of a binary merger in a triple system (Pasquali et al., 2000; Podsiadlowski et al., 2006; Wu et al., 2020).
BLOeM 3-031 and BLOeM 5-071 show spectra which resemble those of Be + bloated stripped-star binaries such as LB-1 (Irrgang et al., 2020; Abdul-Masih et al., 2020; Shenar et al., 2020; El-Badry & Quataert, 2021) and HR~6819 (Bodensteiner et al., 2020; Frost et al., 2022). Such objects feature strong Balmer emission characteristic of Be stars, in combination with narrow and RV variable absorption features which stem from a putative bloated stripped-star companion.
BLOeM 2-104 and BLOeM 4-039 were classified as Of?p stars by Evans et al. (2004), which refers to the presence of emission lines associated with N iii and C iii in the range 4630–4660Å (Walborn, 1972). We cannot independently verify this since the BLOeM data lack this range, thoug we do confirm the presence of emission in the Balmer lines characteristic of such stars, and so we adopt this classification. All known Galactic Of?p stars possess strong, global magnetic fields (Grunhut et al., 2017). It is therefore plausible that BLOeM 2-104 and BLOeM 4-039 are magnetic as well, although spectropolarimetric data are needed to verify this.
Nine targets are identified as significant X-ray emitters from a cross-match with various X-ray catalogues described in Appendix B, one of which may be a spurious X-ray detection (BLOeM 6-116). Four of those have been classified as high-mass X-ray binaries (HMXBs) in the past: BLOeM 2-055, 2-82, BLOeM 2-116, BLOeM 4-026 and BLOeM 4-113. BLOeM 3-042 may be a colliding-wind binary (CWB) given its classification (O6 III + O7.5). The origin of X-rays in BLOeM 8-029 (B1 IV:) and BLOeM 1-102 (O6 III(n)) is not yet clear. More details regarding the cross-match process and X-ray luminosities of the objects can be found in Appendix B.
Finally, cross-matching with the OGLE catalogue of photometrically variable stars in the SMC (Pawlak et al., 2016), 74/929 (i.e., 8%) of the BLOeM targets are found to be eclipsing binaries, while eight are classified as ellipsoidal variables (Sect. B.4).
5.4 Hertzsprung–Russell Diagram
Spect. | Dwarf | Giant | Supergiant | |||
Type | ||||||
kK | mag | kK | mag | kK | mag | |
O4 | 46.0 | –4.99 | 41.7 | –4.87 | 40.2 | –4.78 |
O5 | 41.3 | –4.84 | 39.9 | –4.63 | 38.5 | –4.44 |
O6 | 39.5 | –4.72 | 38.0 | –4.50 | 36.5 | –4.27 |
O7 | 38.7 | –4.58 | 36.5 | –4.37 | 34.5 | –4.15 |
O8 | 36.4 | –4.30 | 35.0 | –4.15 | 32.5 | –4.01 |
O9 | 33.2 | –3.97 | 31.7 | –3.94 | 30.0 | –3.63 |
O9.5 | 32.1 | –3.82 | 30.5 | –3.75 | 29.0 | –3.48 |
B0 | 31.0 | –3.74 | 29.5 | –3.61 | 27.2 | –3.29 |
B0.5 | 29.6 | –3.57 | 28.5 | –3.47 | 24.3 | –2.89 |
B1 | 27.3 | –3.28 | 23.9 | –2.93 | 22.3 | –2.62 |
B1.5 | 26.1 | –3.17 | 22.5 | –2.70 | 20.6 | –2.38 |
B2 | 24.9 | –3.07 | 21.2 | –2.53 | 18.9 | –2.18 |
B2.5 | 23.9 | –2.95 | 19.8 | –2.34 | 17.2 | –1.78 |
B3 | 21.5 | –2.59 | 18.4 | –2.05 | 15.5 | –1.45 |
B5 | 16.7 | –1.65 | 13.8 | –1.04 | ||
B8 | 12.0 | –0.62 | ||||
B9 | 10.5 | –0.27 | ||||
A0 | 9.5 | +0.02 | ||||
A2 | 8.5 | +0.30 | ||||
A5 | 7.7 | +0.35 | ||||
A7 | 7.2 | +0.49 | ||||
F0 | 6.7 | +0.62 | ||||
F2 | 6.4 | +0.72 | ||||
F5 | 5.9 | +0.85 |
Spectral classification of the BLOeM sample permits coarse estimates of surface temperatures of normal OBAF stars presented in Table 1. These are adapted from the SMC spectral type-temperature calibration of Dufton et al. (2019), incorporating results for SMC O stars from Bouret et al. (2013), Bouret et al. (2021) and SMC late B and AF supergiants from Evans & Howarth (2003).
Bolometric luminosities are estimated from Ks band point spread function (PSF) photometry fitting from VISTA/VMC (Cioni et al., 2011), a distance modulus of 18.95 mag for the SMC (Hilditch et al., 2005; Graczyk et al., 2020), intrinsic V-Ks colours from Martins & Plez (2006), Cox (2000) and Koornneef (1983), plus V-band bolometric corrections drawn from Bouret et al. (2013) and Bouret et al. (2021) for O stars, Lanz & Hubeny (2007) for B stars with effective temperatures 15kK, Kudritzki et al. (2008) for late B and A supergiants (8kK 12kK) and Cox (2000) otherwise. The use of Ks band photometry avoids the need to correct for interstellar extinction, which is significantly smaller at the Ks band than at the band. Given this and the typically negligible extinction towards SMC sightlines (see, e.g., Schootemeijer et al., 2021), we assume here; this may affect the luminosities of individual objects, but does not affect our general conclusions.
We present the resulting Hertzsprung–Russell diagram (HRD) of the BLOeM sample in Fig. 10, together with evolutionary tracks and isochrones for SMC metallicity stars from the extended grid of Schootemeijer et al. (2019). As described in Appendix B of Hastings et al. (2021), these models have efficient semiconvection () and mass-dependent overshooting ( linearly increases from 0.1 at to 0.3 at 20 , and remains constant at 0.3 for higher masses). Apart from and , this grid has the same physics assumptions as Brott et al. (2011). We adopt the dwarf temperature calibration for sub-giants, and the giant temperature calibration for bright giants. The BLOeM sample includes dwarfs and sub-giants in the mass range 10–70 , plus a few giants and supergiants down to 7–8 . The three sgB[e] objects (BLOeM 2-116, BLOeM 3-012, and BLOeM 4-055) are excluded from the HRD, since standard calibrations cannot be used on them.
The errors on and can only be roughly estimated. The spectral types can be estimated to within one spectral bin roughly, corresponding to ; this is larger than typical calibration errors. However, the impact of binary contamination is not possible to quantify without further analysis; we set the error in Fig. 10 to 0.03 dex. Similarly, errors on bolometric correction can be estimated from neighbouring spectral-type bins. Depending on the spectral type, they typically range between , corresponding roughly to dex. Added to this are potential sources contaminating the Ks magnitudes (e.g., infrared excess), binary contamination, and reddening. To remain conservative, we adopt an uncertainty of dex in Fig. 10.
The HRD reveals the span of parameters covered by our sample in terms of initial mass and age. Qualitatively, it is similar to that presented by Humphreys & McElroy (1984), though the BLOeM sample goes substantially deeper in magnitude (mag). This HRD suggests that the sample probes the range , though only a minority of targets exceed initial masses . In terms of ages, a visual comparison with isochrones implies that the BLOeM sample probes stars as young as Myr, extending to ages of the order of 20 Myr or more. The lack of very massive stars in the SMC was noted in the past (e.g., Ramachandran et al., 2019; Schootemeijer et al., 2021), and could be related to a peak of star formation Myr ago (Antoniou et al., 2010; Rubele et al., 2015; Schootemeijer et al., 2021). The Humphreys–Davidson (H-D) limit (Humphreys & Davidson, 1979; Humphreys & McElroy, 1984), which marks the absence of bright stars in the upper-right part of the HRD, is clearly seen above , in agreement with recent evaluations of the H–D limit in the SMC (Ramachandran et al., 2019; Davies et al., 2018; Gilkis et al., 2021; Sabhahit et al., 2021).
We note that these estimates are subject to uncertainty given the uncertainties on , and could also change depending on the set of evolutionary tracks being used (e.g., Georgy et al., 2013; Choi et al., 2016; Marigo et al., 2017; Keszthelyi et al., 2022). Moreover, an unknown fraction of the sample stars will be affected either by past binary interactions or bright companions, while we only make use of single-star tracks in this first characterisation effort. A full analysis of the mass and age distribution will require treatment of multiplicity (e.g., binary identification, orbital analysis, spectral disentangling), which will be the subject of subsequent papers.
6 Conclusions
This work presented the rationale, target selection, and first characterisation of the sample spectroscopically monitored with FLAMES/VLT in the framework of the Binarity at LOw Metallicity (BLOeM) ESO Large Programme. BLOeM will collect 25 epochs of spectroscopy of 929 massive stars in the low metallicity conditions of the SMC in the period from October 2023 to September 2025. The sample populates eight fields within the SMC, probing several of its environments, though limited to field stars.
The goals of the survey are to use the time-dependent RVs of all targets to derive the observed and intrinsic binary fraction as a function of stellar mass and age, derive the orbits of all identified binaries, and, through this, establish fundamental properties such as the initial mass function, star formation history, surface abundance pattern, and orbital-parameter distributions of massive stars at low metallicity. Moreover, the survey will enable the identification of unique evolved binaries such as dormant OB+BH binaries, and provide testbeds for single-star evolution models via dynamical-mass measurements of eclipsing binaries.
This study was based on 9/25 epochs thus far acquired in the framework of the BLOeM survey during the first semester, with the main goal of describing the sample selection (Sect. 2), developing a data-reduction pipeline (Sect. 3), performing a spectral classification (Sect. 4), and investigating the mass and age domains probed by the sample (Sect. 5).
The sample spans a wide range of spectral types, extending from O4 for the earliest subtypes to F5 for the latest ones, though it is dominated by B0-2 stars. In total, there are 159 O-type stars, 324 early B-type (B0–B3) dwarfs, subgiants, and giants, 309 early B-type bright giants and supergiants, and 137 late-type supergiants. From spectral-type calibrations and usage of Ks magnitudes, we derived the effective temperatures and luminosities of the targets, assuming they are all single stars. The sample covers the regime kK and . From comparison to evolution tracks and isochrones extended from Schootemeijer et al. (2019), this roughly corresponds to initial masses in the range and ages mainly in the range Myr. The sample corroborates the blue region of the Humphreys-Davidson limit.
We highlighted a few peculiar objects in our sample: eight sources are confirmed as X-ray bright, four of which have been classified as HMXBs in the past, and one of which is a promising CWB candidate. Three objects are classified as sgB[e]/LBV-like stars, and two as Be + bloated stripped-star binary candidates. Two candidate magnetic stars, classified as Of?p stars previously, are also included in the sample.
We identify 75 stars as Oe/Be stars. Since we lack the diagnostic H line, this should be considered a lower limit, although usage of low resolution Gaia spectra which cover the H line only yielded a few more candidates, amounting to an Oe/Be fraction of relative to the non-supergiant OB sample. The Oe fraction among all O stars in the sample is 13%. However, the Be fraction is merely 10%, which is lower than reported in the literature. It is likely that we are biased against Be stars in this sample, since these tend to be redder compared to a ”standard” evolutionary track, which we used to select our sample (see Sect. 5.2). The Oe sample however is likely unbiased, and so this fraction should represent the Oe fraction in the SMC.
Finally, 74 eclipsing binaries are identified, as well as eight ellipsoidal variables (Sect. 5.3 and Appendix B).
A systematic analysis of the first batch of data for RV variability, for the various subsamples of the study, is currently underway.
Acknowledgements.
[…]
- Abbott et al. (2016) Abbott, B. P., Abbott, R., Abbott, T. D., et al. 2016, Phys. Rev. Lett., 116, 061102
- Abdul-Masih et al. (2020) Abdul-Masih, M., Banyard, G., Bodensteiner, J., et al. 2020, Nature, 580, E11
- Abt (1983) Abt, H. A. 1983, ARA&A, 21, 343
- Almeida et al. (2017) Almeida, L. A., Sana, H., Taylor, W., et al. 2017, A&A, 598, A84
- Antoniou et al. (2010) Antoniou, V., Zezas, A., Hatzidimitriou, D., & Kalogera, V. 2010, ApJ, 716, L140
- Ardeberg & Maurice (1977) Ardeberg, A. & Maurice, E. 1977, A&AS, 30, 261
- Atri et al. (2019) Atri, P., Miller-Jones, J. C. A., Bahramian, A., et al. 2019, MNRAS, 489, 3116
- Azzopardi & Vigneau (1979) Azzopardi, M. & Vigneau, J. 1979, A&AS, 35, 353
- Azzopardi & Vigneau (1982) Azzopardi, M. & Vigneau, J. 1982, A&AS, 50, 291
- Banagiri et al. (2023) Banagiri, S., Doctor, Z., Kalogera, V., Kimball, C., & Andrews, J. J. 2023, ApJ, 959, 106
- Banyard et al. (2022) Banyard, G., Sana, H., Mahy, L., et al. 2022, A&A, 658, A69
- Barkat et al. (1967) Barkat, Z., Rakavy, G., & Sack, N. 1967, Phys. Rev. Lett., 18, 379
- Bodensteiner et al. (2021) Bodensteiner, J., Sana, H., Wang, C., et al. 2021, A&A, 652, A70
- Bodensteiner et al. (2020) Bodensteiner, J., Shenar, T., & Sana, H. 2020, A&A, 641, A42
- Bonanos et al. (2010) Bonanos, A. Z., Lennon, D. J., Köhlinger, F., et al. 2010, AJ, 140, 416
- Bordier et al. (2023) Bordier, E., de Wit, W. J., Frost, A. J., et al. 2023, arXiv e-prints, arXiv:2311.06131
- Bouret et al. (2013) Bouret, J. C., Lanz, T., Martins, F., et al. 2013, A&A, 555, A1
- Bouret et al. (2021) Bouret, J. C., Martins, F., Hillier, D. J., et al. 2021, A&A, 647, A134
- Bowman (2024) Bowman, D. 2024, Astronomy and Geophysics, 65, 2.20
- Bowman et al. (2019) Bowman, D. M., Burssens, S., Pedersen, M. G., et al. 2019, Nature Astronomy, 3, 760
- Breivik et al. (2017) Breivik, K., Chatterjee, S., & Larson, S. L. 2017, ApJ, 850, L13
- Britavskiy et al. (2023) Britavskiy, N., Simón-Díaz, S., Holgado, G., et al. 2023, A&A, 672, A22
- Broekgaarden et al. (2021) Broekgaarden, F. S., Berger, E., Neijssel, C. J., et al. 2021, MNRAS, 508, 5028
- Brott et al. (2011) Brott, I., de Mink, S. E., Cantiello, M., et al. 2011, A&A, 530, A115
- Chen et al. (2017) Chen, T.-W., Smartt, S. J., Yates, R. M., et al. 2017, MNRAS, 470, 3566
- Choi et al. (2016) Choi, J., Dotter, A., Conroy, C., et al. 2016, ApJ, 823, 102
- Cioni et al. (2011) Cioni, M. R., Clementini, G., Girardi, L., et al. 2011, The Messenger, 144, 25
- Clark et al. (2013) Clark, J. S., Bartlett, E. S., Coe, M. J., et al. 2013, A&A, 560, A10
- Coe & Kirk (2015) Coe, M. J. & Kirk, J. 2015, MNRAS, 452, 969
- Corcoran et al. (1996) Corcoran, M. F., Stevens, I. R., Pollock, A. M. T., et al. 1996, ApJ, 464, 434
- Corral-Santana et al. (2016) Corral-Santana, J. M., Casares, J., Muñoz-Darias, T., et al. 2016, A&A, 587, A61
- Cox (2000) Cox, A. N. 2000, Allen’s astrophysical quantities
- Dachs (1970) Dachs, J. 1970, A&A, 9, 95
- Davies et al. (2018) Davies, B., Crowther, P. A., & Beasor, E. R. 2018, MNRAS, 478, 3138
- De Angeli et al. (2023) De Angeli, F., Weiler, M., Montegriffo, P., et al. 2023, A&A, 674, A2
- de Mink et al. (2013) de Mink, S. E., Langer, N., Izzard, R. G., Sana, H., & de Koter, A. 2013, ApJ, 764, 166
- de Mink & Mandel (2016) de Mink, S. E. & Mandel, I. 2016, MNRAS, 460, 3545
- de Mink et al. (2014) de Mink, S. E., Sana, H., Langer, N., Izzard, R. G., & Schneider, F. R. N. 2014, ApJ, 782, 7
- Dotter (2016) Dotter, A. 2016, ApJS, 222, 8
- Drout et al. (2023) Drout, M. R., Götberg, Y., Ludwig, B. A., et al. 2023, Science, 382, 1287
- Dufton et al. (2019) Dufton, P. L., Evans, C. J., Hunter, I., Lennon, D. J., & Schneider, F. R. N. 2019, A&A, 626, A50
- Dunstall et al. (2015) Dunstall, P. R., Dufton, P. L., Sana, H., et al. 2015, A&A, 580, A93
- El-Badry (2024) El-Badry, K. 2024, The Open Journal of Astrophysics, 7, 38
- El-Badry & Quataert (2021) El-Badry, K. & Quataert, E. 2021, MNRAS, 502, 3436
- El-Badry et al. (2023a) El-Badry, K., Rix, H.-W., Cendes, Y., et al. 2023a, MNRAS, 521, 4323
- El-Badry et al. (2023b) El-Badry, K., Rix, H.-W., Quataert, E., et al. 2023b, MNRAS, 518, 1057
- Evans et al. (2019) Evans, C. J., Castro, N., Gonzalez, O. A., et al. 2019, A&A, 622, A129
- Evans & Howarth (2003) Evans, C. J. & Howarth, I. D. 2003, MNRAS, 345, 1223
- Evans et al. (2004) Evans, C. J., Howarth, I. D., Irwin, M. J., Burnley, A. W., & Harries, T. J. 2004, MNRAS, 353, 601
- Evans et al. (2006) Evans, C. J., Lennon, D. J., Smartt, S. J., & Trundle, C. 2006, A&A, 456, 623
- Evans et al. (2011) Evans, C. J., Taylor, W. D., Hénault-Brunet, V., et al. 2011, A&A, 530, A108
- Evans et al. (2010) Evans, I. N., Primini, F. A., Glotfelty, K. J., et al. 2010, ApJS, 189, 37
- Farmer et al. (2019) Farmer, R., Renzo, M., de Mink, S. E., Marchant, P., & Justham, S. 2019, ApJ, 887, 53
- Ferrario et al. (2009) Ferrario, L., Pringle, J. E., Tout, C. A., & Wickramasinghe, D. T. 2009, MNRAS, 400, L71
- Foellmi et al. (2003) Foellmi, C., Moffat, A. F. J., & Guerrero, M. A. 2003, MNRAS, 338, 360
- Frost et al. (2022) Frost, A. J., Bodensteiner, J., Rivinius, T., et al. 2022, A&A, 659, L3
- Frost et al. (2024) Frost, A. J., Sana, H., Mahy, L., et al. 2024, Science, 384, 214
- Fryer et al. (2001) Fryer, C. L., Woosley, S. E., & Heger, A. 2001, ApJ, 550, 372
- Gaia Collaboration et al. (2024) Gaia Collaboration, Panuzzo, P., Mazeh, T., et al. 2024, A&A, 686, L2
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Gal-Yam (2012) Gal-Yam, A. 2012, Science, 337, 927
- Garmany et al. (1987) Garmany, C. D., Conti, P. S., & Massey, P. 1987, AJ, 93, 1070
- Georgy et al. (2013) Georgy, C., Ekström, S., Eggenberger, P., et al. 2013, A&A, 558, A103
- Giacobbo et al. (2018) Giacobbo, N., Mapelli, M., & Spera, M. 2018, MNRAS, 474, 2959
- Giesers et al. (2018) Giesers, B., Dreizler, S., Husser, T.-O., et al. 2018, MNRAS, 475, L15
- Gilkis et al. (2021) Gilkis, A., Shenar, T., Ramachandran, V., et al. 2021, arXiv e-prints, arXiv:2102.03102
- Götberg et al. (2018) Götberg, Y., de Mink, S. E., Groh, J. H., et al. 2018, A&A, 615, A78
- Graczyk et al. (2020) Graczyk, D., Pietrzyński, G., Thompson, I. B., et al. 2020, ApJ, 904, 13
- Graham & Fruchter (2017) Graham, J. F. & Fruchter, A. S. 2017, ApJ, 834, 170
- Graus et al. (2012) Graus, A. S., Lamb, J. B., & Oey, M. S. 2012, ApJ, 759, 10
- Grunhut et al. (2017) Grunhut, J. H., Wade, G. A., Neiner, C., et al. 2017, MNRAS, 465, 2432
- Haberl & Sturm (2016) Haberl, F. & Sturm, R. 2016, A&A, 586, A81
- Hainich et al. (2015) Hainich, R., Pasemann, D., Todt, H., et al. 2015, A&A, 581, A21
- Harmanec (1988) Harmanec, P. 1988, Bulletin of the Astronomical Institutes of Czechoslovakia, 39, 329
- Hastings et al. (2021) Hastings, B., Langer, N., Wang, C., Schootemeijer, A., & Milone, A. P. 2021, A&A, 653, A144
- Hilditch et al. (2005) Hilditch, R. W., Howarth, I. D., & Harries, T. J. 2005, MNRAS, 357, 304
- Hubeny & Lanz (1995) Hubeny, I. & Lanz, T. 1995, ApJ, 439, 875
- Humphreys (1983) Humphreys, R. M. 1983, ApJ, 265, 176
- Humphreys & Davidson (1979) Humphreys, R. M. & Davidson, K. 1979, ApJ, 232, 409
- Humphreys et al. (1991) Humphreys, R. M., Kudritzki, R. P., & Groth, H. G. 1991, A&A, 245, 593
- Humphreys & McElroy (1984) Humphreys, R. M. & McElroy, D. B. 1984, ApJ, 284, 565
- Hunter et al. (2007) Hunter, I., Dufton, P. L., Smartt, S. J., et al. 2007, A&A, 466, 277
- Irrgang et al. (2020) Irrgang, A., Geier, S., Kreuzer, S., Pelisoli, I., & Heber, U. 2020, A&A, 633, L5
- Janssens et al. (2022) Janssens, S., Shenar, T., Sana, H., et al. 2022, A&A, 658, A129
- Janssens et al. (2023) Janssens, S., Shenar, T., Sana, H., & Marchant, P. 2023, A&A, 670, A79
- Jermyn et al. (2023) Jermyn, A. S., Bauer, E. B., Schwab, J., et al. 2023, ApJS, 265, 15
- Keszthelyi et al. (2022) Keszthelyi, Z., de Koter, A., Götberg, Y., et al. 2022, MNRAS, 517, 2028
- Kiminki & Kobulnicky (2012) Kiminki, D. C. & Kobulnicky, H. A. 2012, ApJ, 751, 4
- Klencki et al. (2018) Klencki, J., Moe, M., Gladysz, W., et al. 2018, A&A, 619, A77
- Kobulnicky & Fryer (2007) Kobulnicky, H. A. & Fryer, C. L. 2007, ApJ, 670, 747
- Koornneef (1983) Koornneef, J. 1983, A&A, 128, 84
- Kraus (2019) Kraus, M. 2019, Galaxies, 7, 83
- Kroupa (2001) Kroupa, P. 2001, MNRAS, 322, 231
- Kudritzki et al. (2008) Kudritzki, R.-P., Urbaneja, M. A., Bresolin, F., et al. 2008, ApJ, 681, 269
- Kummer et al. (2023) Kummer, F., Toonen, S., & de Koter, A. 2023, A&A, 678, A60
- Lamb et al. (2013) Lamb, J. B., Oey, M. S., Graus, A. S., Adams, F. C., & Segura-Cox, D. M. 2013, ApJ, 763, 101
- Lamb et al. (2016) Lamb, J. B., Oey, M. S., Segura-Cox, D. M., et al. 2016, ApJ, 817, 113
- Lamers et al. (1998) Lamers, H. J. G. L. M., Zickgraf, F.-J., de Winter, D., Houziaux, L., & Zorec, J. 1998, A&A, 340, 117
- Langer (2012) Langer, N. 2012, ARA&A, 50, 107
- Langer et al. (2020) Langer, N., Baade, D., Bodensteiner, J., et al. 2020, A&A, 633, A40
- Langer et al. (2007) Langer, N., Norman, C. A., de Koter, A., et al. 2007, A&A, 475, L19
- Lanz & Hubeny (2003) Lanz, T. & Hubeny, I. 2003, ApJS, 146, 417
- Lanz & Hubeny (2007) Lanz, T. & Hubeny, I. 2007, ApJS, 169, 83
- Laycock et al. (2010) Laycock, S., Zezas, A., Hong, J., Drake, J. J., & Antoniou, V. 2010, ApJ, 716, 1217
- Lennon (1997) Lennon, D. J. 1997, A&A, 317, 871
- Lorenzo et al. (2022) Lorenzo, M., Garcia, M., Najarro, F., et al. 2022, MNRAS, 516, 4164
- Machida (2008) Machida, M. N. 2008, ApJ, 682, L1
- Mahy et al. (2020a) Mahy, L., Almeida, L. A., Sana, H., et al. 2020a, A&A, 634, A119
- Mahy et al. (2020b) Mahy, L., Sana, H., Abdul-Masih, M., et al. 2020b, A&A, 634, A118
- Mahy et al. (2022) Mahy, L., Sana, H., Shenar, T., et al. 2022, A&A, 664, A159
- Maíz Apellániz et al. (2016) Maíz Apellániz, J., Sota, A., Arias, J. I., et al. 2016, ApJS, 224, 4
- Mandel & Broekgaarden (2022) Mandel, I. & Broekgaarden, F. S. 2022, Living Reviews in Relativity, 25, 1
- Mandel & Farmer (2022) Mandel, I. & Farmer, A. 2022, Phys. Rep, 955, 1
- Maravelias et al. (2014) Maravelias, G., Zezas, A., Antoniou, V., & Hatzidimitriou, D. 2014, MNRAS, 438, 2005
- Marchant & Bodensteiner (2023) Marchant, P. & Bodensteiner, J. 2023, arXiv e-prints, arXiv:2311.01865
- Marchant et al. (2016) Marchant, P., Langer, N., Podsiadlowski, P., Tauris, T. M., & Moriya, T. J. 2016, A&A, 588, A50
- Marigo et al. (2017) Marigo, P., Girardi, L., Bressan, A., et al. 2017, ApJ, 835, 77
- Marín Pina et al. (2024) Marín Pina, D., Rastello, S., Gieles, M., et al. 2024, arXiv e-prints, arXiv:2404.13036
- Marks et al. (2012) Marks, M., Kroupa, P., Dabringhausen, J., & Pawlowski, M. S. 2012, MNRAS, 422, 2246
- Martayan et al. (2007) Martayan, C., Frémat, Y., Hubert, A. M., et al. 2007, A&A, 462, 683
- Martins & Plez (2006) Martins, F. & Plez, B. 2006, A&A, 457, 637
- Martins et al. (2005) Martins, F., Schaerer, D., & Hillier, D. J. 2005, A&A, 436, 1049
- Mashian & Loeb (2017) Mashian, N. & Loeb, A. 2017, MNRAS, 470, 2611
- Mason et al. (2009) Mason, B. D., Hartkopf, W. I., Gies, D. R., Henry, T. J., & Helsel, J. W. 2009, AJ, 137, 3358
- Massey (2002) Massey, P. 2002, ApJS, 141, 81
- Massey et al. (2004) Massey, P., Bresolin, F., Kudritzki, R. P., Puls, J., & Pauldrach, A. W. A. 2004, ApJ, 608, 1001
- Massey et al. (1995) Massey, P., Lang, C. C., Degioia-Eastwood, K., & Garmany, C. D. 1995, ApJ, 438, 188
- Massey et al. (2000) Massey, P., Waterhouse, E., & DeGioia-Eastwood, K. 2000, AJ, 119, 2214
- Massey et al. (2009) Massey, P., Zangari, A. M., Morrell, N. I., et al. 2009, ApJ, 692, 618
- McQuinn et al. (2015) McQuinn, K. B. W., Skillman, E. D., Dolphin, A., et al. 2015, ApJ, 812, 158
- Mennickent et al. (2006) Mennickent, R. E., Cidale, L., Pietrzyński, G., Gieren, W., & Sabogal, B. 2006, A&A, 457, 949
- Millward & Walker (1985) Millward, C. G. & Walker, G. A. H. 1985, ApJS, 57, 63
- Mirabel & Rodrigues (2003) Mirabel, I. F. & Rodrigues, I. 2003, Science, 300, 1119
- Modjaz et al. (2008) Modjaz, M., Kewley, L., Kirshner, R. P., et al. 2008, AJ, 135, 1136
- Moe & Di Stefano (2013) Moe, M. & Di Stefano, R. 2013, ApJ, 778, 95
- Moe et al. (2019) Moe, M., Kratter, K. M., & Badenes, C. 2019, ApJ, 875, 61
- Mokiem et al. (2006) Mokiem, M. R., de Koter, A., Evans, C. J., et al. 2006, A&A, 456, 1131
- Morrell et al. (2003) Morrell, N., Ostrov, P., Massey, P., & Gamen, R. 2003, MNRAS, 341, 583
- Nazé et al. (2007) Nazé, Y., Corcoran, M. F., Koenigsberger, G., & Moffat, A. F. J. 2007, ApJ, 658, L25
- Nazé et al. (2004) Nazé, Y., Manfroid, J., Stevens, I. R., Corcoran, M. F., & Flores, A. 2004, ApJ, 608, 208
- Negueruela et al. (2024) Negueruela, I., Simón-Díaz, S., de Burgos, A., Casasbuenas, A., & Beck, P. G. 2024, arXiv e-prints, arXiv:2407.04163
- Neugent et al. (2020) Neugent, K. F., Levesque, E. M., Massey, P., Morrell, N. I., & Drout, M. R. 2020, ApJ, 900, 118
- Neugent et al. (2018) Neugent, K. F., Massey, P., & Morrell, N. 2018, ApJ, 863, 181
- Neugent et al. (2010) Neugent, K. F., Massey, P., Skiff, B., et al. 2010, ApJ, 719, 1784
- Offner et al. (2023) Offner, S. S. R., Moe, M., Kratter, K. M., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 275
- Oskinova (2005) Oskinova, L. M. 2005, MNRAS, 361, 679
- Oskinova et al. (2013) Oskinova, L. M., Sun, W., Evans, C. J., et al. 2013, ApJ, 765, 73
- Paczyński (1967) Paczyński, B. 1967, Acta Astron., 17, 355
- Pasquali et al. (2000) Pasquali, A., Nota, A., Langer, N., Schulte-Ladbeck, R. E., & Clampin, M. 2000, AJ, 119, 1352
- Pasquini et al. (2002) Pasquini, L., Avila, G., Blecha, A., et al. 2002, The Messenger, 110, 1
- Patrick et al. (2019) Patrick, L. R., Lennon, D. J., Britavskiy, N., et al. 2019, A&A, 624, A129
- Paul et al. (2012) Paul, K. T., Subramaniam, A., Mathew, B., Mennickent, R. E., & Sabogal, B. 2012, MNRAS, 421, 3622
- Pauli et al. (2022a) Pauli, D., Langer, N., Aguilera-Dena, D. R., Wang, C., & Marchant, P. 2022a, A&A, 667, A58
- Pauli et al. (2022b) Pauli, D., Oskinova, L. M., Hamann, W. R., et al. 2022b, A&A, 659, A9
- Pawlak et al. (2016) Pawlak, M., Soszyński, I., Udalski, A., et al. 2016, Acta Astron., 66, 421
- Paxton et al. (2011) Paxton, B., Bildsten, L., Dotter, A., et al. 2011, ApJS, 192, 3
- Paxton et al. (2013) Paxton, B., Cantiello, M., Arras, P., et al. 2013, ApJS, 208, 4
- Paxton et al. (2015) Paxton, B., Marchant, P., Schwab, J., et al. 2015, ApJS, 220, 15
- Paxton et al. (2018) Paxton, B., Schwab, J., Bauer, E. B., et al. 2018, ApJS, 234, 34
- Paxton et al. (2019) Paxton, B., Smolec, R., Schwab, J., et al. 2019, ApJS, 243, 10
- Podsiadlowski et al. (2006) Podsiadlowski, P., Morris, T. S., & Ivanova, N. 2006, in Astronomical Society of the Pacific Conference Series, Vol. 355, Stars with the B[e] Phenomenon, ed. M. Kraus & A. S. Miroshnichenko, 259
- Pols et al. (1991) Pols, O., Cote, J., Waters, L. B. F. M., & Heise, J. 1991, A&A, 241, 419
- Portegies Zwart & Verbunt (1996) Portegies Zwart, S. F. & Verbunt, F. 1996, A&A, 309, 179
- Price-Whelan et al. (2020) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020, ApJ, 895, 2
- Prinja (1987) Prinja, R. K. 1987, MNRAS, 228, 173
- Quimby et al. (2011) Quimby, R. M., Kulkarni, S. R., Kasliwal, M. M., et al. 2011, Nature, 474, 487
- Ramachandran et al. (2019) Ramachandran, V., Hamann, W. R., Oskinova, L. M., et al. 2019, A&A, 625, A104
- Rastello et al. (2023) Rastello, S., Iorio, G., Mapelli, M., et al. 2023, MNRAS, 526, 740
- Renzo & Götberg (2021) Renzo, M. & Götberg, Y. 2021, ApJ, 923, 277
- Renzo et al. (2019) Renzo, M., Zapartas, E., de Mink, S. E., et al. 2019, A&A, 624, A66
- Ricker et al. (2015) Ricker, G. R., Winn, J. N., Vanderspek, R., et al. 2015, Journal of Astronomical Telescopes, Instruments, and Systems, 1, 014003
- Rivinius et al. (2013) Rivinius, T., Carciofi, A. C., & Martayan, C. 2013, A&A Rev., 21, 69
- Rubele et al. (2015) Rubele, S., Girardi, L., Kerber, L., et al. 2015, MNRAS, 449, 639
- Sabhahit et al. (2021) Sabhahit, G. N., Vink, J. S., Higgins, E. R., & Sander, A. A. C. 2021, MNRAS, 506, 4473
- Saigo et al. (2004) Saigo, K., Matsumoto, T., & Umemura, M. 2004, ApJ, 615, L65
- Sana et al. (2013) Sana, H., de Koter, A., de Mink, S. E., et al. 2013, A&A, 550, A107
- Sana et al. (2012) Sana, H., de Mink, S. E., de Koter, A., et al. 2012, Science, 337, 444
- Sana et al. (2014) Sana, H., Le Bouquin, J. B., Lacour, S., et al. 2014, ApJS, 215, 15
- Sana et al. (2006) Sana, H., Rauw, G., Nazé, Y., Gosset, E., & Vreux, J. M. 2006, MNRAS, 372, 661
- Sanduleak (1968) Sanduleak, N. 1968, AJ, 73, 246
- Schneider et al. (2019) Schneider, F. R. N., Ohlmann, S. T., Podsiadlowski, P., et al. 2019, Nature, 574, 211
- Schootemeijer et al. (2019) Schootemeijer, A., Langer, N., Grin, N. J., & Wang, C. 2019, A&A, 625, A132
- Schootemeijer et al. (2021) Schootemeijer, A., Langer, N., Lennon, D., et al. 2021, A&A, 646, A106
- Schootemeijer et al. (2022) Schootemeijer, A., Lennon, D. J., Garcia, M., et al. 2022, A&A, 667, A100
- Schootemeijer et al. (2024) Schootemeijer, A., Shenar, T., Langer, N., et al. 2024, arXiv e-prints, arXiv:2406.01420
- Shahaf et al. (2023) Shahaf, S., Bashi, D., Mazeh, T., et al. 2023, MNRAS, 518, 2991
- Sheets et al. (2013) Sheets, H. A., Bolatto, A. D., van Loon, J. T., et al. 2013, ApJ, 771, 111
- Shenar et al. (2020) Shenar, T., Bodensteiner, J., Abdul-Masih, M., et al. 2020, A&A, 639, L6
- Shenar et al. (2016) Shenar, T., Hainich, R., Todt, H., et al. 2016, A&A, 591, A22
- Shenar et al. (2022a) Shenar, T., Sana, H., Mahy, L., et al. 2022a, Nature Astronomy, 6, 1085
- Shenar et al. (2022b) Shenar, T., Sana, H., Mahy, L., et al. 2022b, A&A, 665, A148
- Shenar et al. (2023) Shenar, T., Wade, G. A., Marchant, P., et al. 2023, Science, 381, 761
- Shore et al. (1987) Shore, S. N., Sanduleak, N., & Allen, D. A. 1987, A&A, 176, 59
- Skillman et al. (1989) Skillman, E. D., Kennicutt, R. C., & Hodge, P. W. 1989, ApJ, 347, 875
- Smith Neubig & Bruhweiler (1999) Smith Neubig, M. M. & Bruhweiler, F. C. 1999, AJ, 117, 2856
- Sota et al. (2011) Sota, A., Maíz Apellániz, J., Walborn, N. R., et al. 2011, ApJS, 193, 24
- Stegmann et al. (2022) Stegmann, J., Antonini, F., & Moe, M. 2022, MNRAS, 516, 1406
- Sturm et al. (2013) Sturm, R., Haberl, F., Pietsch, W., et al. 2013, A&A, 558, A3
- Tauris et al. (2017) Tauris, T. M., Kramer, M., Freire, P. C. C., et al. 2017, ApJ, 846, 170
- Telford et al. (2023) Telford, O. G., McQuinn, K. B. W., Chisholm, J., & Berg, D. A. 2023, ApJ, 943, 65
- Testor (2001) Testor, G. 2001, A&A, 372, 667
- Testor & Lortet (1987) Testor, G. & Lortet, M. C. 1987, A&A, 178, 25
- Torres et al. (2010) Torres, G., Andersen, J., & Giménez, A. 2010, A&A Rev., 18, 67
- Traulsen et al. (2020) Traulsen, I., Schwope, A. D., Lamer, G., et al. 2020, A&A, 641, A137
- Udalski (2003) Udalski, A. 2003, Acta Astron., 53, 291
- Udalski et al. (2015) Udalski, A., Szymański, M. K., & Szymański, G. 2015, Acta Astron., 65, 1
- van Son et al. (2022) van Son, L. A. C., de Mink, S. E., Renzo, M., et al. 2022, ApJ, 940, 184
- Vigna-Gómez et al. (2024) Vigna-Gómez, A., Willcox, R., Tamborra, I., et al. 2024, Phys. Rev. Lett., 132, 191403
- Villaseñor et al. (2021) Villaseñor, J. I., Taylor, W. D., Evans, C. J., et al. 2021, MNRAS, 507, 5348
- Vink et al. (2023) Vink, J. S., Mehner, A., Crowther, P. A., et al. 2023, A&A, 675, A154
- Walborn (1972) Walborn, N. R. 1972, AJ, 77, 312
- Walborn (1983) Walborn, N. R. 1983, ApJ, 265, 716
- Walborn & Fitzpatrick (1990) Walborn, N. R. & Fitzpatrick, E. L. 1990, PASP, 102, 379
- Walborn et al. (2002) Walborn, N. R., Fullerton, A. W., Crowther, P. A., et al. 2002, ApJS, 141, 443
- Walborn et al. (2000) Walborn, N. R., Lennon, D. J., Heap, S. R., et al. 2000, PASP, 112, 1243
- Walborn et al. (2014) Walborn, N. R., Sana, H., Simón-Díaz, S., et al. 2014, A&A, 564, A40
- Wang et al. (2020) Wang, C., Langer, N., Schootemeijer, A., et al. 2020, ApJ, 888, L12
- Wang et al. (2017) Wang, L., Gies, D. R., & Peters, G. J. 2017, ApJ, 843, 60
- Wang et al. (2016) Wang, S., Liu, J., Qiu, Y., et al. 2016, ApJS, 224, 40
- Webb et al. (2020) Webb, N. A., Coriat, M., Traulsen, I., et al. 2020, A&A, 641, A136
- Weiler et al. (2023) Weiler, M., Carrasco, J. M., Fabricius, C., & Jordi, C. 2023, A&A, 671, A52
- Woosley (2017) Woosley, S. E. 2017, ApJ, 836, 244
- Woosley & Heger (2006) Woosley, S. E. & Heger, A. 2006, ApJ, 637, 914
- Wu et al. (2020) Wu, S., Everson, R. W., Schneider, F. R. N., Podsiadlowski, P., & Ramirez-Ruiz, E. 2020, ApJ, 901, 44
- Yang et al. (2019) Yang, M., Bonanos, A. Z., Jiang, B.-W., et al. 2019, A&A, 629, A91
- Yoon & Langer (2005) Yoon, S. C. & Langer, N. 2005, A&A, 443, 643
- Zasche et al. (2014) Zasche, P., Wolf, M., Vraštil, J., et al. 2014, A&A, 572, A71
- Zickgraf et al. (1996) Zickgraf, F. J., Kovacs, J., Wolf, B., et al. 1996, A&A, 309, 505
- Zivick et al. (2018) Zivick, P., Kallivayalil, N., van der Marel, R. P., et al. 2018, ApJ, 864, 55
Appendix A Epoch dates and spectral types
Field 1 | Field 2 | Field 3 | Field 4 | Field 5 | Field 6 | Field 7 | Field 8 | |
RA center | 01:02:53.8 | 00:52:01.0 | 00:48:20.2 | 00:59:41.8 | 01:05:56.2 | 01:15:08.2 | 00:59:51.4 | 00:52:44.2 |
DEC center | -72:05:26.1 | -72:41:26.1 | -73:16:14.1 | -72:11:26.1 | -72:19:50.1 | -73:15:02.1 | -72:37:50.1 | -72:15:02.1 |
Epoch 1 [MJD] | 60219.10 | 60242.07 | 60242.10 | 60242.12 | 60242.15 | 60261.16 | 60246.12 | 60246.14 |
Epoch 2 [MJD] | 60220.26 | 60247.19 | 60247.21 | 60247.24 | 60247.26 | 60267.20 | 60248.17 | 60261.18 |
Epoch 3 [MJD] | 60242.05 | 60256.20 | 60256.24 | 60254.35 | 60256.22 | 60280.02 | 60256.18 | 60267.06 |
Epoch 4 [MJD] | 60245.02 | 60261.09 | 60261.11 | 60256.27 | 60261.03 | 60282.12 | 60261.07 | 60268.11 |
Epoch 5 [MJD] | 60248.15 | 60267.15 | 60262.13 | 60261.14 | 60262.11 | 60285.09 | 60267.04 | 60280.12 |
Epoch 6 [MJD] | 60252.21 | 60281.10 | 60267.09 | 60267.12 | 60267.17 | 60286.22 | 60268.13 | 60280.14 |
Epoch 7 [MJD] | 60256.29 | 60285.16 | 60270.03 | 60281.05 | 60270.05 | 60288.14 | 60280.16 | 60282.10 |
Epoch 8 [MJD] | 60261.05 | 60289.05 | 60281.08 | 60285.18 | 60280.04 | 60290.05 | 60286.19 | 60285.11 |
Epoch 9 [MJD] | 60285.20 | 60290.09 | 60285.13 | 60289.07 | 60281.12 | 60291.09 | 60288.16 | 60288.19 |
Table 2 provides the field centres and MJD values of the mid-exposures of the combined two sub-exposures per field and epoch. Table LABEL:tab:SpT provides Gaia DR3 IDs, common aliases, Ks magnitudes from the VISTA/VMC catalogue (Cioni et al. 2011), previous spectral-type classifications, newly derived spectral types, flags for H-emitters in low-resolution Gaia spectra (Sect. 5.2), and additional comments. Newly derived spectral types are based on the BLOeM dataset unless otherwise stated in the comments.
ID | Gaia DR3 ID | Alias | K[mag] | SpT (old) | SpT (new) | H emit. | Comment |
1-001 | 4690503998385774848 | AzV 76F | 13.69 | B8 [Azz79] | B9 Iab | no | - |
1-002 | 4690519082313236608 | [M2002] SMC 49825 | 15.85 | B0.2 III [Lam16] | B0 IV: | no | - |
1-003 | 4690519151032663296 | - | 15.52 | - | B2.5 II: | no | - |
1-004 | 4690501146531455744 | Sk 92, AzV 260 | 13.57 | B1.5 Iab [Len97] | B1 Ib | no | - |
1-005 | 4690505510214025472 | [M2002] SMC 50031 | 15.45 | B0 [She13] | B1 II | no | - |
1-006 | 4690520010026062336 | OGLE SMC725.15.018759 | 15.12 | - | O9.5 V + early B | no | - |
1-007 | 4690506781524350336 | - | 16.31 | - | B1.5: II: | no | - |
1-008 | 4690501180887206656 | Sk 93, AzV 263 | 12.79 | B9 Ia [Len97] | B9 Iab | no | - |
1-009 | 4690520387983090432 | Sk 94, AzV 264 | 12.78 | B1 Ia [Len97] | B1 Ia | no | - |
1-010 | 4690505883854563840 | - | 15.27 | - | B1.5 III: | no | - |
1-011 | 4690506712804881792 | OGLE SMC108.3 21525 | 14.50 | - | B1.5 II | no | - |
1-012 | 4690507022042458624 | AzV 267 | 15.65 | O8 Vn [Mas02] | O7.5 Vn | no | LC from [Mas02] |
1-013 | 4690506747164575872 | - | 16.66 | - | B1.5 III: | no | - |
1-014 | 4690520074427341184 | - | 15.20 | - | B1 II | no | - |
1-015 | 4690505986961967872 | - | 15.74 | - | B0.5 III | no | - |
1-016 | 4690506914646214912 | Sk 99, AzV 273 | 11.78 | A0 I [Neu10] | A2 Ib | no | - |
1-017 | 4690507193841111296 | - | 15.71 | - | B1 II: | no | - |
1-018 | 4690506025610129152 | LHA 115-S 34, AzV 276 | 14.32 | B0e [Azz75] | B2 II: e | yes | - |
1-019 | 4690501558844407936 | AzV 277 | 14.47 | B1-2 (II) [Eva04] | B2.5 Ib | no | - |
1-020 | 4690519047953407232 | AzV 279 | 14.77 | B0 III [Gar87] | B0 III | no | - |
1-021 | 4690519425910500352 | AzV 280 | 14.16 | B pec [Gar87] | B1 Ib | no | - |
1-022 | 4690519700788344064 | - | 17.07 | - | B1 III | no | - |
1-023 | 4690502452197510272 | AzV 282 | 15.42 | O7 V [Eva04] | O7.5V(n) | no | LC from [Eva04] |
1-024 | 4690506437926915840 | 2dFS 1609 | 15.76 | O7 II(f) [Eva04] | O8 II(f) | no | LC from [Eva04] |
1-025 | 4690502452197478528 | Sk 101, AzV 287 | 13.61 | O9.5 I [Wal83] | O9.2 Ib(n) | no | - |
1-026 | 4690519455996745216 | - | 15.09 | - | B1 Ib | no | - |
1-027 | 4690501661923520768 | AzV 296 | 15.04 | O8.5 V((f)) [Mas04] | O7.5 V((f))n | no | LC from [Mas04] |
1-028 | 4690519769509887488 | RMC 23, AzV 297 | 12.21 | B8 Ia [Len97] | B8 Iab/Ia | no | - |
1-029 | 4690519803867505024 | - | 15.35 | - | B2 II | no | - |
1-030 | 4690506266128220160 | - | 17.06 | - | B1 II | no | - |
1-031 | 4690507670561677696 | RMC 24, AzV 298 | 12.33 | A0 I [Neu18] | A0 Iab | no | - |
1-032 | 4690506541006084224 | 2dFS 5105 | 15.96 | B0 (V) [Eva04] | B1.5 III: | no | - |
1-033 | 4690502860220750208 | AzV 300 | 15.18 | B0 IIw [Gar87] | O9.5 V:n | no | - |
1-034 | 4690519803867499648 | AzV 301 | 14.98 | B1 V [Mas95] | B1 II | no | - |
1-035 | 4690507709237145856 | [M2002] SMC 53417 | 15.39 | - | B1.5 III-II | no | - |
1-036 | 4690519593389669632 | RMC 25, AzV 303 | 13.26 | B1.5 Iab [Len97] | B1.5 Ib | no | - |
1-037 | 4690506300487960576 | OGLE SMC113.6 5437 | 16.20 | - | B1.5 III | no | - |
1-038 | 4690513653473762176 | - | 14.15 | - | B2.5 Ib | no | - |
1-039 | 4690503306873686912 | Dachs SMC 2-2 | 12.40 | A3 I [Neu10] | A2 II/Ib | no | - |
1-040 | 4690525438865707904 | LIN 393 | 13.23 | - | O9.7 III:n e | yes | - |
1-041 | 4690512485242560000 | [M2002] SMC 54325 | 15.03 | - | O9.7: V:n + O9.7: | no | - |
1-042 | 4690512485242543104 | - | 13.83 | - | B2 Ib | no | - |
1-043 | 4690526641456527360 | AzV 311 | 15.24 | O7-9 [She13] | O9.7 III: + O9.7 | no | - |
1-044 | 4690512485242552960 | [M2002] SMC 54432 | 15.75 | - | B1 II | no | - |
1-045 | 4690501971160817792 | AzV 312 | 14.47 | B0 (III) [Eva04] | B0.7 II | no | - |
1-046 | 4687499475466452864 | - | 15.95 | - | B1 III | no | - |
1-047 | 4690512279084139392 | - | 15.10 | - | B1 III-II | no | - |
1-048 | 4690513687833493760 | AzV 313 | 15.10 | B1 [Azz75] | B2 III-II | no | - |
1-049 | 4690503375593898624 | - | 14.67 | - | B1.5 II: + early B | no | - |
1-050 | 4690503036312684544 | 2dFS 1707 | 12.64 | F0 [Eva04] | A7 Ib | no | - |
1-051 | 4690502280398446720 | RMC 27, Sk 106, AzV 315 | 10.56 | A0 Ia [Len97] | A0 Ia | no | - |
1-052 | 4690512588321738496 | - | 15.31 | - | B1 II | no | - |
1-053 | 4690502280398443776 | AzV 318 | 14.09 | B2 I [Gar87] | B1 Ib | no | - |
1-054 | 4690512584031941504 | [M2002] SMC 54931 | 14.92 | - | B0 II | no | - |
1-055 | 4690513516034604032 | - | 15.57 | - | B1.5 III-II | no | - |
1-056 | 4690508980549304704 | AzV 321, 2dFS 1720 | 14.30 | O9.5 Ib [Eva04] | O9.5 Ibn | no | LC from [Eva04] |
1-057 | 4690513344235943040 | - | 15.08 | - | B1.5 II: | no | - |
1-058 | 4690512313443855360 | LIN 404 | 13.31 | - | O9.5 II: pe | yes | - |
1-059 | 4690508151597375616 | [M2002] SMC 55168 | 15.86 | - | B0 IV | no | - |
1-060 | 4690508907539818112 | OGLE J010302.31-720836.1 | 14.12 | - | B1.5 Ib | no | - |
1-061 | 4690512244724379264 | OGLE SMC-SC9 175326 | 14.30 | B1-2 (II) [Eva04] | B1.5 Ib: | no | - |
1-062 | 4690507881037706112 | Dachs SMC 2-8, AzV 324 | 13.09 | B4 Iab: [Smi97] | B8 Iab | no | - |
1-063 | 4690509289786901504 | Cl* NGC 371 KAG V12 | 16.94 | - | B1.5 II: e | no | - |
1-064 | 4690508087196130304 | 2dFS 1734 | 14.66 | B9 (II) [Eva04] | B9 Ib | no | - |
1-065 | 4687505389671548672 | AzV 90F | 12.45 | F0 Ib [Neu10] | A7 Ib | no | - |
1-066 | 4690512691400904448 | RMC 28, AzV 327 | 13.81 | O9.5 II-Ibw [Wal00] | O9.7 II-Ib | no | - |
1-067 | 4690508946189561728 | [M2002] SMC 55556 | 15.48 | - | B0.7 III | no | - |
1-068 | 4690527088132999552 | [M2002] SMC 55948 | 14.72 | OBe? [She13] | O9.7 III: | no | - |
1-069 | 4690514134509862528 | [M2002] SMC 55952 | 15.15 | B0.2 V [Lam16] | B0.7 III | no | - |
1-070 | 4690512347803574400 | [MWD2000] h53-40 | 14.83 | B2 I [Mas00] | B1.5 II | no | - |
1-071 | 4687505320952170880 | - | 16.17 | - | B0.7 III | no | - |
1-072 | 4690509324146632704 | MOA J010321.3-720538 | 14.27 | O6 V +O4-5 III(f) [Mor03] | O5 V(n) + O6.5(n) | no | - |
1-073 | 4690526057340871040 | [MA93] 1373 | 14.68 | - | B1 II: e | yes | - |
1-074 | 4690509117967865088 | - | 16.54 | - | B1 II | no | - |
1-075 | 4690526057340870272 | AzV 330 | 13.02 | Be [Mas95] | O9.2 II:n pe | yes | - |
1-076 | 4690509049268759808 | [MWD2000] h53-60 | 14.54 | O8 III [Mas00] | O9 III: | no | - |
1-077 | 4690512450882740480 | RMC 30, AzV 331 | 12.57 | A2 I [Mas00] | A2 II/Ib | no | - |
1-078 | 4690513859631967616 | - | 15.05 | - | O9.7 III: | no | - |
1-079 | 4690512828839860096 | [MA93] 1380 | 13.31 | - | O9.2 III:(n) pe | yes | - |
1-080 | 4690509358506360320 | [MWD2000] h53-91 | 14.78 | O8.5 V [Mas00] | O8: V: + (early B + early B) | no | - |
1-081 | 4690509427225802624 | - | 16.95 | - | B1.5: III: e | no | - |
1-082 | 4690525988621394816 | [M2002] SMC 56680 | 15.57 | - | B1.5 III-II | no | - |
1-083 | 4690512760120397312 | [MWD2000] h53-101 | 14.84 | - | B0.5: V: + early B | no | - |
1-084 | 4687505527110195072 | [M2002] SMC 56785 | 15.44 | - | B1 II | no | - |
1-085 | 4687503946562429184 | Dachs SMC 2-14, AzV 84F | 12.46 | A7 II [Eva04] | A7 Ib/ab | no | - |
1-086 | 4687505454063447936 | [M2002] SMC 56834 | 15.44 | O7-9 [She13] | B0 IV: | no | - |
1-087 | 4690509732152391680 | [M2002] SMC 56942 | 15.07 | - | B1 II: | no | - |
1-088 | 4690508258994548352 | [M2002] SMC 57134 | 15.53 | - | B1.5 III + early B | no | - |
1-089 | 4690514340668263552 | [M2002] SMC 57144 | 14.47 | - | B1 II | no | - |
1-090 | 4687505252232336512 | OGLE SMC-SC10 28913 | 14.40 | B9 (II) [Eva04] | A1 Ib | no | - |
1-091 | 4690512794480131072 | [MWD2000] h53-137 | 14.20 | O8.5 III [Mas00] | O9.5 III: + O9.2 | no | - |
1-092 | 4690513997070898944 | [M2002] SMC 57269 | 15.29 | - | B2 II | no | - |
1-093 | 4690509766512143232 | AzV 339a | 11.94 | F5 [Azz82] | A7 Ib | no | - |
1-094 | 4687505217872848000 | OGLE SMC-SC10 28919 | 15.68 | B1-3 (III) [Eva04] | B1.5 II: | no | - |
1-095 | 4690508632639753984 | RMC 33, AzV 349 | 12.82 | B0 IIIw [Gar87] | B1 Ia | no | - |
1-096 | 4690509495924424960 | - | 16.60 | - | B1 II | no | - |
1-097 | 4690509942621639680 | RMC 35, AzV 342 | 13.30 | B2.5 Iab [Len97] | B2.5 Ib | no | - |
1-098 | 4690512966278545792 | - | 14.76 | - | B1 Ib | no | - |
1-099 | 4690508671311394816 | - | 15.78 | - | B2: II: | no | - |
1-100 | 4690508740030862208 | [M2002] SMC 58016 | 14.86 | - | B1 II | no | - |
1-101 | 4690508671311389824 | [M2002] SMC 58132 | 15.78 | - | B1.5 III-II | no | - |
1-102 | 4690509599024293376 | AzV 345a | 14.29 | O7 III [Gar87] | O6 III(n) | no | - |
1-103 | 4690508774390594944 | [MWD2000] h53-197 | 14.78 | B1 I [Mas00] | B1 II: | no | - |
1-104 | 4690514409387557376 | [M2002] SMC 58803 | 15.46 | - | O9 V: | no | - |
1-105 | 4687504874275456896 | - | 16.42 | - | B2: II | no | - |
1-106 | 4687505801988031232 | LHA 115-S 41, AzV 351 | 13.23 | B1-2 (Ib)e [Eva04] | B1.5 Ib: e | yes | - |
1-107 | 4690508808750315776 | - | 16.49 | - | B0.5 III-II | no | - |
1-108 | 4687504977354372864 | - | 15.49 | - | B1 II: | no | - |
1-109 | 4690510286219053312 | AzV 356 | 14.03 | B0 Iw [Gar87] | B1.5 Ib | no | - |
1-110 | 4690510286219047936 | AzV 358a | 14.53 | B0 I [Azz79] | B1 Ib | no | - |
1-111 | 4687507584364583040 | RMC 36, Sk 114, AzV 362 | 11.34 | B3 Ia [Len97] | B3 Ia | no | - |
1-112 | 4687507245097014144 | RMC 37, AzV 367 | 10.88 | B9 Ia+ [Len97] | B9 Ia | yes | - |
1-113 | 4690511626248698368 | AzV 364 | 14.39 | B2 IIIe [Men06] | B1 II e | yes | - |
1-114 | 4690510492377435776 | Sk 118, AzV 369 | 9.71 | F4 Ia [Ard77] | F5: | no | - |
1-115 | 4690510419368517760 | OGLE SMC113.6 39585 | 15.27 | - | B1 II: | no | - |
1-116 | 4687507554335030656 | OGLE SMC-SC10 106172 | 15.20 | B0.5 (IV) [Eva04] | B1 II | no | - |
2-001 | 4688959562535054976 | - | 16.55 | - | O9.2 V: | no | - |
2-002 | 4688959356376632960 | - | 16.43 | - | B2 II | no | - |
2-003 | 4688981823353120896 | 2dFS 745 | 14.94 | B1-2 (II) [Eva04] | B1 II | no | - |
2-004 | 4688965575504047744 | 2dFS 5036 | 15.65 | B1-5 (III) [Eva04] | B3 II: | no | - |
2-005 | 4688983614358317184 | AzV 67 | 14.17 | O9 III [Eva04] | O8.5 II:(n) | no | - |
2-006 | 4688981857676434816 | - | 15.40 | - | B1.5 III | no | - |
2-007 | 4688978838366817280 | Sk 35, AzV 70 | 12.88 | O9.5 Ib [Wal02] | O9.5 II-I | no | - |
2-008 | 4688982235684518016 | OGLE SMC719.12.035620 | 13.84 | O9 III: [Lam13] | O9 II: | no | - |
2-009 | 4688977463965651712 | [M2002] SMC 16769 | 16.00 | - | O9.7 IV(n) | no | - |
2-010 | 4688964506004003840 | - | 15.23 | - | B1.5 III-II | no | - |
2-011 | 4688977837586212608 | - | 16.24 | - | B1 II: | no | - |
2-012 | 4688979250671773184 | - | 13.87 | - | A0 Ib | no | - |
2-013 | 4688982343048079488 | - | 15.59 | - | B1 II | no | - |
2-014 | 4688964613419753984 | - | 15.65 | - | B1 II | no | - |
2-015 | 4688978219879759360 | - | 15.54 | - | B2.5 II: | no | - |
2-016 | 4688964334254663168 | AzV 80 | 13.77 | O4-6n(f)p [Wal00] | O6 III:nn(f)p | no | (f)p from [Wal00] |
2-017 | 4688982652297652352 | - | 14.12 | - | B2 II e | yes | - |
2-018 | 4688966090888390784 | LIN 157, AzV 82 | 14.70 | B2e [Azz75] | O6.5 III: neb | yes | - |
2-019 | 4688964918322093312 | - | 15.13 | - | O9.7 V: | no | - |
2-020 | 4688966125248119808 | AzV 83 | 13.99 | O7 Iaf+ [Wal00] | O7 Iaf+ | no | f+ from [Wal00] |
2-021 | 4688964922657391744 | LIN 163 | 14.08 | - | B2.5: II: e | yes | - |
2-022 | 4688978254239473408 | - | 13.34 | - | A0 Ib | no | - |
2-023 | 4688978048081059712 | - | 15.26 | - | B2.5 III | no | - |
2-024 | 4688977979361634816 | OGLE SMC-SC5 316725 | 15.15 | O9 + O9 [Har03] | O9.7: V + O9.7 | no | - |
2-025 | 4688978048081047296 | - | 15.45 | - | B3 II: | no | - |
2-026 | 4688965777314555264 | OGLE SMC101.1 39365 | 14.85 | - | B2 II + early B | no | - |
2-027 | 4688961624122202624 | - | 17.00 | - | B1.5 III neb | no | - |
2-028 | 4688977979361594880 | AzV 87 | 14.31 | B0 IIw [Gar87] | B1.5 Ib | no | - |
2-029 | 4688961654195517952 | 2dFS 5047 | 15.42 | O9.5 V [Eva04] | O9.7 V | no | LC from [Eva04] |
2-030 | 4688978082440780672 | [M2002] SMC 19070 | 15.57 | - | B2 II | no | - |
2-031 | 4688966189635061760 | Dachs SMC 1-5, AzV 93 | 14.53 | B1 Ib [Mas95] | B1 II | no | - |
2-032 | 4688964853937446912 | - | 15.05 | - | B1 II | no | - |
2-033 | 4688961452323508736 | - | 16.99 | - | B1 III-II | no | - |
2-034 | 4688961383604048768 | - | 15.63 | - | B1 II | no | - |
2-035 | 4688965060095843072 | AzV 95 | 14.15 | O7 III((f)) [Wal00] | O7.5 III((f)) | no | ((f)) from [Wal00] |
2-036 | 4688960627689838976 | - | 15.98 | - | B2 II: | no | - |
2-037 | 4688961761561107456 | AzV 99 | 13.17 | B2.5 Iab [Len97] | B2.5 Ib | no | - |
2-038 | 4688967735823484800 | - | 14.80 | - | B1.5 II | no | - |
2-039 | 4688961864640239488 | AzV 103 | 13.84 | B0.7 Ib [Len97] | B1 II-Ib | no | - |
2-040 | 4688984026675105280 | - | 14.74 | - | B2 II | no | - |
2-041 | 4688960657713202304 | - | 16.75 | - | B2 II | no | - |
2-042 | 4688960765128725760 | - | 15.49 | - | B2 IV: | no | - |
2-043 | 4688961795920796928 | [M2002] SMC 20706 | 13.68 | - | B1 Iab | no | - |
2-044 | 4688967465277409408 | - | 16.08 | - | B2: II | no | - |
2-045 | 4688960765128723584 | - | 16.26 | - | B2 III: | no | - |
2-046 | 4688967396557954432 | - | 14.94 | - | B1 II | no | - |
2-047 | 4688979594268679296 | AzV 106 | 14.85 | B1 II [Lam16] | B1 Ib | no | - |
2-048 | 4688981003018180992 | 2dFS 821 | 15.63 | O9 V [Eva04] | O9.2 V | no | - |
2-049 | 4688984404616099200 | - | 16.33 | - | B1.5: III | no | - |
2-050 | 4688979972225779968 | AzV 107, 2dFS 822 | 12.52 | F0 [Eva04] | A5 Ib | no | - |
2-051 | 4688960421531447296 | OGLE J005146.64-725121.3 | 16.58 | B1 [Zas14] | B2.5 IV: | no | - |
2-052 | 4688967705795512448 | AzV 109 | 14.66 | O9.5 + B0.5 [Har03] | B0.5 V: + B0-0.7 | no | - |
2-053 | 4688960485917320448 | - | 16.11 | - | B2 III | no | - |
2-054 | 4688966292708802304 | Sk 49, AzV 110 | 11.72 | A3 Ia [Neu10] | A0 Ia | no | - |
2-055 | 4688984061034529408 | AzV 102 | 14.26 | O9.5-B0 IV-V [Coe15] | O9.7 V:n e | no | - |
2-056 | 4688981106097035008 | AzV 111 | 14.07 | B0 II [Gar87] | B2 II: e | no | - |
2-057 | 4688980831219164416 | 2dFS 832 | 14.49 | B0 (III) [Eva04] | B0.7 II | no | - |
2-058 | 4688963303411565056 | Dachs SMC 1-9, AzV 32F | 12.22 | A3 I [Neu10] | A5 Ib | no | - |
2-059 | 4688967602716295168 | AzV 114 | 15.39 | O8 V [Eva04] | O7.5 V(n) | no | LC from [Eva04] |
2-060 | 4688963307749203072 | - | 14.55 | - | B1.5 Ib | no | - |
2-061 | 4688961894666576768 | [MA93] 536 | 14.77 | B1 [She13] | B2 II: e | no | - |
2-062 | 4688963410828391680 | [M2002] SMC 22729, 2dFS 850 | 15.15 | B0-3 (III) [Eva04] | B1 III-II | no | - |
2-063 | 4688963204681907328 | [MA93] 555 | 14.73 | - | B2 III e | yes | - |
2-064 | 4688966537564476288 | MOA J005219.2-724151 | 15.31 | B0.2 + B1 [Har03] | early B + early B | no | - |
2-065 | 4688967568356557440 | Dachs SMC 1-11, AzV 121 | 9.56 | F5 Ia [Hum83] | F5: | no | - |
2-066 | 4688966739421000448 | [M2002] 23352 | 14.18 | B0-2 [Pau12] | O9.7 III:nnn pe+ | yes | - |
2-067 | 4685959235845498240 | AzV 123 | 13.24 | B8 [Azz75] | B8 Ib | no | - |
2-068 | 4688980414561766016 | 2dFS 5057 | 12.79 | B9 (Iab) [Eva04] | B9 Iab | no | - |
2-069 | 4688963887532390528 | 2dFS 875 | 15.44 | O9 V [Eva04] | O9 V | no | - |
2-070 | 4688962414395649792 | [MA93] 584 | 13.88 | B1-2 (II) [Eva04] | B1 II e | no | - |
2-071 | 4688968049392847488 | 2dFS 5058 | 14.86 | B1-3 (III)e? [Eva04] | B1.5 Ib e | no | - |
2-072 | 4685959373284393728 | 2dFS 880 | 13.53 | A2 II [Eva04] | A2 II/Ib | no | - |
2-073 | 4688963548266999808 | Dachs SMC 1-16, AzV 131 | 12.34 | A0 [Azz75] | B9 Ia | no | - |
2-074 | 4688968049392837248 | 2dFS 881 | 15.12 | B1-5 (II) [Eva04] | B3 II: | no | - |
2-075 | 4688968358630475136 | AzV 133 | 14.48 | O6.5 V((f)) [Eva04] | O6 Vn((f)) | no | ((f)) from [Eva04] |
2-076 | 4688966881161489280 | [M2002] SMC 24500 | 15.30 | - | B1 Ib | no | - |
2-077 | 4688963926224354816 | - | 15.25 | - | B1 II | no | - |
2-078 | 4688963651346132224 | Dachs SMC 1-17 | 14.71 | - | B0 II | no | - |
2-079 | 4688963548266951808 | [M2002] SMC 24767 | 15.26 | - | O4 V: + early B | no | - |
2-080 | 4688962517474878720 | - | 13.00 | - | A2 II/Ib | no | - |
2-081 | 4688968461709335040 | [M2002] SMC 24895 | 14.13 | B0-3 [Pau12] | B2 II e | yes | - |
2-082 | 4685960021758355584 | AzV 138 | 13.98 | O9 III-Vpe [Lam16] | O9.2 III pe | no | - |
2-083 | 4688963582626684544 | Sk 53, AzV 137 | 12.64 | B2 Iab [Len97] | B2.5 Ia | no | - |
2-084 | 4688962448755796608 | - | 15.33 | - | B2 II: | no | - |
2-085 | 4688963582627047424 | - | 15.87 | - | O9 V | no | - |
2-086 | 4688962757993027328 | [M2002] SMC 25142 | 15.45 | - | O8 V(n) | no | - |
2-087 | 4688967323503140992 | - | 15.04 | - | O9.5 III | no | - |
2-088 | 4688968461709333632 | 2dFS 898 | 13.16 | A5 II [Eva04] | A5 Ib | no | - |
2-089 | 4688968118111968512 | OGLE SMC-SC6 167473 | 15.27 | - | B2 II | no | - |
2-090 | 4688966984219996288 | [M2002] SMC 25387 | 14.98 | B0 [She13] | O7.5 Vn | no | - |
2-091 | 4688967052939420672 | Dachs SMC 1-21 | 14.32 | - | O8.5 V | no | - |
2-092 | 4688967289176997504 | Dachs SMC 1-22 | 12.17 | - | B8 Iab | no | - |
2-093 | 4688967362197739392 | HD 5291, RMC 11, Sk 56 | 10.58 | B8 Ia+ [Len97] | B8 Ia | no | - |
2-094 | 4688964029303180416 | [M2002] SMC 25741 | 15.27 | - | B0.5 III-II | no | - |
2-095 | 4688963960583732736 | - | 16.20 | - | B0 IV | no | - |
2-096 | 4688962483115141248 | [M2002] SMC 25866 | 14.97 | - | O9.5 III: | no | - |
2-097 | 4689003680451206272 | LIN 230, AzV 141, 2dFS 915 | 14.27 | B5 (II) [Eva04] | B5 Ia: e | yes | - |
2-098 | 4688991585819715712 | 2dFS 5066 | 14.99 | O6 V((f)) [Eva04] | O6.5 V((f)) | no | LC from [Eva04] |
2-099 | 4688991585819679104 | - | 16.30 | - | B0 IV e | no | - |
2-100 | 4688962929791640064 | - | 15.43 | - | B0 V | no | - |
2-101 | 4688991585819641088 | - | 13.53 | - | B9 Ib | no | - |
2-102 | 4688990555027643904 | AzV 143, 2dFS 932 | 14.54 | B0 (III) [Eva04] | B0 II: | no | - |
2-103 | 4688987256492902528 | [M2002] SMC 27080 | 15.32 | - | B0.5 V: + B0: | no | - |
2-104 | 4688987252155037568 | 2dFS 936 | 14.68 | O6.5f?p [Eva04] | O5.5f?pe | yes | f?p from [Eva04] |
2-105 | 4688991688898752896 | AzV 39F | 13.57 | A0 [Azz79] | A2 II/Ib | no | - |
2-106 | 4688990589387328384 | AzV 145 | 13.99 | B1 [Azz75] | B0.7 II | no | - |
2-107 | 4685960129198503936 | - | 14.99 | - | B1.5 III-II | no | - |
2-108 | 4685983665623961216 | - | 14.89 | - | B5 II | no | - |
2-109 | 4688990692466513408 | 2dFS 5071 | 15.00 | B1-5 (II) [Eva04] | B2 II e | no | - |
2-110 | 4688986431859207296 | AzV 148 | 14.97 | O8.5 V [Mas02] | B0 II | no | - |
2-111 | 4685983699984262912 | [MA93] 715 | 14.67 | - | B2 II: e | yes | - |
2-112 | 4688987389594710272 | OGLE SMC-SC6 447961 | 14.57 | - | B2 II: | no | EB [Paw16] |
2-113 | 4685983420730442624 | Sk 57, AzV 151 | 12.36 | B2.5 Ia [Len97] | B2.5 Ia | no | - |
2-114 | 4688986569298097792 | - | 14.58 | - | B1.5 II | no | - |
2-115 | 4688990658106753280 | AzV 153 | 13.71 | B8 [Azz75] | B5 Ib | no | - |
2-116 | 4688986534938381184 | LHA 115-S 18, AzV 154 | 11.11 | B[e] [Lam16] | sgB[e] | yes | - |
3-001 | 4685853682687097088 | Sk 6, AzV 9 | 13.20 | B0 III [Smi97] | B2.5 Ib | no | - |
3-002 | 4685853609694629888 | - | 15.83 | - | O6 V:nn | no | - |
3-003 | 4685854056331151232 | 2dFS 5005 | 15.74 | B1-5 (II) [Eva04] | B1.5 III | no | - |
3-004 | 4685849696957595136 | - | 16.65 | - | O9.7 IV: | no | - |
3-005 | 4685849319000715904 | - | 16.30 | - | B1 III | no | - |
3-006 | 4685849623922204160 | - | 13.65 | - | B5: + early A Ia | no | - |
3-007 | 4685848597446171904 | - | 12.37 | - | A5 Ib | no | - |
3-008 | 4685836880771632512 | - | 16.18 | - | O7.5 V:(n) | no | - |
3-009 | 4685850036241667840 | - | 14.52 | - | B5 II | no | - |
3-010 | 4685836876461619328 | - | 15.54 | - | O9.7 V: | no | - |
3-011 | 4685854369881566848 | LIN 84 | 14.39 | - | B0.5 III: e | yes | - |
3-012 | 4685854571724688768 | HA 115-S 6, RMC 4, Sk 11, AzV 16 | 11.02 | B0[e] [Zic96] | sgB[e] | yes | - |
3-013 | 4685848666165653504 | - | 13.82 | - | A0 Ib | no | - |
3-014 | 4685848734885047680 | [M2002] SMC 7782 | 14.86 | O7.5 V [Lam13] | O8 Vn | no | - |
3-015 | 4685850139318317312 | - | 16.73 | - | B2 III | no | - |
3-016 | 4685849044122619264 | [SBV2013] B11 | 16.82 | - | B2 III e | no | - |
3-017 | 4685836571534016128 | [M2002] SMC 7909 | 14.44 | - | B0.7 II | no | - |
3-018 | 4685836635962991104 | LIN 88 | 13.84 | - | B1.5: II: e | yes | - |
3-019 | 4685948481244298112 | 2dFS 5010 | 15.04 | O8.5 V [Eva04] | O9.2 V | no | LC from [Eva04] |
3-020 | 4685849112841980928 | [M2002] SMC 8017 | 15.42 | - | B0.5 III | no | - |
3-021 | 4685835536431612288 | - | 14.82 | - | B1.5 II | no | - |
3-022 | 4685948412524815616 | OGLE SMC-SC4 120783 | 14.65 | - | B2 II e | no | - |
3-023 | 4685849078482253440 | - | 15.17 | - | B3 II | no | - |
3-024 | 4685850177993655424 | - | 14.88 | - | B5 II | no | - |
3-025 | 4685850242400168960 | [MA93] 167 | 14.50 | - | B2 II e | yes | - |
3-026 | 4685947038135411200 | - | 15.06 | - | B1.5 II-III | no | - |
3-027 | 4685948515603957376 | LIN 91 | 13.44 | B0-2 [Pau12] | B1.5 II e | yes | - |
3-028 | 4685835575101684608 | [M2002] SMC 8609 | 14.52 | B0 III [Lam16] | B0.5 II | no | - |
3-029 | 4685944117556288640 | [M2002] SMC 8649 | 14.34 | - | B1.5 Ib | no | - |
3-030 | 4685947141214597376 | BBB SMC 289, AzV 21 | 14.80 | B1 [Azz75] | B1 II | no | - |
3-031 | 4685836708972891648 | [MA93] 177 | 14.14 | - | B0: III: + OB e | yes | - |
3-032 | 4685836708972886400 | - | 14.56 | - | B1.5 II | no | - |
3-033 | 4685948275085868928 | - | 15.50 | - | O9.5 IV:n e? | no | - |
3-034 | 4685943773959001984 | [M2002] SMC 8909 | 14.72 | - | B0 III | no | - |
3-035 | 4685835884339216512 | - | 16.34 | - | B2 II e | no | - |
3-036 | 4685947519171645952 | 2dFS 5014 | 15.28 | B1-3 (II) [Eva04] | B0.7 III-II | no | - |
3-037 | 4685835884339219456 | Sk 17, AzV 23 | 11.96 | B3 Ia [Len97] | B3 Ia | no | - |
3-038 | 4685930614175426560 | - | 13.16 | - | A0 Ia | no | - |
3-039 | 4685944151915964928 | AzV 25 | 13.25 | B5 [Azz75] | B3 Ib | no | - |
3-040 | 4685942949325301120 | - | 14.00 | - | B2 Ib e | yes | - |
3-041 | 4685944147624292224 | - | 13.25 | - | B8 Ib | no | - |
3-042 | 4685947553531340800 | BBB SMC 266, Sk 18, AzV 26 | 12.94 | O6 I(f) [Mas04] | O6 I(f) + O7.5 | no | LC from [Mas04] |
3-043 | 4685943121123943680 | Sk 19, AzV 27 | 11.63 | B8 I [Neu10] | A2 Iab | no | - |
3-044 | 4685929750832616320 | AzV 28, 2dFS 668 | 13.88 | B1-2 (II) [Eva04] | B1 Ib | no | - |
3-045 | 4685944250703291520 | - | 13.44 | - | A0 Ib | no | - |
3-046 | 4685947209934051200 | [M2002] SMC 9532 | 14.45 | - | O8.5 V(n) | no | SB |
3-047 | 4685944250646464768 | [M2002] SMC 9534 | 14.27 | - | B1 II | no | - |
3-048 | 4685926491006945536 | - | 15.29 | - | B0.7 III | no | - |
3-049 | 4685926147408274304 | [M2002] SMC 9647 | 15.08 | - | O4 I(n) | no | - |
3-050 | 4685943121123938048 | OGLE SMC720.28.039935 | 15.36 | - | B1.5 III | no | EB [Paw16] |
3-051 | 4685942777526639744 | [MA93] 203 | 14.81 | - | O5.5: V neb | no | - |
3-052 | 4685942777526626432 | [M2002] SMC 9845 | 15.16 | - | O7 V: neb + O7.5 | no | - |
3-053 | 4685943082442898432 | LIN 106 | 14.87 | O8 V [Tes01] | O7.5 V(n) neb | no | - |
3-054 | 4685929926980635136 | OGLE SMC720.28.000244 | 15.05 | - | O7 V(n) neb | no | - |
3-055 | 4685944285006166144 | [M2002] SMC 10202 | 15.42 | - | B0.7 IV | no | - |
3-056 | 4685947652249342080 | OGLE SMC-ECL- 1376 | 14.46 | - | B1.5 Ib | no | EB [Paw13] |
3-057 | 4685926181767968000 | [M2002] SMC 10209 | 15.54 | - | B1 II | no | - |
3-058 | 4685926559725017600 | AzV 34 | 13.11 | B2: [Azz75] | A0 Iab | no | - |
3-059 | 4685947897128639744 | - | 14.38 | - | B1 II | no | - |
3-060 | 4685925146630858240 | [M2002] SMC 10505 | 15.31 | O7-9 [She13] | O6 Vn: | no | - |
3-061 | 4685943254211136000 | AzV 28F | 13.00 | B5 [Azz79] | B3 Ib | no | - |
3-062 | 4685948064629440768 | Sk 22 | 14.01 | OB: [San68] | O9.7 II e | no | - |
3-063 | 4685947622250808832 | [M2002] SMC 10818 | 14.70 | - | O9.2 II(n) | no | - |
3-064 | 4685926353526652160 | - | 14.99 | - | B3 II | no | - |
3-065 | 4685943254211144832 | [M2002] SMC 10947 | 14.66 | - | B0.7 II | no | - |
3-066 | 4685926319206869632 | - | 14.95 | - | O7 III: | no | - |
3-067 | 4685948000207808384 | OGLE SMC-SC4 175149 | 15.09 | - | B0.2 IV | no | - |
3-068 | 4685930334952220416 | - | 13.75 | - | B5 Ib | no | - |
3-069 | 4685931468820884864 | - | 14.97 | - | B1.5 II: | no | - |
3-070 | 4685926662804412800 | - | 15.36 | - | B1 II | no | - |
3-071 | 4685925563292648576 | [M2002] SMC 11294 | 16.17 | - | B0.2 V | no | - |
3-072 | 4685943288573729920 | 2dFS 5021 | 15.36 | B1-5 (II) [Eva04] | B1 III | no | EB [Paw13] |
3-073 | 4685943219851406208 | OGLE SMC-SC5 26508 | 13.20 | - | B3 Ib | no | EB [Paw13] |
3-074 | 4685943670879523840 | [SBV2013] B37 | 15.19 | B1 [She13] | B1 II | no | - |
3-075 | 4685925872530280192 | AzV 44 | 14.45 | B0 IIww [Gar87] | O6.5 V:nnn | no | EB [Paw13], SB2? |
3-076 | 4685925872530300544 | - | 16.10 | - | O8 Vn | no | - |
3-077 | 4685947828409081088 | [MA93] 256 | 14.32 | B0-2 [Pau12] | B1 II e | yes | - |
3-078 | 4685925666371896576 | AzV 47 | 14.16 | O8 III((f)) [Wal00] | O8 III((f)) | no | LC from [Wal00] |
3-079 | 4685925936902711424 | [M2002] SMC 12022 | 15.28 | - | B1.5 III + B2.5: | no | - |
3-080 | 4685931232650460032 | - | 15.36 | - | B1 III-II | no | - |
3-081 | 4685946282220487424 | [MA93] 265 | 14.34 | - | O6 III: | no | - |
3-082 | 4685926800243109760 | OGLE SMC100.8 45085 | 14.60 | - | B1 II + B1.5: | no | - |
3-083 | 4685943636519760512 | - | 13.71 | - | A0 Ib | no | - |
3-084 | 4685931335729612288 | [M2002] SMC 12787 | 14.54 | - | B0 Ib: | no | - |
3-085 | 4685926009969170816 | AzV 49 | 13.54 | A0 [Azz75] | B9 Iab | no | - |
3-086 | 4685931267010164480 | - | 14.76 | - | B1 III: + B2.5: | no | - |
3-087 | 4685944839111584256 | OGLE J004915.62-731210.4 | 12.57 | - | A2 II/Ib | no | - |
3-088 | 4685945968691683456 | - | 13.12 | - | A2 II | no | - |
3-089 | 4685926933335252992 | - | 16.50 | - | B2 III: | no | - |
3-090 | 4685932744479939584 | - | 13.76 | - | B0.2 Ia | no | - |
3-091 | 4685927212559921536 | - | 15.20 | - | B1 III-II | no | - |
3-092 | 4685933053717525632 | - | 14.44 | - | B5 II | no | - |
3-093 | 4685932843197054208 | - | 16.62 | - | B1 II: | no | - |
3-094 | 4685945247096928256 | [M2002] SMC 13774 | 16.17 | O7 [She13] | O8: V(n) | no | - |
3-095 | 4685925833826106368 | [M2002] SMC 13986 | 15.10 | - | B2 Ib: e | no | - |
3-096 | 4685928410806791936 | - | 14.18 | - | B2.5 II | no | - |
3-097 | 4685928415151555712 | 2dFS 5031 | 14.55 | B1-3 (II) [Eva04] | B2.5 II | no | - |
3-098 | 4685931679327948800 | [M2002] SMC 14324 | 14.25 | O6 V((f))e [Lam16] | O6 V((f))n e | no | LC from [Lam16] |
3-099 | 4685928724389105536 | AzV 55 | 13.24 | B5 [Azz75] | B9 Iab | no | - |
3-100 | 4685928449511258112 | - | 15.22 | - | B3 II | no | - |
3-101 | 4685928445163783552 | OGLE SMC100.8 52871 | 14.82 | O9 + O9 [Hil05] | O9.7 V: + O9.7 | no | - |
3-102 | 4685932052995296768 | - | 14.65 | - | B8 II-Ib | no | - |
3-103 | 4685932126004485376 | - | 15.16 | - | B1 II-Ib | no | - |
3-104 | 4685933191156274432 | BBB SMC 126, AzV 59 | 13.27 | A0 [Azz75] | B8 Iab | no | - |
3-105 | 4685927620532759808 | OGLE SMC-SC5 95214 | 15.19 | B0 (IV) [Eva04] | B0.5 III-II | no | - |
3-106 | 4685945010910000640 | AzV 63 | 13.34 | A0 [Azz75] | B5 Ib | no | - |
3-107 | 4685880277148119040 | AzV 64 | 14.32 | B1 [Azz75] | O9.7 II:(n) e? | yes | - |
3-108 | 4685927727956781568 | - | - | - | B0 III | no | - |
3-109 | 4685931816766886656 | AzV 66 | 14.10 | B0. 5V [Mas95] | B0.2 Ib | no | - |
3-110 | 4685933569113371904 | - | 14.28 | - | B8 II-Ib | no | - |
3-111 | 4685931812407476864 | - | 13.96 | - | B1 Ib | no | - |
3-112 | 4685931816766868224 | OGLE SMC100.2 30 | 14.78 | - | O6.5 V-III | no | - |
3-113 | 4685927453078935680 | - | 14.77 | - | B2 II: | no | - |
3-114 | 4685928793108545152 | ISO-MCMS J005013.7-731710 | 12.00 | - | F2: | no | - |
3-115 | 4685928861827989120 | OGLE SMC100.2 35 | 14.66 | - | B1 II | no | - |
3-116 | 4685928071554123264 | - | 15.04 | - | B3 II | no | - |
4-001 | 4689018661298288000 | [BLK2010] flames1004 | 13.04 | A2 Ib [Duf19] | A1 Ib | no | - |
4-002 | 4689016049958328576 | AzV 191, Cl* NGC 346 ELS 4 | 13.21 | Be(B1:) [Eva06] | B1.5: Ia e | yes | - |
4-003 | 4689019073615200768 | 2dFS 1252 | 15.69 | B0.5 III [Duf19] | B1 II | no | - |
4-004 | 4689067585206690176 | - | 15.40 | - | B5 II: | no | - |
4-005 | 4689016049958297600 | 2dFS 1259,Cl* NGC 346 ELS 20 | 15.41 | B1 V+early B [Eva06] | B1 III-II | no | SB2 [Eva06] |
4-006 | 4689002851458876800 | AzV 197 | 10.58 | F0 I [Neu10] | F2: | no | SB2? |
4-007 | 4689014709928718464 | [BLK2010] flames1028 | 15.16 | B2.5 III [Duf19] | B1.5 III | no | - |
4-008 | 4689019589011024128 | AzV 68F, Cl* NGC 346 ELS 2 | 12.99 | A2: Iab [Eva06] | A0 Ib | no | - |
4-009 | 4689014778648125824 | Cl* NGC 346 ELS 52 | 15.50 | B1.5 V [Eva06] | B2 III | no | SB1 [Eva06] |
4-010 | 4689016221756877184 | AzV 201 | 12.99 | B1e [Duf19] | B1 Ib e | yes | - |
4-011 | 4689014808648096896 | OGLE SMC-SC8 107285 | 15.48 | B0 III [Duf19] | B1 II | no | - |
4-012 | 4689020722882360192 | Cl* NGC 346 ELS 84, 2dFS 1296 | 16.46 | B1 V [Eva06] | B2 III | no | - |
4-013 | 4689019245413731840 | Cl* NGC 346 MPG 11, ELS 43 | 15.87 | B0 V [Eva06] | B0.2 IV | no | - |
4-014 | 4689016221756858624 | Cl* NGC 346 MPG 12, ELS 26 | 15.33 | B0 IV (Nstr) [Eva06] | B0 III | no | - |
4-015 | 4689019898248677888 | AzV 202, Cl* NGC 346 ELS 12 | 14.94 | B1 Ib [Eva06] | B1 II-Ib | no | - |
4-016 | 4689020825961546496 | [BLK2010] flames1046 | 15.38 | B0.5 V [Duf19] | B0.7 III | no | - |
4-017 | 4689002787095701248 | [BLK2010] flames1041 | 14.58 | B2 III-II [Duf19] | B1.5 Ib | no | - |
4-018 | 4689002787095693568 | [BLK2010] flames1067 | 16.13 | O9 V [Duf19] | O9.5 IV | no | - |
4-019 | 4689002718376234112 | OGLE SMC-SC8 104191 | 15.20 | B1 III [Duf19] | B1 II | no | - |
4-020 | 4689003027613801216 | Sk 73, AzV 210 | 12.78 | B1.5 Ia [Len97] | B1 Iab-Ib | no | RV variable [Duf19] |
4-021 | 4689015294044085248 | [SBV2013] B135 | 15.47 | O7-9 [She13] | O9.2 V | no | - |
4-022 | 4689015500202439296 | LIN 336, Cl* NGC 346 MPG 184 | 14.73 | B2-3e [Duf19] | B2.5 II: e | yes | - |
4-023 | 4689015225324629120 | OGLE SMC-SC8 110238 | 15.50 | B2.5 III-II [Duf19] | B2 II: | no | - |
4-024 | 4689003126336982272 | OGLE SMC-SC8 107276 | 15.22 | B1.5 II [Duf19] | B1 II | no | - |
4-025 | 4689001721943896192 | - | 15.38 | - | B1 III | no | - |
4-026 | 4689015328403760128 | Cl* NGC 346 MPG 217, ELS 18 | 14.60 | O9.5 IIIe [Eva06] | O9.5 IIIpe | no | LC from [Eva06] |
4-027 | 4689001721943908096 | - | 16.90 | - | B2 II | no | - |
4-028 | 4689001653225111680 | OGLE SMC-SC8 104178 | 15.71 | B1-3 (III) [Eva04] | B2 II | no | - |
4-029 | 4689020070047305472 | Cl* NGC 346 ELS 38 | 15.73 | B1 V [Eva06] | B1.5 IV | no | EB [Paw13] |
4-030 | 4689015259684292608 | AzV 214, Cl* NGC 346 MPG 293 | 13.34 | B1 Ia [Duf19] | B1 Ia | no | - |
4-031 | 4689016462274935296 | Cl* NGC 346 MPG 304, ELS 93 | 16.16 | B0 V [Eva06] | B0 V neb | no | - |
4-032 | 4689003508650071040 | [BLK2010] flames1023 | 14.53 | B0 V [Duf19] | O9.7 III: | no | EB SB1 [Duf19] |
4-033 | 4689016840231937792 | Cl* NGC 346 ELS 27 | 15.60 | B0.5 V [Eva06] | B0.7 III | no | - |
4-034 | 4689015736378946688 | LIN 345, Cl* NGC 346 MPG 482 | 14.28 | B0.5e [Duf19] | B2 II pe | yes | - |
4-035 | 4689016771512472960 | Cl* NGC 346 ELS 24 | 14.90 | B2:(Be-Fe) [Eva06] | B3 Ib e | yes | - |
4-036 | 4689002409138503040 | AzV 219 | 15.10 | B1 III [Duf19] | B1.5 II | no | SB1 [Duf19] |
4-037 | 4689015667659473920 | Cl* NGC 346 NMC 13, MPG 549 | 13.14 | O9.5 V [Duf19] | O9.7 IV neb | no | - |
4-038 | 4689002134260639232 | - | 16.08 | - | B1.5 III-II | no | - |
4-039 | 4690521006458654464 | AzV 220 | 15.03 | O6.5f?p [Wal00] | O6.5f?pe | yes | f?p from [Wal00] |
4-040 | 4689015736378951936 | Cl* NGC 346 ELS 33 | 13.12 | O8 V [Eva06] | O7 V(n) neb | no | - |
4-041 | 4689015465788811008 | Cl* NGC 346 MPG 602 | 15.19 | O6.5 V((f)) [Duf19] | O6.5 V: neb | - | - |
4-042 | 4688999041884354944 | AzV 222 | 13.45 | B3 [Mas00] | B2.5 Ib | no | - |
4-043 | 4689015397123159424 | Cl* NGC 346 MPG 637, ELS 29 | 15.37 | B0 V [Eva06] | O9.5 V neb | no | SB1 [Eva06] |
4-044 | 4689002237339841792 | [BLK2010] flames1176 | 16.85 | B1.5 V [Duf19] | B1.5 III neb | no | - |
4-045 | 4690521109537837440 | AzV 224, Cl* NGC 346 ELS 8 | 14.73 | B1e [Eva06] | B1 Iab | no | - |
4-046 | 4688998835725922176 | [MA93] 1135, 2dFS 1375 | 14.15 | B1-5 (II)e? [Eva04] | B2 II: | yes | - |
4-047 | 4690518085880908544 | Cl* NGC 346 ELS 37 | 15.62 | B3 III [Eva06] | B2 II | no | - |
4-048 | 4689003371211130624 | Cl* NGC 346 ELS 21, 2dFS 5099 | 15.33 | B1 III [Eva06] | B1 II | no | - |
4-049 | 4689002443498253440 | AzV 226, Cl* NGC 346 ELS 10 | 15.02 | O7 IIIn((f)) [Eva 06] | O7 IIIn((f)) | no | ((f)) from [Eva06] |
4-050 | 4690521487494893952 | Cl* NGC 346 ELS 103 | 16.91 | B0.5 V [Eva06] | B1 II: | no | - |
4-051 | 4688999140607234944 | - | 14.94 | - | B1 II | no | - |
4-052 | 4690521040818521600 | Cl* NGC 346 ELS 32 | 15.73 | B0.5 V [Eva06] | B1.5 III: | no | - |
4-053 | 4690504857379258496 | Cl* NGC 346 MPG 770 | 15.20 | B0 V [Duf19] | B0 V | no | - |
4-054 | 4688998938805087232 | AzV 231 | 14.08 | A0 [Azz75] | B9 Ib | no | - |
4-055 | 4690522273449262592 | LHA 115-S 29, RMC 15, Sk 79 | 12.10 | B 7I[e] [Gra12] | sgB[e] | yes | - |
4-056 | 4690517742283562624 | [BLK2010] flames1059 | 16.02 | B0 V [Duf19] | B0 V: e | no | SB1 [Duf19] |
4-057 | 4690504582501410432 | Cl* NGC 346 ELS 46 | 16.06 | O7 Vn [Eva06] | O6.5 Vnn | no | - |
4-058 | 4690516677131714432 | Sk 80, AzV 232 | 12.70 | O7 Iaf+ [Wal77] | O7 Iaf+ | no | from [Wal77], Binary [Eva06] |
4-059 | 4690503689148323328 | 2dFS 5100, Cl* NGC 346 ELS 18 | 15.22 | B0.5 Vn [Eva06] | B1: III: + early B | no | SB2 [Eva06], EB [Paw13] |
4-060 | 4690516711491448960 | Cl* NGC 346 MPG 802 | 14.45 | B8 Ib [Duf19] | B8 II-Ib | no | - |
4-061 | 4689002271699550720 | OGLE SMC-SC8 157867 | 16.08 | B2(shell) [Duf19] | B2 II | no | - |
4-062 | 4688998973164819200 | - | 16.27 | - | B2: III: + B2: | no | - |
4-063 | 4690504582501405696 | [M2002] SMC 46892 | 12.00 | A5 V [Neu10] | A7 Ib | no | - |
4-064 | 4688999248042623744 | - | 16.38 | - | B2 IV: | no | - |
4-065 | 4690517089448534272 | [BLK2010] flames1021 | 15.01 | B2 II [Duf19] | B2 II: | no | - |
4-066 | 4690518223319786624 | Sk 81, AzV 234 | 13.19 | B3 Iab [Duf19] | B2.5 Ib | no | - |
4-067 | 4690518154600349184 | 2dFS 1418, Cl* NGC 346 ELS 35 | 15.62 | B1 V [Eva06] | B2 II: | no | SB2 [Eva06] |
4-068 | 4690518154600325760 | [BLK2010] flames1029 | 15.30 | B2.5 III [Duf19] | B2 II: | no | - |
4-069 | 4690517879722483328 | - | 17.05 | - | B1 II neb | no | - |
4-070 | 4690516745851186048 | Cl* NGC 346 MPG 842 | 16.95 | B1 V [Duf19] | B2 II neb | no | - |
4-071 | 4690504926098661504 | [MA93] 1167, Cl* NGC 346 ELS 9 | 14.18 | B0e [Eva06] | O9.7 II: e? | no | - |
4-072 | 4690500184455105536 | RMC 16, Sk 83, AzV 237 | 12.32 | B9 Ia [Eva04] | B9 Ia | no | - |
4-073 | 4690504994818102016 | Cl* NGC 346 ELS 25, MPG 848 | 15.45 | O9 V [Eva06] | O9.2 V | no | SB1 [Eva06] |
4-074 | 4690518532557434496 | Cl* NGC 346 ELS 31 | 15.74 | O8 Vz [Eva06] | O9 V | no | - |
4-075 | 4690504651220842880 | Cl* NGC 346 MPG 859, 2dFS 5102 | 13.25 | A2 Ib [Duf19] | A0 Ib | no | - |
4-076 | 4690503792227414400 | AzV 238 | 14.24 | O9 III [Duf19] | O9.7 III | no | SB1 [Duf19] |
4-077 | 4690517055088782848 | [BLK2010] flames1014 | 14.88 | B3 Ib + mid A? [Duf19] | B3 Ib | no | SB2? [Duf19] |
4-078 | 4690503826587160832 | RMC 18, Sk 85, AzV 242 | 12.35 | B1 Ia [Len97] | B1 Ia | no | - |
4-079 | 4690521762372716160 | - | 13.10 | - | A0 Ib | no | - |
4-080 | 4690503929666296960 | MOA J010015.9-721244 | 14.81 | O8 + O9 [Hil05] | O9.7 + (O8-8.5 + B) | no | SB2 [Duf19] EB [Paw13] |
4-081 | 4690503860946856832 | [BLK2010] flames1091 | 16.28 | B0.5: + B0.5: [Duf19] | B0.5: V | no | SB2 [Duf19] |
4-082 | 4690517256934074624 | [BLK2010] flames1114 | 16.64 | O9 V [Duf19] | O9.5 IV:(n) neb | no | - |
4-083 | 4690500317576535040 | [M2002] SMC 48657 | 15.05 | - | B1 II: | no | - |
4-084 | 4690522002890799488 | Sk 86, AzV 250 | 12.57 | A0 :I [San68] | A0 Iab | no | - |
4-085 | 4690505097897766144 | Sk 87, AzV 252 | 13.27 | B2.5 Iab [Len97] | B2.5 Ib | no | - |
4-086 | 4690504341983110272 | OGLE SMC-SC9 41748 | 13.90 | A0 (Ib) [Eva04] | A2 II/Iab | no | - |
4-087 | 4690504341983098368 | - | 15.76 | - | B1 II: | no | - |
4-088 | 4690500734210820352 | - | 16.26 | - | B0.5: IV: | no | - |
4-089 | 4690500665491363456 | - | 15.85 | - | O9 V | no | - |
4-090 | 4690503998385814272 | OGLE SMC-SC9 41745 | 15.31 | B1-3 (II) [Eva04] | B1 II | no | - |
4-091 | 4690505475854319360 | - | 13.44 | - | B8 Ib/Iab | no | - |
4-092 | 4690518429478167680 | - | 16.52 | - | B2 III | no | - |
4-093 | 4690517604844488832 | - | 15.36 | - | B1 III: | no | - |
4-094 | 4690517329966645632 | - | 15.39 | - | B1 III | no | - |
4-095 | 4690505578937619712 | OGLE SMC-SC9 47454 | 16.17 | B1 + B1-2 [Hil05] | B2 III: | no | EB [Paw13] |
4-096 | 4690504170184413952 | [MA93] 1233 | 14.86 | B0-5 (II) [Eva04] | B1.5 II: | yes | - |
4-097 | 4690504204544113152 | - | 16.24 | - | B2 III: | no | - |
4-098 | 4690500802930242816 | OGLE J010103.75-721544.4 | 14.18 | - | B2.5 II: | no | - |
4-099 | 4690500802930226432 | [M2002] SMC 50390 | 15.38 | - | B0.7 III | no | - |
4-100 | 4690506712804884352 | OGLE SMC-SC9 44811 | 16.01 | B1-2 (III) [Eva04] | B1 III | no | - |
4-101 | 4690499978296571392 | OGLE SMC-SC9 38950 | 15.30 | B1-2 (II) [Eva04] | B1.5 II: | no | - |
4-102 | 4690500871649654656 | OGLE SMC-SC9 41776 | 16.16 | B0-5 (IV) [Eva04] | O7.5 V:n | no | - |
4-103 | 4690500047016009728 | AzV 80F | 12.20 | A3 V [Neu18] | A5 Ib | no | - |
4-104 | 4690518841795059840 | - | 15.15 | - | B1.5 II: | no | - |
4-105 | 4690501180887186944 | Sk 96, AzV 268, 2dFS 1550 | 13.56 | B2.5 Iab [Len97] | B2: Ib: | no | - |
4-106 | 4690507125121661696 | - | 16.43 | - | B2-2.5 III | no | - |
4-107 | 4690506747164583424 | - | 15.87 | - | B1 II | no | - |
4-108 | 4690505986933925888 | - | 15.99 | - | B1.5 III: | no | - |
4-109 | 4690502417837769344 | - | 16.18 | - | B0.7 III | no | - |
4-110 | 4690505819451697920 | - | 15.83 | - | O7 V:(n) | no | - |
4-111 | 4690518979233951104 | OGLE SMC113.6 22632 | 15.98 | - | B3: II | no | - |
4-112 | 4690501623246789376 | AzV 83F, 2dFS 1607 | 14.88 | B5 (II) [Eva04] | B5 II | no | - |
4-113 | 4690502761435017984 | OGLE SMC-SC9 131970 | 15.33 | B1-3 (III) [Eva04] | B2.5 II pe | no | - |
4-114 | 4690501352685999616 | AzV 291 | 14.92 | B1 [Azz75] | B5 II | no | - |
4-115 | 4690503208111546496 | - | 16.47 | - | B2 III: | no | - |
4-116 | 4690506266128219904 | - | 16.39 | - | B1.5 III: | no | - |
5-001 | 4687500991624952064 | OGLE SMC-SC10 24428 | 15.05 | O9 Ve [Lam16] | O9.7 Vn + B | no | LC from [Lam16] |
5-002 | 4687504324519462528 | OGLE SMC113.7 34708 | 15.94 | - | B2 IV | no | - |
5-003 | 4687504358879209088 | - | 16.81 | - | B2.5 III | no | - |
5-004 | 4687499922153760256 | - | 15.95 | - | B2.5 III-II | no | - |
5-005 | 4687504182822540416 | SV* HV 1954 | 12.16 | - | F2: | no | - |
5-006 | 4687487763123080704 | - | 15.94 | - | B2 II: | no | - |
5-007 | 4687499685954929664 | AzV 346 | 14.34 | B1-5 (II) [Eva04] | B2.5 II | no | - |
5-008 | 4687501537095707264 | - | 15.99 | - | B1 II: | no | - |
5-009 | 4687500063912170112 | OGLE SMC-SC10 63718 | 15.38 | B2.5 (III) [Eva04] | B2.5 II | no | - |
5-010 | 4687501468343335808 | - | 15.33 | - | B3 II | no | - |
5-011 | 4687487660043860096 | - | 16.10 | - | B2 III: | no | - |
5-012 | 4687499720314663296 | Dachs SMC 2-33 | 12.91 | - | A2 II/Ib | no | - |
5-013 | 4687487625684130688 | OGLE SMC-SC10 61625 | 16.25 | B3 (III) [Eva04] | B2 III | no | EB [Paw13] |
5-014 | 4687487900566335744 | Dachs SMC 2-34,AzV 352,2dFS 1898 | 12.65 | A0 (Iab) [Eva04] | A0 Ib | no | - |
5-015 | 4687505733268579968 | - | 15.92 | - | B2 III: | no | - |
5-016 | 4687486556213576960 | AzV 354, 2dFS 1904 | 15.18 | B1-3 (II) [Eva04] | B1.5 II | no | - |
5-017 | 4687501678819589888 | OGLE J010437.78-721352.7 | 16.18 | - | B0.2 IV | no | - |
5-018 | 4687505763370767104 | - | 14.62 | - | B3 II | no | - |
5-019 | 4687501678819562880 | - | 16.20 | - | B2 III | no | - |
5-020 | 4687486526172590720 | LIN 435, AzV 359, 2dFS 1922 | 14.17 | B0-3 (II)e [Eva04] | B1.5 II e | yes | - |
5-021 | 4687487999358589824 | Dachs SMC 2-41,AzV 360,2dFS 1923 | 15.04 | B0.5 (IV) [Eva04] | B1 II | no | - |
5-022 | 4687500579308038016 | 2dFS 1925 | 15.15 | B0-5 (II) [Eva04] | B3 II | no | - |
5-023 | 4687487247726980864 | AzV 363 | 14.88 | B0.5 III [Mas00] | B1.5 III e | yes | - |
5-024 | 4687483502515662336 | AzV 365, 2dFS 1934 | 15.02 | B0 (IV) [Eva04] | B0.7 III | no | - |
5-025 | 4687488102403240320 | - | 16.93 | - | B1 III | no | - |
5-026 | 4687500304430157312 | AzV 368, 2dFS 1939 | 15.20 | B0-3 (II) [Eva04] | B1 II | no | - |
5-027 | 4687487282086698240 | AzV 370 | 15.46 | B0 V: [Cra82] | B1.5 III: | no | - |
5-028 | 4687501640142199936 | - | 16.66 | - | B0 V | no | - |
5-029 | 4687487213367243392 | Dachs SMC 2-44, 2dFS 1949 | 12.35 | A5 II [Eva04] | A7 Ib | no | - |
5-030 | 4687501365297007232 | OGLE SMC-ECL- 7623 | 15.69 | - | B1.5 III: | no | - |
5-031 | 4687487007208845056 | LIN 440 | 14.19 | - | B2.5: II:e + early AIa | yes | - |
5-032 | 4687500270070426496 | 2dFS 1967 | 15.77 | B1-3 (III) [Eva04] | B2.5 II | no | - |
5-033 | 4687487385165885568 | - | 16.82 | - | B1 III | no | - |
5-034 | 4687486938489379328 | LIN 444, 2dFS 1977 | 14.82 | B1-5 (II) [Eva04] | B1.5 II: e | yes | - |
5-035 | 4687482780961207424 | [MA93] 1525, 2dFS 1982 | 14.06 | B0-5 (II) [Eva04] | B1.5 II: e | no | - |
5-036 | 4687482884040407936 | 2dFS 1997 | 14.32 | A0 (Ib) [Eva04] | B9 Ib | no | - |
5-037 | 4687487144647743616 | OGLE SMC-ECL- 7650 | 15.65 | - | B0.5 V: + early B | no | - |
5-038 | 4687483738714882304 | [M2002] SMC 61844 | 15.18 | - | B1.5 III-II | no | - |
5-039 | 4687487419525570816 | - | 16.03 | - | B1.5 III | no | - |
5-040 | 4687489927786395776 | AzV 381 | 14.86 | B0 IIw [Gar87] | B1 II | no | EB (Paw13] |
5-041 | 4687482815320929152 | AzV 383, 2dFS 2018 | 14.71 | B1-5 (Ib) [Eva04] | B1 II | no | - |
5-042 | 4687482609162518784 | 2dFS 2023 | 15.50 | O9.5 III [Eva04] | O9.7 III(n) | no | - |
5-043 | 4687482712241716608 | nan | 14.64 | B1-2 (II) [Eva04] | B1.5 II | no | - |
5-044 | 4687488896994307712 | - | 15.60 | O9.5 V [Lam13] | O9.5 IV | no | - |
5-045 | 4687501884978067584 | AzV 387 | 14.62 | O9.5 III [Eva04] | O9.2 III + early B | no | LC from [Eva04] |
5-046 | 4687483777393718400 | LHA 115-S 44 | 13.56 | B2 Ib [Dac70] | B1.5 II: e | yes | - |
5-047 | 4687502228575362688 | [MA93] 1545, 2dFS 2055 | 14.96 | B1-5 (II) [Eva04] | B2.5 III: e | yes | - |
5-048 | 4687482952759849728 | LIN 446 | 13.87 | B0-3 [Pau12] | O9.7 V(n) | no | - |
5-049 | 4687435708116998400 | OGLE SMC110.4 12773 | 16.53 | - | B0.2 IV | no | - |
5-050 | 4687501953697501568 | AzV 389 | 14.33 | B0 IIw [Gar87] | O9.7 V: + early B | no | - |
5-051 | 4687501884978054784 | - | 17.11 | - | B0.2 III | no | - |
5-052 | 4687488862634574848 | AzV 104F | 13.09 | B8 [Azz79] | B9 Iab | no | - |
5-053 | 4687485319263467520 | 2dFS 2068 | 16.24 | B0-5 (III) [Eva04] | B3 II: | no | - |
5-054 | 4687488587757077504 | 2dFS 2069 | 13.25 | A3 II [Eva04] | A5 Ib | no | - |
5-055 | 4687482746601417216 | 2dFS 2070 | 14.02 | A0 (Ib) [Eva04] | A1 II/Ib | no | - |
5-056 | 4687501988057218688 | 2dFS 2074 | 15.42 | B2.5 (III) [Eva04] | B2 II | no | - |
5-057 | 4687488656476146432 | 2dFS 2077 | 15.27 | B0-5 (III) [Eva04] | B1 III: + early B | no | - |
5-058 | 4687490168304521344 | 2dFS 2085 | 15.91 | B1-5 (II) [Eva04] | B1 III-II | no | - |
5-059 | 4687484498948014464 | - | 16.45 | - | B0.7 IV | no | - |
5-060 | 4687502572172727808 | 2dFS 2086 | 14.93 | B1-3 (II) [Eva04] | B1.5 II | no | - |
5-061 | 4687489790347907328 | - | 16.93 | - | B2 III: | no | - |
5-062 | 4687502640892163456 | 2dFS 2090 | 15.74 | B0.5 (V) [Eva04] | B1.5 + early B | no | - |
5-063 | 4687488690835887360 | 2dFS 2098 | 15.73 | B0 (V) [Eva04] | B0 V: + early B | no | - |
5-064 | 4687485564099788544 | AzV 394, 2dFS 2097 | 14.04 | B1-5 (Iab) [Eva04] | B1.5 II | no | - |
5-065 | 4687502366014559232 | AzV 396 | 14.96 | B0 III [Mas04] | B0 IV | no | - |
5-066 | 4687489824707171456 | LHA 115-S 47, AzV 397 | 13.56 | B1-2 (II) [Eva04] | B2 Ib e | yes | - |
5-067 | 4687485387976612736 | Sk 126, AzV 399 | 12.21 | A0 I [Hum83] | A0 Iab | no | - |
5-068 | 4687489000073470720 | - | 16.14 | - | B1 III-II | no | - |
5-069 | 4687502400374028928 | LIN 455, AzV 400, 2dFS 2127 | 14.03 | B2.5 (II) [Eva04] | B3 II e | yes | - |
5-070 | 4687484636386901760 | - | 15.93 | - | B2 III: | no | - |
5-071 | 4687437254305162880 | HA 115-S 49, AzV 402, 2dFS 2139 | 13.54 | O9.7 Ib [Eva04] | O8.5: Ib + OB pe | yes | LC from [Eva04] |
5-072 | 4687484361509061504 | [M2002] SMC 63718 | 14.74 | - | B1 II | no | - |
5-073 | 4687490305743468416 | 2dFS 2148 | 15.72 | B1-5 (II)e? [Eva04] | B2 II: | no | - |
5-074 | 4687502331654573056 | 2dFS 2151 | 15.89 | B0.5 (V) [Eva04] | B0.7 IV | no | - |
5-075 | 4687502434733757440 | SV* HV 2016 | 15.58 | B0.5+B2 [Har03] | B2 II: | no | EB [Paw13], SB2? |
5-076 | 4687502709611614336 | - | 16.97 | - | B1.5 III: | no | - |
5-077 | 4687485491055825536 | Sk 128, AzV 404, 2dFS 2174 | 12.47 | B2.5 Iab [Len97] | B2.5 Ia | no | - |
5-078 | 4687502434733761280 | - | 15.48 | - | B1.5 III: | no | - |
5-079 | 4687490305743456000 | - | 16.53 | - | B2 III: | no | - |
5-080 | 4687485873337367936 | - | 16.09 | - | B2 III: | no | - |
5-081 | 4687502537812927744 | - | 17.02 | - | B0.7 III | no | - |
5-082 | 4687489206231868800 | - | 17.12 | - | B1.5 III | no | - |
5-083 | 4687489584188971392 | [M2002] SMC 64573 | 16.47 | - | B1 II | no | - |
5-084 | 4687484773825841408 | 2dFS 2211 | 15.47 | B1-3 (III) [Eva04] | B2 II: | no | - |
5-085 | 4687489579873557120 | - | 16.30 | - | B0.7 III | no | - |
5-086 | 4687486182574932352 | 2dFS 2215 | 14.47 | B9 (Ib) [Eva04] | B9 Ib | no | - |
5-087 | 4687489240591624832 | 2dFS 2216 | 15.56 | B2 (II) [Eva04] | B1 II | no | - |
5-088 | 4687489579867156864 | - | 15.33 | - | B1.5 III-II | no | - |
5-089 | 4687437494823799168 | - | 16.07 | - | B2 III: e | yes | - |
5-090 | 4687513910877142528 | AzV 411, 2dFS 2238 | 14.53 | O9 III [Eva04] | O9.5 III | no | - |
5-091 | 4687437426103845376 | AzV 412, 2dFS 2243 | 13.11 | B8 (Iab) [Eva04] | B8 Ib | no | - |
5-092 | 4687437529183051136 | 2dFS 2249 | 14.04 | B0-5 (II) [Eva04] | B1 II: pe | no | - |
5-093 | 4687437907140147712 | 2dFS 2257 | 14.98 | B1-5 (II) [Eva04] | B1 II | no | - |
5-094 | 4687485976416952832 | LIN 468, AzV 413 | 14.74 | Be [Gar85] | B1 Ib: e | yes | - |
5-095 | 4687436700232929536 | [M2002] SMC 65523 | 15.95 | - | B1.5 III: | no | - |
5-096 | 4687437834104778624 | [MA93] 1620, 2dFS 2265 | 15.06 | B1-3 (II) [Eva04] | B1.5 III-II | yes | - |
5-097 | 4687513876517417856 | 2dFS 2266 | 16.02 | OC7 II(f) [Eva04] | O8 II(f) | no | LC from [Eva04] |
5-098 | 4687437632262257024 | AzV 414 | 13.13 | B8 [Azz75] | B9 Ia | no | - |
5-099 | 4687512845725520640 | 2dFS 2281 | 16.24 | O9.5 III [Eva04] | O9.7 III | no | LC from [Eva04] |
5-100 | 4687514048316067968 | 2dFS 2283 | 15.42 | B0.5 (IV) [Eva04] | B0.2 V | no | EB [Paw13] |
5-101 | 4687437563542778880 | - | 12.98 | - | A5 Ib | no | - |
5-102 | 4687508413319210368 | AzV 417, 2dFS 2289 | 12.63 | F5 [Eva04] | F0: | no | - |
5-103 | 4687509753348904832 | AzV 418 | 15.04 | B2: [Azz75] | B3 II | no | - |
5-104 | 4687512948804514176 | AzV 419 | 13.54 | B2 [Azz75] | B2.5 Ia | yes | - |
5-105 | 4687509749039520768 | Sk 131, AzV 420 | 13.76 | B0.5 Ia [Len97] | B0.7 II | no | - |
5-106 | 4687508344599737216 | AzV 421, 2dFS 2299 | 15.14 | B2 (II) [Eva04] | B2 II | no | - |
5-107 | 4687512880085044608 | LIN 469, AzV 422, 2dFS 2316 | 14.30 | B1-5 (II)e [Eva04] | B2.5 II + early A | yes | - |
5-108 | 4687437597895960448 | 2dFS 2317 | 16.42 | B0 (V) [Eva04] | O9.7 IV: | no | - |
5-109 | 4687513567279735552 | 2dFS 2340 | 12.19 | F0 ([Eva04] | A7 Iab | no | - |
5-110 | 4687513361121326080 | - | 16.82 | - | B1 III | no | - |
5-111 | 4687508825635978752 | OGLE SMC-SC11 89163 | 16.50 | - | B2 III: | no | EB [Paw13] |
5-112 | 4687461099956943744 | 2dFS 2350 | 14.46 | B8 (II) [Eva04] | A0 Ib | no | - |
5-113 | 4687508550758084864 | - | 14.24 | - | B1 II | no | - |
5-114 | 4687508585117812736 | AzV 432 | 14.43 | B0 [Azz75] | B1.5 II | no | - |
5-115 | 4687460481482120832 | LHA 115-S 54, AzV 436, 2dFS 2413 | 13.40 | B0 (II)e [Eva04] | O9.5: V: pe | yes | - |
5-116 | 4687461649712663296 | - | 16.24 | - | B2 II | no | - |
6-001 | 4687162582543422080 | [MA93] 1760 | 14.53 | B0 Ve [Lam13] | B0.5 III: e | yes | - |
6-002 | 4687161929708394752 | Dachs SMC 3-1, AzV 122F, 2dFS 2876 | 12.47 | A0 (Iab) [Eva04] | A2 Ib | no | - |
6-003 | 4687161654830500480 | 2MASS J01124381-7318457, 2dFS 2877 | 15.91 | B2 (III) [Eva04] | B1.5 III-II | no | - |
6-004 | 4687161860988920960 | 2MASS J01124660-7317497, 2dFS 2883 | 15.66 | B2 (III) [Eva04] | B2 III: | no | - |
6-005 | 4687161895348646400 | Sk 151, AzV 471, 2dFS 2905 | 13.87 | O9.5 I [Lam16] | O9.7 II-Ib(n) | no | - |
6-006 | 4687158592506675328 | Dachs SMC 3-4, AzV 473 | 11.20 | F8 I [Neu10] | F2: | no | - |
6-007 | 4687157286836619520 | 2MASS J01132216-7324067, 2dFS 2937 | 15.24 | B3 (II) [Eva04] | B5 II | no | - |
6-008 | 4687159868124113152 | HD 7583, RMC 45, Sk 152, AzV 475 | 9.62 | A0 Ia-0 [Fea60] | A2 Ia | no | - |
6-009 | 4687163235378435328 | Dachs SMC 3-9 | 12.89 | A3 V: [Dac70] | A2 II/Ib | no | - |
6-010 | 4687158768612502528 | OGLE SMC116.7 9681 | 16.73 | - | B2 IV: | no | - |
6-011 | 4687160280440948608 | - | 15.36 | - | B2 III: e | no | - |
6-012 | 4687160280440940672 | - | 15.88 | - | B1 III neb | no | - |
6-013 | 4687160207423517312 | WBBe NGC 456 2 | 14.65 | O6-7 V [Tes87] | O7.5 V: + O9.5 neb | no | - |
6-014 | 4687160211708489088 | - | 15.18 | - | O9.5 V(n) neb | no | - |
6-015 | 4686408111398825728 | [BKB2004b] 182 | 11.74 | F0 I [Neu18] | F2: | no | - |
6-016 | 4687160246081207680 | OGLE SMC116.6 11072 | 15.40 | - | B0.7 IV neb | no | - |
6-017 | 4687165125163949952 | Sk 153, AzV 477, 2dFS 2980 | 13.50 | B2 (Ib) [Eva04] | B1 Ib | no | - |
6-018 | 4687176979273664512 | 2dFS 2983 | 16.97 | B0-5 (IV) [Eva04] | B2 IV-III | no | - |
6-019 | 4687160246081200256 | LIN 503 | 15.06 | - | O9.7 V neb | no | - |
6-020 | 4687164128731574400 | 2MASS J01141596-7311537, 2dFS 2995 | 15.73 | B2.5 (III) [Eva04] | B2.5 III | no | - |
6-021 | 4687164060012111616 | 2MASS J01141722-7313014, 2dFS 2998 | 15.93 | O9.5 III [Eva04] | B0 V-IV | no | - |
6-022 | 4687183159725860736 | - | 17.00 | - | B2 III-II | no | - |
6-023 | 4687160727117511168 | OGLE SMC116.3 37 | 15.96 | - | O9.2 V | no | - |
6-024 | 4687164055705068288 | Sk 154, AzV 478 | 11.09 | A2-3 I [Hum91] | A5 Iab | no | - |
6-025 | 4687163853853686144 | GSC 09142-00856, 2dFS 3006 | 15.27 | O8.5 V [Eva04] | O9.2 V | no | - |
6-026 | 4687159593246187136 | - | 16.58 | - | B1 III neb | no | - |
6-027 | 4687164919005535616 | 2dFS 3010 | 15.90 | B1-3 (III) [Eva04] | B1.5 III | no | - |
6-028 | 4687159558886453760 | - | 16.29 | - | B0.5 V | no | - |
6-029 | 4687160486599363712 | 2dFS 3023 | 16.87 | O9.5 III [Eva04] | B0.5 III | no | - |
6-030 | 4687163819493940992 | - | 16.46 | - | B0 IV: | no | - |
6-031 | 4687159661965655552 | OGLE SMC733.30.000053 | 15.48 | - | O8.5 V: | no | EB [Paw16] |
6-032 | 4687159455807235072 | - | 16.35 | - | O9 V:(n) neb | no | - |
6-033 | 4687159657667699712 | [MA93] 1794 | 15.21 | - | O4.5 V: neb | no | - |
6-034 | 4686408313260119552 | 2dFS 3036 | 15.23 | B1-3 (III) [Eva04] | B2 III-II e | yes | EB [Paw13] |
6-035 | 4687171206837590272 | [M2002] SMC 75626 | 15.81 | O9 III-V [Lam16] | B0 IV | no | - |
6-036 | 4686408691230920448 | - | 15.25 | - | B0 V neb | no | - |
6-037 | 4687159387087763968 | - | 16.35 | - | B2 III: neb | no | - |
6-038 | 4686408764233799168 | OGLE SMC-ECL- 5820 | 16.60 | - | B2 IV: neb | no | EB [Paw13] |
6-039 | 4686408764233794816 | - | 16.69 | - | B2 III: neb | no | - |
6-040 | 4687159421447500160 | - | 16.64 | - | B1.5 III: neb | no | - |
6-041 | 4686408729874075008 | 2dFS 3066 | 16.70 | B1-2 (IV) [Eva04] | B1.5 III: | no | |
6-042 | 4686408729874061952 | AzV 482, 2dFS 3077 | 15.05 | B1-3 (II) [Eva04] | B0.7 III | no | - |
6-043 | 4686407389844312704 | - | 16.01 | - | B2 III | no | - |
6-044 | 4686414674108779008 | - | 16.13 | - | B1.5 IV-III neb | - | - |
6-045 | 4687170966319428992 | 2dFS 3082 | 15.43 | B1-2 (III) [Eva04] | B1 III-II | no | - |
6-046 | 4687169935527321088 | 2dFS 3083 | 16.83 | B0.5 (V) [Eva04] | B1 III | no | - |
6-047 | 4687170141685733376 | 2dFS 3086 | 16.23 | B1-3 (III) [Eva04] | B1.5 III neb | no | - |
6-048 | 4686414399230889600 | - | 16.75 | - | B0.7 III | no | - |
6-049 | 4687170588362310912 | 2dFS 3088 | 16.74 | B1-5 (III) | B2 III: | no | EB [Paw13] |
6-050 | 4686414708468514688 | [M2002] SMC 76150 | 15.35 | - | B1 III | no | - |
6-051 | 4687169862500848256 | 2dFS 3092 | 15.93 | B0.5 (V) [Eva04] | B0.7 IV | no | - |
6-052 | 4687165709279578368 | OM 3, 2dFS 3094 | 12.25 | A3 II [Eva94] | A2 II/Ib | no | - |
6-053 | 4687170588362312192 | 2dFS 3095 | 17.04 | B0-5 (IV) [Eva04] | B2 III: | no | - |
6-054 | 4686414429294260608 | - | 15.27 | - | B0.7 IV | no | - |
6-055 | 4687166843150898304 | 2dFS 3102 | 16.79 | B1-5 (III) [Eva04] | B2 III: | no | - |
6-056 | 4687165709279571968 | 2dFS 3103 | 16.04 | B0 (V) [Eva04] | B0.7 III | no | - |
6-057 | 4686414330511426048 | - | 16.06 | - | B0.7 V | no | - |
6-058 | 4687170244764966912 | - | 16.42 | - | B1 III | no | - |
6-059 | 4687170244764958336 | - | 16.14 | - | B2 III: | no | - |
6-060 | 4687165503121154944 | 2dFS 3106 | 16.32 | O9 V [Eva04] | O9.7 IV | no | - |
6-061 | 4687166843150894208 | 2dFS 3107 | 15.09 | B0 (IV) [Eva04] | B0 V | no | EB [Paw13] |
6-062 | 4686413505877718784 | 2dFS 3108 | 16.19 | B1-3 (III) [Eva04] | B2 III-IV | no | EB [Paw13] |
6-063 | 4686413230999833856 | 2dFS 3109 | 16.51 | B0-5 (IV) [Eva04] | B2 III: | no | - |
6-064 | 4687166636992467328 | [M2002] SMC 76350 | 15.98 | - | B0 V | no | - |
6-065 | 4686414364871151360 | - | 16.84 | - | B1.5 III: | no | - |
6-066 | 4687165812358767104 | 2dFS 3117 | 16.43 | B0 (V) [Eva04] | B0.7 IV | no | - |
6-067 | 4686413230999834496 | AzV 484, 2dFS 3118 | 14.82 | O9.5 III [Eva04] | O9.7 III | no | - |
6-068 | 4686414742828249472 | 2dFS 3120 | 15.77 | B0 (V) [Eva04] | B0.2 IV: | no | - |
6-069 | 4686413608956928640 | 2dFS 3121 | 16.43 | B0-3 (IV) [Eva04] | B1 II | no | - |
6-070 | 4686414364871146624 | 2dFS 3130 | 14.77 | B1-2 (II) [Eva04] | B1: II | no | EB [Paw13] |
6-071 | 4686414467950361344 | - | 15.52 | - | B1.5 III: | no | - |
6-072 | 4686413540237697408 | 2dFS 3134 | 14.59 | B3 (II) [Eva04] | B3 II | no | - |
6-073 | 4686413540237697152 | - | 16.66 | - | B2 III-II | no | - |
6-074 | 4686414532373479680 | - | 16.46 | - | B1.5 III neb | no | - |
6-075 | 4686413712036121088 | - | 16.07 | - | B0.7: V | no | - |
6-076 | 4686414467950355072 | Sk 158, AzV 487 | 13.18 | BC0 Ia [Len97] | B0.2 Ib | no | - |
6-077 | 4686413712036117376 | - | 16.49 | - | B2 III: | no | - |
6-078 | 4687165915437998848 | 2dFS 3150 | 16.72 | B0-5 (IV) [Eva04] | B2 II | no | - |
6-079 | 4686413712029003904 | 2dFS 3152 | 16.05 | B1-3 (III) [Eva04] | B0.7 III | no | - |
6-080 | 4686413437158238208 | Sk 159, AzV 488 | 12.22 | B0.5 Iaw [Wal83] | B0.2 Ia | no | - |
6-081 | 4686414566732495872 | 2dFS 3165 | 16.37 | B0-3 (IV) [Eva04] | B1.5 IV: e | no | - |
6-082 | 4686413677676380672 | - | 17.05 | - | B1.5 IV-III | no | - |
6-083 | 4687165949797725696 | - | 16.57 | - | B1 III: | no | - |
6-084 | 4687167152388524032 | 2dFS 3170 | 14.35 | O9.5 Ib [Eva04] | O9.7 I:(n) | no | - |
6-085 | 4686413402798497920 | - | 16.87 | - | B1.5 IV | no | - |
6-086 | 4687166323447798016 | - | 16.07 | - | B1.5 III: | no | - |
6-087 | 4687172271989430912 | 2dFS 3187 | 16.22 | B1-3 (III) [Eva04] | B1.5 III-II | no | - |
6-088 | 4686413398501396736 | 2dFS 3188 | 16.30 | B1-5 (III) [Eva04] | B1.5 III-II | no | - |
6-089 | 4686413849475073920 | 2dFS 3191 | 16.33 | B0-3 (IV) [Eva04] | B0 V: | no | - |
6-090 | 4687167186748236544 | 2dFS 3192 | 16.33 | B1-3 (IV) [Eva04] | B2 III | no | - |
6-091 | 4686415361303520512 | [M2002] SMC 76980 | 15.88 | - | O9.5 IV: | no | - |
6-092 | 4687168870375394944 | 2dFS 3193 | 16.30 | B1-3 (III) [Eva04] | B1 III | no | EB [Paw13] |
6-093 | 4686413810818254464 | 2dFS 3199 | 14.58 | O9 V [Eva04] | O8.5 V | no | - |
6-094 | 4686413879538156544 | 2dFS 3202 | 16.70 | B1-5 (III) [Eva04] | B2 III: | no | - |
6-095 | 4686410550940229888 | - | 13.82 | - | A0 Ib | no | - |
6-096 | 4686414193072434944 | 2dFS 3208 | 16.86 | B1-5 (IV) [Eva04] | B2 IV | no | - |
6-097 | 4687172100190744448 | - | 14.89 | - | B8 II-Ib | no | - |
6-098 | 4686410787161979008 | - | 15.81 | - | O8.5 V: | no | - |
6-099 | 4686414261791903360 | 2dFS 3213 | 17.00 | B2 (IV) [Eva04] | B2 II | no | - |
6-100 | 4687168629857232000 | [M2002] SMC 77248 | 15.42 | B0 [She13] | B1 II | no | - |
6-101 | 4687167873943001600 | 2dFS 3216 | 16.18 | B0 (V) [Eva04] | B0 V | no | - |
6-102 | 4686416873131995776 | 2dFS 3217 | 16.26 | B1-5 (III) [Eva04] | B2 II | no | - |
6-103 | 4687167873943006208 | - | 16.59 | - | B1 III: | no | - |
6-104 | 4687168732936450048 | 2dFS 3222 | 16.58 | O9.5 V [Eva04] | B0 IV | no | - |
6-105 | 4686410997616777984 | 2dFS 3225 | 14.70 | O6.5 V [Eva04] | O6 V:n | no | - |
6-106 | 4687167908302747136 | AzV 489 | 15.35 | O8.5 V [Mas95] | O9.7 V | no | - |
6-107 | 4686416907491729664 | - | 16.21 | - | O8.5 V: | no | - |
6-108 | 4686410688379157888 | - | 13.05 | - | A0 Iab | no | - |
6-109 | 4687167633424856320 | OGLE SMC-ECL- 7941 | 16.54 | nan | B2 III: | no | EB [Paw16] |
6-110 | 4686416941851468160 | 2dFS 3248 | 16.05 | B1-3 (III) [Eva04] | B1.5 III: | no | - |
6-111 | 4686416941851466880 | Sk 161, AzV 492, 2dFS 3252 | 13.12 | B0.5 I [Lam13] | B0.5 Ib | no | - |
6-112 | 4686412475085517440 | - | 16.42 | - | B1 III-II | no | - |
6-113 | 4686415430023013760 | AzV 493 | 13.78 | Ope pec [Lam16] | Onnpe | no | - |
6-114 | 4686416976211204992 | 2dFS 3267 | 16.84 | B1-5 (III) [Eva04] | B2 IV-III | no | - |
6-115 | 4686415842339856256 | 2dFS 3269 | 15.40 | B0-3 (III) [Eva04] | B2 III-II | no | EB [Paw13] |
6-116 | 4686417148009897344 | 2dFS 3274 | 12.80 | A7 II [Eva04] | A7 Iab | no | - |
7-001 | 4685992530438263808 | AzV 186, CL* NGC 330 ELS 13 | 14.52 | O8.5 III((f)) [Mas95] | O8.5 III((f)) | no | LC from [Mas95] |
7-002 | 4685991224768443392 | - | 15.30 | - | B3 II: | no | - |
7-003 | 4685988196744734080 | - | 16.51 | - | B1.5 III-II | no | - |
7-004 | 4685987857514223744 | AzV 195 | 13.51 | A0 [Azz75] | A0 Iab | no | - |
7-005 | 4685988132392086016 | - | 16.04 | - | B2 III-II | no | - |
7-006 | 4688995468472128128 | 2dFS 1268 | 15.17 | B2 (II) [Eva04] | B1 II | no | - |
7-007 | 4685987960593409024 | [MA93] 1042 | 15.06 | - | B1 II neb | no | - |
7-008 | 4685992663601032576 | TIC 181882730 | 15.06 | - | B2 II: | yes | - |
7-009 | 4685987960593372928 | - | 16.68 | - | B1.5 III neb | no | - |
7-010 | 4685991838904940928 | [VA82] II-8 | 15.84 | - | B1 III | no | - |
7-011 | 4688995640270760960 | Cl* NGC 330 ELS 53 | 16.07 | B0.5 V [Eva04] | B1 III | no | - |
7-012 | 4685991632807805312 | - | 14.61 | - | B5 II | - | - |
7-013 | 4685988544708868864 | LIN 329, 2dFS 1283 | 14.33 | B1-3 (II) [Eva04] | B1.5 III e | yes | - |
7-014 | 4685988265464182528 | - | 16.36 | - | B1 III | no | - |
7-015 | 4685986792362398976 | 2dFS 1288 | 14.88 | B1-2 (II) [Eva04] | B1 II | no | - |
7-016 | 4685991602725468800 | - | 16.18 | - | B1 III: | no | - |
7-017 | 4685991602725840128 | - | 13.46 | - | A2 Ib | no | - |
7-018 | 4685987513916795776 | [M2002] SMC 42648 | 15.15 | O9 [She13] | B0 IV | no | - |
7-019 | 4685986792362380928 | [M2002] SMC 42686 | 15.64 | - | B1.5: II | no | - |
7-020 | 4685988304190737536 | - | 15.27 | - | O7.5 V: neb | no | - |
7-021 | 4685987165948226432 | - | 16.55 | - | B1.5 III | no | - |
7-022 | 4685988647787974784 | - | 15.89 | - | B1 II: | no | - |
7-023 | 4685991705804639360 | OGLE SMC-SC8 90998 | 16.14 | nan | B1.5 III-II | no | EB [Paw13] |
7-024 | 4685993183273028224 | - | 16.04 | - | O9.7 V:n | no | - |
7-025 | 4685987475279399936 | - | 15.46 | - | O7.5 V:n neb | no | - |
7-026 | 4685987479557047808 | - | 16.23 | - | B1 II | no | - |
7-027 | 4685975350565410048 | 2dFS 1313 | 15.33 | B0 (IV) [Eva04] | O9.7 Vn | no | - |
7-028 | 4685987651355702528 | - | 16.08 | - | B1 III: | no | - |
7-029 | 4685975453644517248 | GEN# +6.20138214, [VA82] II-14 | 15.32 | - | B2 II | no | - |
7-030 | 4685975110047325568 | - | 16.57 | - | B0.5: V | no | - |
7-031 | 4685994695101439360 | - | 15.93 | - | B1.5 III | no | - |
7-032 | 4685975488004257920 | OGLE SMC105.3 89 | 16.39 | - | B2 III: | no | - |
7-033 | 4685975110047300352 | [M2002] SMC 44302 | 15.43 | - | B0.7 III | no | - |
7-034 | 4685993217632942720 | [M2002] SMC 44634 | 16.02 | - | B0 V | no | - |
7-035 | 4685975213126487424 | AzV 212, 2dFS 1348 | 15.15 | B0.5 (IV) [Eva04] | B1 II: | no | - |
7-036 | 4685987410837520256 | [VA82] II-16 | 15.96 | - | B1.5 III-II | no | - |
7-037 | 4685974457212254336 | OGLE J005854.41-724352.3 | 15.52 | - | B1.5 III: e | no | - |
7-038 | 4685975213126426368 | OGLE J005855.40-724300.8 | 14.76 | - | B0 II: e | yes | - |
7-039 | 4685974388492830080 | - | 16.19 | - | B1 IV: | no | - |
7-040 | 4685988888306231936 | - | 15.90 | - | B1.5 III-II | no | - |
7-041 | 4685994454583299840 | - | 15.59 | B1 Ve [Mas95] | B2 II: | no | - |
7-042 | 4685974491571903744 | - | 16.52 | - | B1.5 III-II | no | - |
7-043 | 4685993423791257984 | Sk 77, AzV 221 | 13.67 | B0 Ia [Gar87] | B1.5 Ib | no | - |
7-044 | 4685988819586784896 | AzV 223 | 14.33 | O9.5 II [Mas09] | O9.7 III | no | - |
7-045 | 4685993419427102848 | - | 17.01 | - | B2 III: e | no | - |
7-046 | 4685993733028817664 | AzV 225, 2dFS 1376 | 14.14 | B8 (Ib) [Eva04] | B5 II | no | - |
7-047 | 4685975625443431808 | AzV 228, 2dFS 1384 | 14.49 | B0.5 (III) [Eva04] | B0.2 III: pe | yes | - |
7-048 | 4685990056537133824 | - | 15.92 | - | B2 III | no | - |
7-049 | 4685975964679436672 | [M2002] SMC 46115 | 15.05 | - | B0.5 III | no | - |
7-050 | 4685994110985897472 | - | 16.88 | - | B2 III: | no | - |
7-051 | 4685990434494050048 | [M2002] SMC 46241 | 15.33 | B1 V [Lam16] | B1 Ib: e | no | - |
7-052 | 4685976759314509824 | - | 16.05 | - | B1.5 III-II | no | - |
7-053 | 4685972597426070016 | - | 16.47 | - | B0.2 IV | no | - |
7-054 | 4685975694162792448 | 2dFS 1392 | 16.03 | B1-2 (IV) [Eva04] | B2 III: | no | EB [Paw13] |
7-055 | 4685972322546968448 | - | 16.63 | - | B0.2 V | no | - |
7-056 | 4685971566632674304 | - | 16.96 | - | B1.5 III | no | - |
7-057 | 4685972391266444800 | OGLE SMC-SC8 129157 | 14.90 | - | B0.2 V | no | - |
7-058 | 4685975934680882304 | AzV 233 | 15.18 | B1 V [Mas95] | B1 II: | no | - |
7-059 | 4685972670505925888 | - | 14.90 | - | B3 II | no | - |
7-060 | 4685972601786502784 | - | 16.04 | - | B2 III: | no | - |
7-061 | 4685976926752244480 | - | 16.57 | - | B2: II: | no | - |
7-062 | 4685972631870071808 | - | 16.01 | - | B2 IV | no | - |
7-063 | 4685972636146192128 | OGLE SMC-SC8 132562 | 15.25 | O9.5 III [Eva04] | O9.5 III(n) | no | LC from [Eva04] |
7-064 | 4685972636146149504 | RMC 17, Sk 82, AzV 235 | 12.25 | B0 Ia [Wal83] | B0 Ia | yes | - |
7-065 | 4685975762891732096 | AzV 236 | 15.27 | B0 III [Mas95] | B1 II | no | EB [Paw13] |
7-066 | 4685976067758705280 | 2dFS 1430 | 16.38 | B0.5 (V) [Eva04] | B1 II | no | - |
7-067 | 4685993904827461504 | AzV 74F, 2dFS 1434 | 13.39 | A0 (Ib) [Eva04] | A0 Ib | no | - |
7-068 | 4685993904827455616 | - | 12.59 | - | A5 Ib | no | - |
7-069 | 4685971639713881344 | Sk 84, AzV 243 | 14.50 | O6 V [Mok06] | O6.5 V | no | LC from [Mok06] |
7-070 | 4685989850378489472 | - | 16.46 | - | B1.5 III: | no | - |
7-071 | 4685976209558662016 | AzV 249, 2dFS 1458 | 14.22 | B0 (III) [Eva04] | B0 II: + B0 | no | - |
7-072 | 4685990709381204864 | AzV 251 | 15.39 | O7 Vn [Cra82] | O8 Vnn | no | - |
7-073 | 4685989816018758272 | - | 15.34 | - | B2 II: | no | - |
7-074 | 4685976656235152640 | 2dFS 1471 | 15.17 | B0-5 (III) [Eva04] | B2 III: | no | - |
7-075 | 4685976621875402368 | - | 12.99 | - | B8 Ib | no | - |
7-076 | 4687491886297246336 | AzV 256, 2dFS 1497 | 15.33 | B0 (IV) [Eva04] | O9.7 I(n) | yes | - |
7-077 | 4687491126099277696 | 2dFS 1503 | 16.29 | B1-5 (III) [Eva04] | B2 II: | no | - |
7-078 | 4687491886293188608 | AzV 258 | 13.46 | O9 e [Azz75] | B0 e | yes | - |
7-079 | 4685976346997537536 | - | 16.13 | - | B2 III: | no | - |
7-080 | 4685976621876337664 | - | 16.36 | - | B1.5 III: | no | - |
7-081 | 4685924803028734080 | OGLE SMC-SC9 10098 | 16.10 | O9 + B0 [Hil05] | O9.7 Vnn | no | EB [Paw13] |
7-082 | 4687491164738701056 | AzV 261, 2dFS 1527 | 13.75 | B2 (Ib)e [Eva04] | B2 Ib: e | yes | - |
7-083 | 4687490958580285440 | 2dFS 1528 | 15.69 | B0-5 (III) [Eva04] | B2 II: | no | - |
7-084 | 4687477558279928320 | AzV 262, 2dFS 1532 | 13.49 | O9.7 Iab [Eva04] | O9.7 Iab | no | - |
7-085 | 4687478692151243776 | - | 16.30 | - | B1.5 III-II | no | - |
7-086 | 4687479070108329216 | - | 16.54 | - | B2 II | no | - |
7-087 | 4687477661359126144 | 2dFS 1543 | 15.74 | B0-5 (III) [Eva04] | B1.5 III-II | no | - |
7-088 | 4685971948951359488 | AzV 265, 2dFS 1544 | 14.66 | B2.5 (II) [Eva04] | B2.5 II | no | - |
7-089 | 4687478863949904384 | - | 16.01 | - | B1.5 III-II | no | - |
7-090 | 4687477588328496128 | - | 16.32 | - | B2 III: | no | - |
7-091 | 4687479173187523584 | Sk 97, AzV 269 | 10.90 | A 5I [San68] | A5 Iab | no | - |
7-092 | 4687478863949902336 | - | 15.91 | - | B2 II | no | - |
7-093 | 4687477592639651072 | [MA93] 1254, 2dFS 1555 | 14.68 | B1-5 (II) [Eva04] | B2 II e | yes | - |
7-094 | 4687477970596745344 | [MA93] 1255 | 15.45 | - | O9.7 V(n) | yes | - |
7-095 | 4687473435111194240 | 2dFS 1557 | 16.32 | B1-5 (III) [Eva04] | B1.5 III: | no | - |
7-096 | 4687490924220536576 | - | 15.95 | - | B2 II: | no | - |
7-097 | 4687474568982696960 | - | 12.80 | - | A7 Ib | no | - |
7-098 | 4687490924220548736 | - | 15.27 | - | B2 II: | no | - |
7-099 | 4687477764438325376 | 2dFS 1561 | 14.92 | B0.5 (IV) [Eva04] | B1 II | yes | - |
7-100 | 4687479104468055552 | - | 16.67 | - | B2 II | no | - |
7-101 | 4687426156109067008 | - | 16.32 | - | B2 III: | no | - |
7-102 | 4687477798798048640 | [M2002] SMC 51419 | 16.07 | O9 III [Lam16] | B0 V-IV | no | - |
7-103 | 4687473641269604352 | AzV 278 | 14.92 | B0 [Azz75] | B1 II | no | - |
7-104 | 4687473366391708288 | OGLE J010140.61-724251.5 | 16.72 | - | B1.5 III: | no | - |
7-105 | 4687474397183820800 | - | 16.89 | - | B1 II | no | - |
7-106 | 4687478451632815104 | - | 12.90 | - | B8 Ib | no | - |
7-107 | 4687475084378539136 | - | 16.56 | - | B1.5 III-II | no | - |
7-108 | 4687426087389568768 | AzV 283 | 14.02 | B3 [Azz75] | B2.5 Ia | no | - |
7-109 | 4687426396627189760 | - | 16.59 | - | B2 III: | no | - |
7-110 | 4687426396627182080 | - | 16.52 | - | B1.5 III: | no | - |
7-111 | 4687478138084449536 | Sk 103, AzV 289 | 12.74 | B0 Ia [Wal83] | B0 Ia | no | SB1 |
7-112 | 4687478142395205888 | [M2002] SMC 52555 | 14.17 | - | B1 Ib | no | - |
7-113 | 4687475187457720448 | 2dFS 1620 | 14.81 | B1-5 (II) [Eva04] | B3 II | no | - |
7-114 | 4687475118738255616 | - | 15.74 | - | B2 II e | no | - |
7-115 | 4687478245474401024 | 2dFS 1635 | 15.99 | O9.5 III [Eva04] | B0 IV | no | - |
7-116 | 4687474942629350272 | OGLE SMC110.6 17837 | 16.02 | - | B3 II: + B1.5 V: | no | - |
8-001 | 4689054670286883968 | Sk 37, AzV 72, 2dFS 765 | 12.98 | B8 (Iab) [Eva04] | B8 Ib | no | - |
8-002 | 4689049894283437696 | - | 15.99 | - | B2 II | no | - |
8-003 | 4689051131233907840 | - | 16.51 | - | B2 III-II | no | - |
8-004 | 4689038177610689664 | - | 16.48 | - | B2 III | no | - |
8-005 | 4689054876445285760 | AzV 84, 2dFS 786 | 15.38 | B0.5 (IV) [Eva04] | B0.7 III | no | - |
8-006 | 4689037181200454656 | AzV 92 | 13.88 | B3 [Azz75] | B2.5 II-Ib | no | - |
8-007 | 4689032778796455808 | - | 15.48 | - | B2 III: | no | - |
8-008 | 4689054945164735616 | Sk 46, AzV 96 | 12.97 | B1.5 Ia [Len97] | B1 Iab | no | - |
8-009 | 4689055009551389440 | [MA93] 462 | 14.85 | - | A5: I: + B3: Ib: | yes | - |
8-010 | 4689033539046230400 | Sk 45, AzV 98 | 11.23 | A0 Ia [Len97] | A0 Ia | no | - |
8-011 | 4689038413838504704 | - | 16.16 | - | B2 III-II | no | - |
8-012 | 4689050890715677952 | - | 16.17 | - | B2 II: | no | - |
8-013 | 4689038418150993920 | - | 16.02 | - | B2 II | no | - |
8-014 | 4689034221906397184 | - | 16.74 | - | B1.5 III-II | no | - |
8-015 | 4689075217413489408 | 2dFS 825 | 15.75 | B0.5 (V) [Eva04] | B0.5 V | no | - |
8-016 | 4689075213077405824 | - | 16.92 | - | B2 IV: | no | - |
8-017 | 4689050577143317888 | - | 15.99 | - | B2 III-II | no | - |
8-018 | 4689075492291367936 | - | 16.38 | - | B2 II: | no | - |
8-019 | 4689074392779777280 | - | 16.36 | - | B1.5 III: | no | - |
8-020 | 4689037662215249792 | OGLE SMC101.4 21947 | 15.82 | O7 Vz [Lam13] | O8 V | no | - |
8-021 | 4689032851851488000 | AzV 113 | 13.78 | O7: V [Eva04] | O7 V-IIInnn pe | yes | - |
8-022 | 4689062298150064640 | AzV 115, 2dFS 837 | 14.01 | B0.5 (II) [Eva04] | B0.7 II | no | - |
8-023 | 4689033229778815360 | - | 16.48 | - | B1.5 III: | no | - |
8-024 | 4689037456056174592 | - | 16.34 | - | O9.5 V:(n) | no | - |
8-025 | 4689033023650128256 | OGLE SMC101.4 8338 | 16.22 | - | B1.5 III | no | - |
8-026 | 4689037456056165376 | 2dFS 840 | 15.62 | B1-5 (II) [Eva04] | B2.5 II | no | - |
8-027 | 4689075625394544512 | - | 16.36 | - | B1 III: | no | - |
8-028 | 4689075183074291072 | Sk 50, AzV 116, 2dFS 843 | 14.00 | B0 (II) [Eva04] | O9.7 II-Ib(n) | no | - |
8-029 | 4689033023650129536 | - | 16.51 | - | B1 IV: | no | - |
8-030 | 4689032954930686720 | - | 15.53 | - | O6.5 Vn | no | - |
8-031 | 4689075354852418176 | - | 15.80 | - | O9.5 V | no | - |
8-032 | 4689061954552698112 | - | 15.76 | - | B1 III-II | no | - |
8-033 | 4689032950664343552 | - | 16.29 | - | B2 III: | - | - |
8-034 | 4689074358420039680 | 2dFS 854 | 15.29 | B0 (IV) [Eva04] | B0.5 III | no | - |
8-035 | 4689074461499247744 | LHA 115-S 15, Sk 51, AzV 118 | 12.49 | B0 III [Pri87] | O9.7 III: e | yes | - |
8-036 | 4689074461499243648 | - | 17.09 | - | B1 III: | no | - |
8-037 | 4688985985180031744 | - | 15.80 | - | B2 II | no | - |
8-038 | 4689074530194683520 | OGLE SMC102.1 11822 | 16.51 | - | B2 III-II | no | - |
8-039 | 4689062435588998400 | - | 15.82 | - | B0.5 V | no | - |
8-040 | 4689062435588992640 | - | 16.56 | - | B2 IV | no | - |
8-041 | 4689056525713977472 | - | 15.31 | - | B1.5 III-II | no | - |
8-042 | 4689009521606825856 | 2dFS 873 | 13.32 | A3 II [Eva04] | A5 Ib | no | - |
8-043 | 4689057934463327616 | LIN 207 | 14.71 | B1e [Lam16] | B2 Ib: e | yes | - |
8-044 | 4689062160711095808 | LIN 209, AzV 128 | 14.85 | O7 V [Mas95] | O7 Vn neb | no | - |
8-045 | 4689009521606804736 | [M2002] SMC 24096 | 15.62 | B0 V [Lam16] | B0.2 IV | no | - |
8-046 | 4689056766232136064 | AzV 129 | 14.60 | B1 [Azz75] | B1.5 II | no | - |
8-047 | 4689009452887349504 | - | 16.21 | - | B2 IV | no | - |
8-048 | 4689056456994593408 | - | 15.18 | - | B3 II | no | - |
8-049 | 4689062195070825216 | - | 16.31 | - | B1.5 III: | no | - |
8-050 | 4689062126351355776 | [M2002] SMC 24631 | 16.09 | - | O9.7 IV | no | - |
8-051 | 4689057968822902912 | - | 16.56 | - | B1.5 III | no | - |
8-052 | 4689062195070824320 | 2dFS 888 | 15.58 | O8.5 V [Eva04] | O9.2 V | no | - |
8-053 | 4689074598938187776 | AzV 135, 2dFS 889 | 14.79 | O8 III [Eva04] | O9 III | no | - |
8-054 | 4689058071902092928 | 2dFS 890 | 16.04 | B0-5 (III) [Eva04] | B2 II | no | - |
8-055 | 4689057693945011072 | 2dFS 892 | 16.01 | B1-5 (III) [Eva04] | B2 III-II | no | - |
8-056 | 4689074598938189568 | - | 15.99 | - | A2: I: + B3: | no | - |
8-057 | 4689058071902087808 | - | 16.11 | - | B2 III: | no | - |
8-058 | 4689056869311331840 | OGLE SMC101.4 8384 | 16.39 | - | B2 III: | no | - |
8-059 | 4689061748394242944 | LIN 225 | 14.98 | - | B2 II: e | no | - |
8-060 | 4689008177223333888 | [MA93] 645, 2dFS 910 | 15.17 | B1-5 (II) [Eva04] | B2 II: | no | - |
8-061 | 4689056903671040512 | [MFH2007] SMC5-78415 | 14.12 | B8 II-III [Mar07] | B9 Ib | no | - |
8-062 | 4689056697512629376 | MACHO 207.16147.27 | 16.02 | B2 IV [Mar07] | B2 III: | no | - |
8-063 | 4689008559534153856 | AzV 36F | 13.65 | B3 [Azz79] | B3 Ib | no | - |
8-064 | 4689056628793468928 | [M2002] SMC 26108 | 15.15 | B2 III [Mar07] | B3 II e | yes | - |
8-065 | 4689062916625049216 | 2dFS 918 | 15.68 | B0 (V) [Eva04] | B1 III-II | no | - |
8-066 | 4689058174981307392 | - | 15.14 | - | B2 II: | no | - |
8-067 | 4689009620370577024 | AzV 142 | 12.00 | A5 I [Neu10] | A7 Ib | no | - |
8-068 | 4689058415499454080 | - | 16.40 | - | B1.5 III: | no | - |
8-069 | 4689061507875858560 | - | 16.30 | - | B1.5 III: | no | - |
8-070 | 4689058484219114496 | 2dFS 924 | 16.14 | B0.5 (V) [Eva04] | B0.7 IV | no | - |
8-071 | 4689009349808102016 | - | 16.52 | - | B1.5 III: | no | - |
8-072 | 4689057346006300544 | TYC 9138-1910-1 | 10.98 | F0 I [Neu10] | F2: | no | - |
8-073 | 4689061812778239488 | - | 16.16 | - | O9.7 V(n) | no | - |
8-074 | 4689061507875864064 | 2dFS 926 | 16.04 | B0-5 (III) [Eva04] | B2 III-II | no | - |
8-075 | 4689061817113476608 | - | 14.86 | - | B1.5 III-II | no | - |
8-076 | 4689009796484689792 | - | 12.99 | - | A5 Ib | no | - |
8-077 | 4689061817113484416 | [MFH2007] SMC5-82923 | 14.47 | B3 III [Mar07] | B5 II | no | - |
8-078 | 4689010067007622656 | [MFH2007] SMC5-74471 | 14.49 | B2 III [Mar07] | B2 II: | yes | - |
8-079 | 4689058415499441280 | - | 16.15 | - | B1 III | no | - |
8-080 | 4689058209341023104 | OGLE SMC101.4 29381 | 16.08 | - | B2 III: | no | - |
8-081 | 4689057109829464448 | [MFH2007] SMC5-22612 | 16.54 | B2 IV [Mar07] | B2 III: | no | - |
8-082 | 4689061920192649984 | AzV 38F, 2dFS 938 | 13.06 | B9 (Iab) [Eva04] | B9 Iab | no | - |
8-083 | 4689058243700777088 | - | 12.82 | - | F0: | no | - |
8-084 | 4689010101367347456 | - | 16.08 | - | B2 IV | no | - |
8-085 | 4689057041109984128 | - | 14.79 | - | B1 II | no | - |
8-086 | 4689008903131074176 | - | 12.22 | - | A7 Ib | no | - |
8-087 | 4689061576595320064 | - | 16.24 | - | O9.7 V(n) | no | SB |
8-088 | 4689058548602468352 | - | 16.50 | - | B0 IV | no | - |
8-089 | 4689064497172976896 | - | 12.35 | - | F2: | no | - |
8-090 | 4689057487786300288 | 2dFS 950 | 15.30 | B1-5 (II) [Eva04] | B1 II | no | - |
8-091 | 4689008937490759040 | - | 13.62 | - | A0 II | no | - |
8-092 | 4689057453426569984 | [M2002] SMC 28153 | 15.64 | - | O9.2 V(n) | no | - |
8-093 | 4689063122783487872 | - | 15.95 | - | B1.5 III: | no | - |
8-094 | 4689059824248453888 | AzV 41F | 13.27 | B5 [Azz79] | B3 Ib | no | - |
8-095 | 4689060099126312192 | - | 16.71 | - | B2 III: | no | - |
8-096 | 4689057453426553728 | - | 12.68 | - | A7 Ib | no | - |
8-097 | 4689058999614762368 | - | 12.62 | - | F0: | no | - |
8-098 | 4689058656017416320 | 2dFS 972 | 15.10 | B1-5 (II) [Eva04] | B3 II | no | - |
8-099 | 4689058724736877824 | - | 15.90 | - | B2 II: | no | - |
8-100 | 4689009070622046336 | [MA93] 742 | 15.17 | - | B2 II e | yes | - |
8-101 | 4689008868771226240 | - | 12.50 | - | F0: | no | - |
8-102 | 4689009143649631232 | OGLE SMC108.8 33982 | 16.01 | - | O9.7 V | no | - |
8-103 | 4689058965255025280 | [MFH2007] SMC5-5045 | 16.17 | B2 IV [Mar07] | B2 III | no | - |
8-104 | 4689011445751390848 | 2dFS 983 | 15.90 | B0.5: (V) [Eva04] | B1.5 III: | no | - |
8-105 | 4689009070575136512 | OGLE SMC-ECL- 2967 | 15.02 | B0.5 III | no | EB [Paw13] | |
8-106 | 4689059029649262464 | [MFH2007] SMC5-38564 | 16.59 | O9 V [Mar07] | O9.7 Vnn | no | - |
8-107 | 4689011651909783424 | AzV 156 | 14.05 | A0 [Azz75] | A0 Ib | no | - |
8-108 | 4689059892967905920 | 2dFS 985 | 14.93 | B0.5 (IV) [Eva04] | B0.5 V | no | - |
8-109 | 4689011754988965888 | LHA 115-S 20, AzV 157 | 13.40 | B0e [Azz75] | B0 III: pe | yes | - |
8-110 | 4689011445751392000 | OGLE SMC108.8 34068 | 16.38 | - | B1 III: pe | no | - |
8-111 | 4689010346239909888 | - | 16.92 | - | B2 III | no | - |
8-112 | 4689011411391693440 | [M2002] SMC 30095 | 16.26 | - | B1.5 III: | no | - |
8-113 | 4689058690377131520 | AzV 159 | 12.54 | F2 [Azz75] | A5 Ib | no | - |
8-114 | 4689011411391652992 | - | 16.45 | - | B2.5 III: | no | - |
8-115 | 4689058793456333824 | - | 16.02 | - | B1 II | no | - |
8-116 | 4689058892200836096 | - | 12.73 | - | A2 Ib | no | - |
8-117 | 4689011583190270720 | - | 14.98 | - | B1.5 III-II | no | - |
Ard77: Ardeberg & Maurice (1977); Azz79: Azzopardi & Vigneau (1979); Azz82: Azzopardi & Vigneau (1982); Coe15: Coe & Kirk (2015); Dac70: Dachs (1970); Eva04: Evans et al. (2004); Eva06: Evans et al. (2006); Gar87: Garmany et al. (1987); Gra12: Graus et al. (2012); Hil05: Hilditch et al. (2005); Hum83: Humphreys (1983); Hum91: Humphreys et al. (1991); Lam13: Lamb et al. (2013); Mar07: Martayan et al. (2007); Mas95: Massey et al. (1995); Mas00: Massey et al. (2000); Mas02: Massey (2002); Mas04: Massey et al. (2004); Mas09: Massey et al. (2009); Men06: Mennickent et al. (2006); Mok06: Mokiem et al. (2006); Mor03: Morrell et al. (2003); Neu10: Neugent et al. (2010); Neu18: Neugent et al. (2018); Pau12: Paul et al. (2012); Paw16: Pawlak et al. (2016); Pri87: Prinja (1987); San68: Sanduleak (1968); She13: Sheets et al. (2013); Smi97: Smith Neubig & Bruhweiler (1999); Tes87: Testor & Lortet (1987); Tes01: Testor (2001); Wal83: Walborn (1983); Wal00: Walborn et al. (2000); Wal02: Walborn et al. (2002); Zas14: Zasche et al. (2014); Zic96: Zickgraf et al. (1996);
Appendix B Cross matches with additional catalogues
B.1 ESO archive
We cross-matched the BLOeM catalogue with spectroscopic databases in the ESO archive, defining a search radius of 3” per target. We retrieve a total of 1,988 spectra for 202 stars out of the 929 in our sample. These spectra were acquired with various instruments of the VLT. These data will be used to improve the quantitative spectroscopy in subsequent papers.
B.2 Hubble UV Legacy Library of Young Stars as Essential Standards (ULLYSES)
ULLYSES is a legacy survey of the Hubble Space Telescope (HST), which includes the acquisition of high-resolution UV spectra for 128 massive stars in the SMC777https://ullyses.stsci.edu/ullyses-targets-smc.html. The programme also includes a follow-up with the X-SHOOTER spectrograph of the VLT to obtain a visual and infrared coverage of the targets Vink et al. (2023). A cross-match of the BLOeM sample with the ULLYSES sample, using a search radius of 3”, resulted in an overlap of 43 targets. From ULLYSES and XShootU there will be broad wavelength coverage of the UV and visible spectrum of these objects that will also be used to inform the analysis of these targets.
B.3 X-ray catalogues
The SMC was extensively observed in X-rays. The largest modern X-ray observatories, XMM-Newton and Chandra, which operate in 0.2-12.0 keV range, conducted surveys of the entire SMC galaxy (Laycock et al. 2010; Sturm et al. 2013). The deep observations of individual fields, such as the SMC Wing and NGC 346 star cluster have also been performed (Nazé et al. 2004; Oskinova et al. 2013). However, despite these efforts X-ray emission of individual “normal” massive OB stars is below current detection limits. On the other hand, X-ray detections of massive stars in the SMC allow to select binary stars. Specifically, X-ray detections are excellent tracers of CWBs, some of which are significantly more X-ray bright compared to single stars (Corcoran et al. 1996; Sana et al. 2006; Oskinova 2005; Nazé et al. 2007). But best of all, X-ray detections are suited to identify high-mass X-ray binaries (HMXB), where a compact object is accreting matter of its OB-type companion. HMXBs are X-ray variable, especially so are BeXRBs where the donor stars have OBe spectral type. BeXRBs are transient X-ray sources, and may remain quiescent over long periods of time and could be detected only during outbursts.
To explore X-ray properties of our targets, the BLOeM catalog was cross-correlated with catalogues produced by the XMM-Newton, Chandra, eROSITA, and Rosat X-ray telescopes. Only eight BLOeM stars are firmly detected, while the positional uncertainty of one X-ray source (6-116) precludes its firm detection. The detected sources are listed in Table 4. There are four already known HMXBs among them. Four other X-ray sources may be either CWBs or newly discovered HMXBs. The BLOeM spectroscopy will shed light on their nature, since HMXBs are SB1 systems while CWBs are likely to be SB2. Some sources listed in Table 4 have different fluxes according to different catalogues. This may reflect true source variability, e.g. in case of BeXRBs. To estimate the X-ray luminosity, we adopted a neutral hydrogen column density cm-2 and a power-law spectrum with for all objects.
BLOeM ID | Alias | Uncertainty | Separation | Spectral Type | Catalog | Remarks | ||
X-ray (′′) | (′′) | [erg cm-1 s-1] | (erg s-1) | |||||
3-042 | AzV 26 | 1.9 | 0.7 | O6 I(f)+O7.5 | 4e32 | CSC v.2 | CWB (?) | |
2-055 | AzV 102 | 0.95 | 0.7 | O9.7 V:n e | 2e34 | SMCDFSCXO | HMXB SXP 8.80 | |
1.0 | 1.8 | 3e33 | 4XMM-DR13s | |||||
1.3 | 0.6 | 1e35 | CXOGSGSRC | |||||
2.6 | 0.2 | 2e34 | CSC v.2 | |||||
8-029 | - | 1.9 | 0.6 | B1 IV: | 3e34 | 4XMM-DR13 | HMXB (?) | |
2-082 | AzV 138 | 1.7 | 0.4 | O9.2 III pe | 6e33 | CSC v.2 | HMXB | |
2-116 | AzV 154 | 0.9 | 2.4 | sgB[e] | 4e33 | 4XMM-DR13s | HMXB | |
1.4 | 1.2 | 6e33 | SMCPSCXMM | |||||
0.6 | 1e33 | CSC v.2 | ||||||
4-026 | Cl* NGC 346 MPG 217 | 0.1 | O9.5 IIIpe | 5e32 | CSC v.2 | HMXB | ||
4-113 | OGLE SMC-SC9 131970 | 0.86 | 1.2 | B2.5 II pe | 2e34 | SMCPSCXMM | HMXB | |
1.5 | 1.2 | 3e34 | CXOGSGSRC | |||||
1.7 | CSC v.2 | |||||||
1-102 | AzV 345a | 0.1 | O6 III(n) | 1e32 | CSC v.2 | CWB ? HMXB? | ||
6-116 | 2dFS 3274 | 1.7 | 1.4 | A7 Iab | 7e33 | SMCPSCXMM | spurious? | |
1.1 | 2.5 | 1e34 | 4XMM-DR13s |
CSC v.2
Chandra Source Catalog v.2 (Evans et al. 2010)
SMCDFSCXO
SMC Deep Fields X-Ray Point Source Catalog (Laycock et al. 2010)
4XMM-DR13s
XMM-Newton Serendipitous Source Catalog from Stacked Observations(Traulsen et al. 2020)
4XMM-DR13
XMM-Newton Serendipitous Source Catalog DR13 (Webb et al. 2020)
CXOGSGSRC
Chandra ACIS GSG Point-Like X-Ray Source Catalog (Wang et al. 2016)
SMCPSCXMM
SMC XMM-Newton Point Source Catalog (Sturm et al. 2013)
B.4 OGLE Photometry
A large number of the BLOeM targets has been monitored by the OGLE photometric survey. Out of the 929 targets, 847 were observed in the OGLE-III (Udalski 2003) and 785 continue to be observed in the OGLE-IV (Udalski et al. 2015). Among them, there are 82 objects identified as binary systems by Pawlak et al. (2016), including 74 eclipsing and 8 ellipsoidal binaries.
B.5 TESS Photometry
All BLOeM targets have been and continue to be periodically observed by the all-sky time-series photometry TESS space mission (Ricker et al. 2015). An initial survey of photometric variability for XShootU targets in the LMC and SMC was performed by Bowman (2024), who found similar stochastic low-frequency (SLF) variability to Galactic OB stars (Bowman et al. 2019). In the future, we will extract light curves for as many BLOeM targets as possible to identify possible eclipsing binaries, stars with rotational modulation, and pulsations. In so doing, this will provide complementary constraints and inform the search for multiplicity for the BLOeM sample.